### 15 Mar 2026 v1

# Mobility Edge for the Anderson Model on Random Regular Graphs

Abstract.
We determine the phase diagram of the Anderson tight-binding model on regular random graphs with Gaussian disorder distribution and sufficiently large degree. In particular, we prove that if the degree is fixed and the number of vertices goes to infinity, the spectrum asymptotically consists of a finite delocalized interval surrounded by two unbounded localized components. Our argument uses a recent description of the spectrum of the tight-binding model on the Bethe lattice (Aggarwal–Lopatto, 2025). By viewing the Bethe lattice as the local limit of a random regular graph, and proving several preliminary estimates, we transfer this characterization of the spectrum of the limiting model to its finite approximants.
1. Introduction
1.1. Background
In a disordered medium, such as a doped semiconductor, the diffusion of waves ceases completely when the disorder is sufficiently strong. This principle, known as Anderson localization, has played a fundamental role in the development of solid state physics [3]. Significant effort has been devoted to studying Anderson localization from first principles, and a detailed picture based on the methods of theoretical physics has emerged, augmented by a substantial body of numerical evidence. However, our mathematically rigorous understanding remains highly incomplete.
A standard setting for studying wave transport in disordered media is the tight-binding model, which served as the basis for Anderson’s original investigations [12]. Let $\mathcal{G}$ be a graph (finite or infinite), let $A$ denote the adjacency operator of $\mathcal{G}$, and let $V$ denote a diagonal operator of independent, identically distributed (i.i.d) random variables. A particle whose time evolution is governed by $A$ moves freely by nearest-neighbor hopping, displaying diffusive behavior. In contrast, the time evolution induced by $V$ traps a particle at its initial site. The tight-binding Hamiltonian is defined as $H_{t}=-tA+V$; it represents a contest between the opposite influences of $A$ and $V$ mediated by $t$, the inverse disorder strength. When $t$ is sufficiently small, $V$ dominates, and particle remains frozen in position space. However, when $t$ is sufficiently large, it is expected that the particle may diffuse, at least for certain energies and graph geometries, leading to a region (or regions) of delocalization in the spectrum separated from regions of localization by sharp transitions called mobility edges [9].
An important special case is when $\mathcal{G}$ is the infinite $d$-regular tree, also known as the Bethe lattice. While simpler than more physically realistic graphs, such as $\mathbb{Z}^{3}$, its lack of cycles renders it more amenable to analysis [2, 1, 29, 28]. Because of this, it has served as a distinguished testbed for theoretical investigations. For instance, the self-consistent theory of Anderson localization was developed using equations which are approximate on a true lattice, but become exact on the Bethe lattice [2, 1]. Much is now known rigorously about this model, including delocalization at small energies [27, 26, 7, 22], localization at large energies [10, 6], and a nearly complete description of the phase diagram at low disorder (omitting only small intervals around the expected locations of the mobility edges) [20]. Further, [8] established a criterion for absolutely continuous spectrum based on the fractional moments of resolvent entries. This was used in [13], along with the Simon–Wolff criterion for localization [32, 6], to provide quantitative bounds on the endpoints of the regions of localization and delocalization, mapping a large part of the spectrum. As the degree of the Bethe lattice tends to infinity, these bounds become increasingly better and characterize the entire the spectrum in the limit (but not for any finite degree). Recently, building on the work of [13, 32, 6, 8, 4], mobility edges separating regions of delocalization and localization were shown to exist in [5] when the degree $d$ is sufficiently large and the potential has unbounded support.
A close cousin of the Bethe lattice is the random regular graph, which is generally understood to be its correct finite-size analog. (Simply truncating the Bethe lattice to a finite tree leads to a graph where most vertices are leaves, which produce undesirable boundary effects.) Since random regular graphs converge locally to the Bethe lattice, their spectral and eigenfunction observables measure how Bethe-lattice physics is modified by long loops and finite-size effects, providing a controlled high-dimensional reference for comparison with finite-dimensional Anderson localization [16, 17, 33]. Further, there has been a resurgence of interest in the tight-binding model on the random regular graph due to its connection with many-body localization, a phenomenon that has recently attracted significant interested in the physics community [30]. Specifically, in suitable many-body problems, certain hierarchical models in Fock space can be approximately mapped to tight-binding models on locally tree-like graphs, for which a random regular graph is a useful proxy [34].
There have been only a few rigorous results for the tight-binding model on random regular graphs. In [23], it was shown that for compactly supported and bounded disorder densities, the integrated density of states converges to the integrated density of states of the Bethe lattice, as $d$ is fixed and the number of vertices $n$ tends to infinity, with a rate of convergence of order $1/\ln n$. This work also showed for that if the eigenvectors in a finite energy window in the absolutely continuous spectrum for the Bethe lattice are averaged, weighted by their squared amplitudes at a uniformly random vertex $x_{0}$, the resulting weighted mass does not concentrate in small balls around $x_{0}$; this is a form of eigenvector delocalization. Further, [24] showed a criterion for the presence of Poisson spectral statistics based on the fractional moment decay of off-diagonal resolvent entries and applied it to the edges of the spectrum (for compactly supported disorder distributions). Finally, it was shown in [11] that in the weak-disorder regime, for energy windows in the absolutely continuous spectrum for the Bethe lattice, most eigenfunctions are asymptotically equidistributed on their coordinates (again for compactly supported disorder distributions).
In this article, we are interested in showing the existence of mobility edges that sharply separate regions of delocalization and localization in random regular graphs. Our main result establishes this for sufficiently large (but fixed) degrees $d$ and a wide range of disorder strengths when the disorder distribution is Gaussian. We choose to concentrate on this setting for brevity, and because it already illustrates the phenomenology predicted by the physics literature. However, our methods are rather flexible, and we comment on possible generalizations below, after stating our main result.
1.2. Main Result
Fix an integer $d\geq 3$, and let $\mathbb{N}$ denote the positive integers. For each $N\in\mathbb{N}$, let $G_{N}$ be a (uniformly) random $d$-regular graph on the vertex set $N]=\{1,\dots,N\}$, and let $A_{N}$ denote its adjacency matrix. Let $(g_{i})_{i=1}^{\infty}$ be i.i.d. Gaussian random variables with mean zero and variance one. Define the diagonal matrix
$V_{N}=\mathrm{diag}(g_{1},\dots,g_{N}).$
For every $t>0$, we define the symmetric random matrix
$W_{N,t}=-tA_{N}+V_{N}.$ (1.1)
We will often abbreviate $W_{N,t}=W_{N}$, as the value of $t$ will never depend on $N$.
Let $(\lambda_{k},u_{k})_{k=1}^{N}$ be an orthonormal eigendecomposition of $W_{N}$, so that
$W_{N}u_{k}=\lambda_{k}u_{k},\qquad\langle u_{k},u_{\ell}\rangle=\delta_{k\ell},$
where $\delta_{k\ell}$ is the Kronecker delta function. The eigenvectors $u_{k}$ are determined only up to a sign, which we may select arbitrarily. We let $\Lambda$ denote the multiset of eigenvalues of $W_{N}$ (which may have duplicated elements due to repeated eigenvalues), and given an interval $I\subset\mathbb{R}$, we let $\Lambda_{I}$ be the multiset of eigenvalues contained in $\mathbb{R}$. While these objects depend on $d$, $t$, and $N$, we suppress this dependence in the notation.
To study (de)localization of the eigenvectors $u_{k}$, we follow [18] and consider averages of (powers of) coordinates of eigenvectors associated with eigenvalues in small intervals of the real line.
Definition 1.1.
Fix a compact interval $I\subset\mathbb{R}$ and index $j\in[N]$. We define
$P_{I}(j)=\frac{N}{|\Lambda_{I}|}\sum_{\lambda_{i}\in\Lambda_{I}}u_{i}(j)^{2},\qquad Q_{I}=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{2}.$
To understand the meaning of these observables, we consider two extreme cases. First, suppose that all eigenvectors with eigenvalues in $I$ are as delocalized as possible, so that $u_{i}(j)=N^{-1/2}$ for all $j\in[N]$ and $i$ such that $\lambda_{i}\in I$. Then $P_{I}(j)=1$ for all $j$, implying $Q_{I}=1$. In particular, $Q_{I}$ stays bounded as $N$ grows. Next, consider the case where $u_{i}(j)=\delta_{ij}$ for all $i,j\in[N]$, in which all eigenvectors are maximally localized. Then $P_{I}(j)=0$ if $\lambda_{j}\notin I$ and $N|\lambda_{I}|^{-1}$ if $\lambda_{j}\in I$, implying $Q_{I}=N|\Lambda_{I}|^{-1}$. If $|\Lambda_{I}|\leq C|I|N$ for all $N$, where $C>0$ is a constant that does not depend on $I$ or $N$, and $|I|$ denotes the length of $I$, then $Q_{I}$ becomes unbounded as $|I|$ tends to zero.
Therefore, it is reasonable to distinguish between delocalization and localization based on the behavior of $Q_{I}$ as $I$ shrinks (assuming the relevant upper bound on $|\Lambda_{I}|$, which will be shown in Lemma 3.4 below). This observation motivates the next definition.
Definition 1.2.
Fix an energy $E\in\mathbb{R}$. We say that delocalization holds at $E$ for $W_{N}$ if there exist constants $C(E)>1$ and $c(E)>0$ such that for all $\varepsilon\in(0,c)$,
$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\leq C\Big)=1.$
where $I(\varepsilon)=[E-\varepsilon,\,E+\varepsilon]$. We say that localization holds at $E$ if for every $D>0$, there exists $c(D)>0$ such that for all $\varepsilon\in(0,c)$,
$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\geq D\Big)=1.$
We now state our main theorem. Let $\rho$ denote the density function of a mean zero, variance one Gaussian random variable. The scaling of the strength disorder as $(d\ln d)^{-1}$ is conventional; it ensures the mobility edge locations tend to a constant-order limit as $d$ grows large.
Theorem 1.3.
For every real number $\mathfrak{L}>0$, there exists a constant $d_{0}(\mathfrak{L})>0$ such that the following holds for all $d\geq d_{0}$. Fix $g\in\mathbb{R}$ such that
$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$
and let $t=g(d\ln d)^{-1}$. Then there exists $\mathfrak{E}>0$ such that the set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$. Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$, localization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|>\mathfrak{M}\}$, and delocalization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|<\mathfrak{M}\}$.
1.3. Proof Ideas
As mentioned previously, the spectral decomposition of the tight-binding model of the Bethe lattice was characterized in [5] (for the disorder regime considered in our main result). Since the Bethe lattice is the local limit of random regular graphs, it natural to attempt to use the Bethe lattice result to prove Theorem 1.3, and this is indeed our strategy. However, this route requires suitable estimates for the tight-binding model on the regular random graph, which we enumerate in Section 3. Specifically, we require concentration for fractional moments of the off-diagonal resolvent entries (Lemma 3.3), upper and lower bounds on the eigenvalue density (Lemma 3.4), convergence of the diagonal resolvent entries of the random regular graph to the diagonal resolvent entry at the root of the Bethe lattice (Lemma 3.5), and, in the delocalized regime, a bound on the second moment of the imaginary of part of the resolvent at the root of the Bethe lattice that is uniform in the imaginary part of the spectral parameter (Lemma 3.6). Of these, the last is perhaps the most technically challenging. We prove it by first establishing a Cauchy-type tail bound on the imaginary part of the relevant resolvent entry, then bootstrapping that bound to an improved polynomial tail decay bound.
Generalizations of our result are possible in at least two directions. First, we have chosen to concentrate on the case of Gaussian disorder to streamline the presentation. The spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] for a wide class of unbounded potentials; in the general case, it is slightly more complicated to describe, with interleaved regions of localization and delocalized determined by the disorder density. The techniques we present suffice to establish the analogous result for random regular graphs with corresponding potentials under the additional assumption that density decays sufficiently rapidly (for example, exponentially), so that the concentration argument for Lemma 3.3 below, based on Gaussian concentration for Lipschitz functions, can be replaced by one based on truncation and Talagrand’s concentration inequality for bounded distributions.
Second, the requirement that $d$ be very large in Theorem 1.3 comes primarily from the analogous Bethe lattice result in [5], not our arguments that show its conclusions transfer to random regular graphs. The only input we require from the limiting model is an understanding of how the imaginary part of the resolvent at the root behaves in the upper half plane as the imaginary part of its spectral parameter goes to zero (see Lemma 2.3), along with a mild lower bound on $d$ (which could be improved). In particular, it seems likely that our methods suffice to transfer the nearly complete description of the phase diagram for the Bethe lattice at low disorder recently established in [20] to random regular graphs.
1.4. Acknowledgments
P. L. thanks A. Aggarwal for several helpful conversations. Both authors were partially supported by NSF grant DMS-2450004.
2. Proof Outline
2.1. Notations and Conventions
We set $\mathbb{H}=\{z\in\mathbb{C}:\operatorname{Im}z>0\}$. For $z\in\mathbb{H}$, we typically use the notation $z=E+\mathrm{i}\eta$, where $E=\operatorname{Re}z$ and $\eta=\operatorname{Im}z$. For all $z\in\mathbb{H}$, we define the resolvent of $W_{N}$ by
$G_{N}(z)=(W_{N}-z)^{-1},$
and we frequently abbreviate $G=G_{N}$.
Given an integer $d\geq 3$, we let $\mathbb{T}=\mathbb{T}_{d}$ be a $d$-regular rooted tree graph. We denote the root by 0, and the vertex set of $\mathbb{T}$ by $\mathbb{V}=\mathbb{V}^{(d)}$. For all $v\in\mathbb{V}$, we let $\delta_{v}$ denote the unit vector supported at $v$, meaning that $\delta_{v}(w)=\delta_{vw}$ for all $w\in\mathbb{V}$.
We define the adjacency operator $A=A^{(d)}$ for $\mathbb{T}$ by setting $\langle\delta_{v},A\delta_{w}\rangle=1$ if there is an edge between $v$ and $w$ in $\mathbb{T}$ and setting $\langle\delta_{v},A\delta_{w}\rangle=0$ otherwise. It extends by linearity to a bounded operator $\ell^{2}(\mathbb{V})$. Further, we let $(g_{v})_{v\in\mathbb{V}}$ be collection of i.i.d. Gaussian random variables with mean zero and variance one. We define an unbounded operator $V=V^{(d)}$ on $\ell^{2}(\mathbb{V})$ by setting $\langle\delta_{v},V\delta_{w}\rangle=g_{v}$ if $v=w$ and $\langle\delta_{v},V\delta_{w}\rangle=0$ otherwise, then extending by linearity. Then, for every $t>0$, we define the operator
$H_{d,t}=-tA^{(d)}+V^{(d)}.$
It is well-known that $H_{d,t}$ is unbounded and self-adjoint on the domain of $V^{(d)}$; see, e.g., [25].
2.2. Mobility Edge on the Bethe Lattice
The following theorem is a specialization of [5, Theorem 1.3] to the case of the Gaussian disorder.
Theorem 2.1.
For every real number $\mathfrak{L}>0$, there exists a constant $d_{0}(\mathfrak{L})>0$ such that the following holds for all $d\geq d_{0}$. Fix $g\in\mathbb{R}$ such that
$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$ (2.1)
and let $t=g(d\ln d)^{-1}$. Then there exists $\mathfrak{E}>0$ such that the solution set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$. Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$ and the following claims hold.
  1. (1)
    For all $E\in\mathbb{R}$ such that $|E|>\mathfrak{M}$, we have $\lim_{\eta\rightarrow 0}\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=0$ in probability.
  2. (2)
    For all $E\in\mathbb{R}$ such that $|E|<\mathfrak{M}$, there exists $c(E)>0$ such that
    $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
Remark 2.2.
In [5], the conclusion of the theorem is stated in a slightly different way, involving absolutely continuous and pure point spectrum. However, it is readily verified by inspecting the proof that the claims in Theorem 2.1 are proved as intermediate steps (see, e.g., [5, Lemma 2.7].) We also note that the proof in [5] is for a tree whose root degree is one less than each other vertex. However, as mentioned in that reference, is it straightforward to see that the same argument gives the conclusion for $\mathbb{T}_{d}$.
2.3. Resolvent Criteria for Localization and Delocalization
The following lemma is the primary novelty of this paper. It shows that the conclusions of Theorem 2.1 regarding the resolvent of the Anderson model at the root of Bethe lattice imply the corresponding (de)localization behavior for the matrices $W_{N}$. The proof is given in Section 3.
Lemma 2.3.
Fix $E\in\mathbb{R}$, $d\in\mathbb{N}$ such that $d\geq 20$, and $t>0$.
  1. (1)
    If $\lim_{\eta\rightarrow 0}\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=0$ in probability, then localization holds at $E$ for $W_{N}$.
  2. (2)
    If there exists $c>0$ such that $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c$, then delocalization holds at $E$.
2.4. Proof of Main Result
We are now able to prove our main result.
Proof of Theorem 1.3.
This is an immediate consequence of Theorem 2.1 and Lemma 2.3. ∎
3. Localization and Delocalization Criteria
This section is devoted to the proof of Lemma 2.3.
3.1. Resolvent Estimates
The following lemma collects some basic facts about $G$, which follow readily from the definition of $G$ and the spectral theorem. Proofs may be found in, e.g., [15].
Lemma 3.1.
Fix $z=E+\mathrm{i}\eta\in\mathbb{H}$.
  1. (1)
    For all $i,j\in[N]$, we have the deterministic bound
    $|G_{ij}(z)|\leq\eta^{-1}.$ (3.1)
  2. (2)
    For every $i\in[N]$, we have
    $\sum_{j=1}^{N}\big|R_{ij}(z)\big|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (3.2)
We also require the next lemma, on the uniform integrability of $R_{00}$.
Lemma 3.2.
Fix $E\in\mathbb{R}$. The set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\}_{\eta\in(0,1)}$ is uniformly integrable.
Proof.
This is immediate from [5, Lemma 3.3] applied to the moments of order $(s+1)/2$. ∎
3.2. Preliminary Lemmas
To prepare for the proof of Lemma 2.3, four main ingredients are needed. First, we state a concentration inequality for the diagonal resolvent entries, which is shown in Section 4.
Lemma 3.3.
Fix $s\in(0,1/2)\cup[1,\infty)$. There exists $C(s)>1$ such that for all $E\in\mathbb{R}$ and $\eta\geq(\ln N)^{-1}$,
$\mathbb{P}\!\left(\left|\frac{1}{N}\sum_{j=1}^{N}\big(\operatorname{Im}G_{jj}(z)\big)^{s}-\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s}\Big]\right|\leq\eta\right)\geq 1-CN^{-100}.$ (3.3)
The following lemma provides upper and lower bounds on the eigenvalue density of $W_{N}$. It is proved in Section 5.
Lemma 3.4.
For all $M>0$, there exists $c(M)>0$ such that the following holds. Fix $E\in[-M,M]$ and $\eta\in(0,1)$, and set $I=[E-\eta,E+\eta]$. Then
$\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\geq cN\eta\right)=1,\qquad\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\leq c^{-1}N\eta\right)=1.$ (3.4)
The next lemma shows that the diagonal resolvent entries of the Anderson model on the $d$-regular graph converge to the resolvent entry at the root of the $d$-regular tree. The proof is postponed to Section 6.
Lemma 3.5.
Fix $d\geq 3$, $t>0$, and $z=E+\mathrm{i}\eta\in\mathbb{C}$ with $\eta>0$. There exists a coupling of $W_{N,t}$ and $H_{d,t}$ such that for all $k\in\mathbb{N}$,
$\lim_{N\rightarrow\infty}G_{kk}(z)=R_{00}(z)$ (3.5)
in probability.
Finally, we state a bound on the second moment of $\operatorname{Im}R_{00}(z)$ that is uniform in $\eta$. The proof is given in Section 7.
Lemma 3.6.
Fix $E\in\mathbb{R}$ and $d\geq 20$. Suppose that there exists $c>0$ such that
$\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
Then
$\sup_{\eta\in(0,1]}\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\big)^{2}\Big]<\infty.$ (3.6)
3.3. Localization
For all $s>0$, we define
$Q_{I}(s)=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{s}.$ (3.7)
Recalling Definition 1.1, we have $Q_{I}=Q_{I}(2)$.
Lemma 3.7.
Fix $s,\delta\in(0,1)$ and $E\in\mathbb{R}$, and suppose that
$\lim_{\eta\rightarrow 0}R_{00}(E+\mathrm{i}\eta)=0$
in probability. Let $I(\eta)=[E-\eta,E+\eta]$. Then there exists $\eta_{0}(s,\delta,E)>0$ such that for all $\eta\in(0,\eta_{0})$,
$\lim_{N\to\infty}\mathbb{P}\!\left(Q_{I}\!\left(\frac{s}{2}\right)\leq\delta\right)=1.$
Proof.
We set $z=E+\mathrm{i}\eta$, where $\eta$ is a parameter that may vary throughout the proof (but never depends on $N$). By the spectral theorem, we have for all $\eta>0$ that
$G_{jl}(z)=\sum_{k=1}^{N}\frac{u_{k}(j)u_{k}(l)}{\lambda_{k}-z},\qquad\operatorname{Im}G_{jj}(z)=\sum_{k=1}^{N}\frac{\eta\,|u_{k}(j)|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}.$ (3.8)
For $\lambda_{k}\in I(\eta)$ we have $(\lambda_{k}-E)^{2}+\eta^{2}\leq 2\eta^{2}$, hence
$\sum_{k\in\Lambda_{I}}\big|u_{k}(j)\big|^{2}\leq\sum_{k=1}^{N}\frac{2\eta^{2}\big|u_{k}(j)\big|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}=2\eta\operatorname{Im}G_{jj}(z).$
Multiplying by $N/|\Lambda_{I}|$ gives
$P_{I}(j)\leq\frac{2N\eta}{|\Lambda_{I}|}\,\operatorname{Im}G_{jj}(z),$ (3.9)
and therefore
$Q_{I}\!\left(\frac{s}{2}\right)=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{s/2}\leq\left(\frac{2N\eta}{|\Lambda_{I}|}\right)^{s/2}\cdot\frac{1}{N}\sum_{j=1}^{N}\bigl(\operatorname{Im}G_{jj}(z)\bigr)^{s/2}.$ (3.10)
Combining (3.3), (3.4), and (3.10) yields a constant $c_{1}(E)>0$ such that
$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\!\left(\frac{s}{2}\right)\leq c_{1}^{-s/2}\left(\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]+\eta\right)\right)=1.$ (3.11)
Next, recalling (3.5) and using $|G_{11}(z)|\leq 1/\eta$ (from (3.1)), the dominated convergence theorem yields
$\lim_{N\to\infty}\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]=\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(z)\big)^{s/2}\Big].$ (3.12)
By Lemma 3.2, the set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta))^{s/2}\}_{\eta\in(0,1)}$ is uniformly integrable. Then the assumption that $\lim_{\eta\rightarrow 0}\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=0$ in probability yields a constant $c_{2}(s,\delta,E)>0$ such that for all $\eta\in(0,c_{2})$, we have
$\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(z)\big)^{s/2}\Big]<\frac{c_{1}^{s/2}\delta}{8}.$
This implies that for sufficiently large $N$,
$\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]\leq\frac{c_{1}^{s/2}\delta}{4}.$
Combining this with (3.11) and taking $\eta_{0}<c_{1}^{s/2}\delta/2$ completes the proof. ∎
Proposition 3.8.
Under the hypotheses of Lemma 3.7, for every $D>0$ there exists $\eta_{0}(D,E)>0$ such that for all $\eta\in(0,\eta_{0})$,
$\lim_{N\to\infty}\mathbb{P}\!\big(Q_{I}\geq D\big)=1.$
Proof.
Let $\varepsilon\in(0,1)$ and $p,q>1$ be parameters such that $1/p+1/q=1$. By Hölder’s inequality,
$N=\sum_{j=1}^{N}P_{I}(j)=\frac{1}{N}\sum_{j=1}^{N}(NP_{I}(j))^{\varepsilon}(NP_{I}(j))^{1-\varepsilon}\leq\Bigl(N^{\varepsilon p}Q_{I}(\varepsilon p)\Bigr)^{1/p}\,\Bigl(N^{(1-\varepsilon)q}Q_{I}((1-\varepsilon)q)\Bigr)^{1/q}.$
With $s\in(0,1)$, we fix
$\varepsilon=\frac{4}{4-s},\qquad p=2-\frac{s}{2}.$
Then $\varepsilon p=s/2$, $(1-\varepsilon)q=2$, and (p/q)=1s/2(p/q)=1-s/2, yielding
$1\leq Q_{I}(s/2)\cdot Q_{I}(2)^{\,1-s/2}=Q_{I}(s/2)\cdot Q_{I}^{\,1-s/2}.$
Thus, if $Q_{I}(s/2)\leq\delta$ then $Q_{I}\geq\delta^{-1/(1-s/2)}$, and the claim follows from Lemma 3.7 by choosing an arbitrary value $s\in(0,1)$ and $\delta$ small enough as a function of $D$. ∎
3.4. Delocalization
Proposition 3.9.
Fix $E\in\mathbb{R}$ and assume that there exists a constant $c>0$ such that
$\liminf_{\eta\rightarrow 0}\mathbb{P}\!\big(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\big)>c.$
Then there exist $\eta_{0}(E)>0$ and $C(E)>1$ such that for all $\eta\in(0,\eta_{0})$,
$\lim_{N\to\infty}\mathbb{P}\!\left(Q_{I}\leq C\right)=1,$
where $I=[E-\eta,E+\eta]$.
Proof.
Using (3.9), we have
$Q_{I}=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{2}\leq\frac{4N^{2}\eta^{2}}{|\Lambda_{I}|^{2}}\cdot\frac{1}{N}\sum_{j=1}^{N}(\operatorname{Im}G_{jj}(z))^{2}.$ (3.13)
Using the concentration bound (3.3) for $r=2$, we obtain
$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\leq 4c_{1}^{-2}\cdot\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]+\eta\right)=1.$ (3.14)
From (3.5) and together with the deterministic bound $|G_{11}|\leq\eta^{-1}$ (from (3.1)), the dominated convergence theorem gives
$\lim_{N\to\infty}\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]=\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(z)\big)^{2}\Big].$ (3.15)
Finally, combining (3.14), (3.15), and Lemma 3.6 finishes the proof. ∎
3.5. Conclusion
Proof of Lemma 2.3.
Combining Proposition 3.8 and Proposition 3.9 completes the proof. ∎
4. Concentration of Resolvent Moments
In this section, we prove Lemma 3.3 using the Gaussian concentration inequality (for the randomness in $V$), a martingale argument based on switching edges in the regular graph (for the randomness in $A$), and a truncation argument.
4.1. Preliminary Reduction
For all symmetric $N\times N$ matrices $M$, $v\in\mathbb{R}^{N}$, $z\in\mathbb{H}$ and $s>0$, define
$F_{s}(z,v,M)=\frac{1}{N}\sum_{i=1}^{N}\big(\operatorname{Im}S_{ii}(z)\big)^{s},\qquad S=\big(M+\mathrm{diag}(v_{1},\dots,v_{N})-z\big)^{-1}.$
We frequently abbreviate $F_{s}(z)=F_{s}(z,v,M)$, and we set $\tilde{F}_{s}=F_{s}(z,V_{N},-tA_{N})$, so that with this choice of arguments, $S(z)=G(z)$. We establish the following preliminary concentration estimate in Section 4.4 below, after some preliminary estimates on $F_{s}(z)$ (and a truncated counterpart). The first exponential term comes from Gaussian concentration in $V_{N}$, while the second comes from a coarse Azuma–Freedman bound for the graph randomness in $A_{N}$.
Lemma 4.1.
Fix $z=E+\mathrm{i}\eta\in\mathbb{H}$ and $d\geq 3$. There exists a constant $C(d)>0$ such that the following hold.
  1. (1)
    For every $s\geq 1$ and every $u>0$,
    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u\right)\leq 2\exp\left(-\frac{N\eta^{2s+2}}{8s^{2}}u^{2}\right)+2\exp\left(-\frac{N\eta^{2s+2}}{Ct^{2}s^{2}}u^{2}\right).$ (4.1)
    Further, the same bound holds for the corresponding averages with $\operatorname{Im}S_{ii}(z)$ replaced by $\operatorname{Re}S_{ii}(z)$.
  2. (2)
    Fix $s\in(0,1/2)$. Then for every $u>0$ and every $\delta\in(0,\eta^{-1}]$,
    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u+2\delta^{s}\right)\leq 2\exp\left(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\right)+2\exp\left(-\frac{N\eta^{4}}{Ct^{2}s^{2}}u^{2}\right).$ (4.2)
Lemma 3.3 is an immediate consequence of Lemma 4.1.
Proof of Lemma 3.3.
We begin with the case $s\geq 1$. This follows from setting $u=\eta$ in (4.1). Next, for $s\in(0,1/2)$, we may suppose that $\eta\leq 1$ (by decreasing $\eta$ if necessary). Then the estimate follows from setting $u=\eta/2$ and $\delta=(\eta/4)^{1/s}$ in (4.2). ∎
4.2. Gaussian Concentration
The following Gaussian concentration inequality is standard (see, e.g., [19, Theorem 5.6]). We say a function $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}$ is $L$-Lipschitz if $|f(x)-f(y)|\leq L\|x-y\|_{2}$ for all $x,y\in\mathbb{R}^{N}$.
Lemma 4.2.
Let $X=(X_{1},\dots,X_{N})$ be a vector of $N$ independent mean zero, variance one Gaussian random variables. Let $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}$ be an $L$-Lipschitz function. For all $t>0$,
$\mathbb{P}\left(f(X)-\mathbb{E}\big[f(X)\big]\geq t\right)\leq\exp\left(-\frac{t^{2}}{2L^{2}}\right).$
We use the notation $\partial_{k}F_{s}(z)$ to denote the partial derivative in $v_{k}$, and we write $\nabla F_{s}(z)$ for the gradient with respect to $v\in\mathbb{R}^{N}$. A straightforward calculation using the definition of $S$ gives for all $i,k\in[N]$ that
$\partial_{k}S_{ii}(z)=-S_{ik}(z)S_{ki}(z).$ (4.3)
We now derive some gradient bounds that are necessary to apply Lemma 4.2. These bounds are uniform in the matrix $M$, and hence may be applied conditionally on $A_{N}$.
Lemma 4.3.
For all $s\geq 1$,
$\|\nabla F_{s}(z)\|_{2}\leq\frac{s}{\eta^{s+1}\sqrt{N}}.$
Proof.
By the chain rule,
$\partial_{k}F_{s}(z)=\frac{1}{N}\sum_{i=1}^{N}s\big(\operatorname{Im}S_{ii}(z)\big)^{s-1}\partial_{k}\big(\operatorname{Im}S_{ii}(z)\big).$
Using (4.3) and the bound $\operatorname{Im}S_{ii}(z)\leq\eta^{-1}$ from (3.1),
$\big|\partial_{k}F_{s}(z)\big|\leq\frac{s}{N}\sum_{i=1}^{N}(\operatorname{Im}S_{ii})^{s-1}|S_{ik}(z)|^{2}\leq\frac{s}{N}\cdot\eta^{-(s-1)}\sum_{i=1}^{N}|S_{ik}(z)|^{2}.$
By the Ward identity (3.2),
$\sum_{i=1}^{N}|G_{ik}(z)|^{2}=\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
Therefore,
$\big|\partial_{k}F_{s}(z)\big|\leq\frac{s}{N}\cdot\eta^{-(s-1)}\cdot\eta^{-2}=\frac{s}{N\eta^{s+1}}.$
Finally,
$\|\nabla F_{s}(z)\|_{2}^{2}=\sum_{k=1}^{N}\big|\partial_{k}F_{s}(z)\big|^{2}\leq N\Big(\frac{s}{N\eta^{s+1}}\Big)^{2}=\frac{s^{2}}{N\eta^{2s+2}},$
which implies the claim. ∎
For every $s\in(0,1/2)$, the map $x\mapsto x^{s}$ is not Lipschitz near $x=0$. We therefore introduce a truncation at level $\delta>0$ and define, for any $\delta\in(0,\eta^{-1}]$,
$F_{s,\delta}(z)=\frac{1}{N}\sum_{i=1}^{N}\big(\operatorname{Im}S_{ii}(z)\vee\delta\big)^{s},$
and define $\tilde{F}_{s,\delta}$ analogously to $\tilde{F}_{s}$.
Lemma 4.4.
Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then
$\|\nabla F_{s,\delta}(z)\|_{2}\leq\frac{s}{\eta^{2}\sqrt{N}}\ \delta^{s-1}.$
Proof.
Let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$. By the chain rule and (4.3),
$\partial_{k}F_{s,\delta}(z)=\frac{1}{N}\sum_{i=1}^{N}\varphi_{\delta}^{\prime}\big(\operatorname{Im}S_{ii}(z)\big)\partial_{k}\big(\operatorname{Im}S_{ii}(z)\big),\qquad\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{1}{N}\sum_{i=1}^{N}\|\varphi_{\delta}^{\prime}\|_{\infty}\big|\partial_{k}\operatorname{Im}S_{ii}(z)\big|.$
Note that (4.3) implies $|\partial_{k}(\operatorname{Im}S_{ii})|\leq|S_{ik}(z)|^{2}$. Then combining this estimate with (3.1) and (3.2), we find
$\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}|S_{ik}(z)|^{2}\leq\frac{s\delta^{s-1}}{N}\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{s\delta^{s-1}}{N\eta^{2}}.$
Summing over $k$ gives
$\|\nabla F_{s,\delta}(z)\|_{2}^{2}\leq N\left(\frac{s}{N\eta^{2}}\delta^{s-1}\right)^{2}=\frac{s^{2}}{N\eta^{4}}\delta^{2s-2}.$
Taking square roots yields the claim. ∎
4.3. Dependence on the Graph
We use the pairing (configuration) model for random $d$-regular graphs. Let
$W=[N]\times[d],\qquad m=\frac{Nd}{2}.$
A pairing of $W$ is a partition of $W$ into unordered sets of cardinality two. Given a pairing $\omega$ of $W$, we define the associated multigraph $\mathcal{G}(\omega)$ on the vertex set $[N]$ as follows. Each element $(i,a)\in W$ is called a half-edge attached to the vertex $i$. For every pair $\{(i,a),(j,b)\}\in\omega$ we place one edge between the vertices $i$ and $j$. Thus, if $i\neq j$, the pair $\{(i,a),(j,b)\}$ contributes one edge joining $i$ and $j$, while if $i=j$, the pair $\{(i,a),(i,b)\}$ contributes one loop at the vertex $i$. Distinct pairs of $\omega$ may project to the same unordered pair $\{i,j\}$, in which case this produces multiple edges.
We write $A(\omega)$ for the adjacency matrix of $A(\omega)$; if a loop occurs, it contributes $2$ to the corresponding diagonal entry, so that every row sum is equal to $d$. Let $\widehat{A}_{N}=A(\omega)$ for $\omega$ chosen uniformly from the set of all pairings of $W$, and let
$\mathcal{S}_{N}=\{\widehat{A}_{N}\text{ is simple}\}.$
Every simple $d$-regular graph on $[N]$ corresponds to exactly $(d!)^{N}$ pairings, hence
$\mathrm{Law}(\widehat{A}_{N}\mid\mathcal{S}_{N})=\mathrm{Law}(A_{N}),$ (4.4)
where $A_{N}$ denotes the adjacency matrix of the uniform simple $d$-regular graph on $[N]$. Moreover, for fixed $d$, it is standard that
$\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{S}_{N})=\exp\left(\frac{1-d^{2}}{4}\right)$
(see, e.g., [35, equation (5)]), and therefore there exists a constant $p_{d}>0$ such that
$\mathbb{P}(\mathcal{S}_{N})\geq p_{d}\qquad\text{for all }N.$ (4.5)
We also record the Azuma–Hoeffding martingale concentration estimate that we use. See, e.g., [19, Chapter 2].
Lemma 4.5.
Let $(M_{k},\mathcal{F}_{k})_{k=0}^{m}$ be a martingale such that
$|M_{k}-M_{k-1}|\leq c_{k}\qquad\text{almost surely for all }k\in[m].$
Then for every $u>0$,
$\mathbb{P}\Big(|M_{m}-M_{0}|\geq u\Big)\leq 2\exp\Big(-\frac{u^{2}}{2\sum_{k=1}^{m}c_{k}^{2}}\Big).$
To apply the previous concentration bound, we require the following lemma on edge switchings of $A$.
Definition 4.6.
Let $\omega$ be a pairing of a finite set of half-edges $W$. Suppose $\{a,b\},\{c,d\}\in\omega$ are two distinct pairs. A switching of the two pairs $\{a,b\}$ and $\{c,d\}$ is the operation of removing these two pairs and replacing them by one of the two alternative pairings of the four half-edges, namely $\{a,c\},\{b,d\}$ or $\{a,d\},\{b,c\}$.
Lemma 4.7.
Let $A$ and $A^{\prime}$ be two symmetric adjacency matrices arising from $d$-regular pairings, and suppose that $A$ and $A^{\prime}$ differ by a switching of two pairs of half-edges. Let
$S=(-tA+\mathrm{diag}(v)-z)^{-1},\qquad S^{\prime}=(-tA^{\prime}+\mathrm{diag}(v)-z)^{-1}.$
Then the following hold.
  1. (1)
    For every $s\geq 1$,
    $\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{8s|t|}{N\eta^{s+1}}.$
  2. (2)
    For every $s\in(0,1/2)$ and every $\delta\in(0,\eta^{-1}]$,
    $\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
Proof.
Set $\Delta=-t(A^{\prime}-A)$. Since $A$ and $A^{\prime}$ differ by one switching of two pairs, the matrix $A^{\prime}-A$ has at most eight nonzero entries, each of magnitude at most $1$. In particular,
$\mathrm{rank}(\Delta)\leq 4,\qquad\|\Delta\|\leq 2|t|.$
By the resolvent identity,
$S^{\prime}-S=S\Delta S^{\prime}.$
Hence
$\mathrm{rank}(S^{\prime}-S)\leq 4,\qquad\|S^{\prime}-S\|\leq\|S\|\,\|\Delta\|\,\|S^{\prime}\|\leq\frac{2|t|}{\eta^{2}}.$
Therefore
$\frac{1}{N}\sum_{i=1}^{N}\big|S^{\prime}_{ii}-S_{ii}\big|\leq\frac{1}{N}\|S^{\prime}-S\|_{*}\leq\frac{4}{N}\|S^{\prime}-S\|\leq\frac{8|t|}{N\eta^{2}},$
where $\|\cdot\|_{*}$ denotes the trace norm.
For $s\geq 1$, the map $x\mapsto x^{s}$ is $s\eta^{-(s-1)}$-Lipschitz on $0,\eta^{-1}]$, so
$\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{s}{N}\eta^{-(s-1)}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{s+1}}.$
This proves the first claim.
Next, let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$, and therefore
$\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
This proves the second claim. ∎
The following is the graph concentration estimate used in the proof of Lemma 4.1.
Lemma 4.8.
Fix $d\geq 3$. There exists a constant $C(d)>0$ such that the following holds. Let $f$ be a real-valued function of the pairing-model adjacency matrix $A$ satisfying
$\big|f(A)-f(A^{\prime})\big|\leq L$
whenever $A$ and $A^{\prime}$ arise from two pairings that differ by one switching of two pairs. Then for every $u>0$,
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)\leq 2\exp\left(-\frac{u^{2}}{CNdL^{2}}\right).$
Proof.
We first prove the corresponding estimate for the configuration-model adjacency matrix $\widehat{A}_{N}$.
Choose the random pairing sequentially as follows. Order the half-edges in $W$ lexicographically. At step $k$, let $a_{k}$ be the smallest currently unmatched half-edge, and choose its partner $Y_{k}$ uniformly from the remaining unmatched half-edges other than $a_{k}$. This produces a uniformly random pairing. Let
$\mathcal{F}_{k}=\sigma(Y_{1},\dots,Y_{k}),\qquad M_{k}=\mathbb{E}\big[f(\widehat{A}_{N})\mid\mathcal{F}_{k}\big],\qquad k=0,1,\dots,m.$
Then $(M_{k},\mathcal{F}_{k})$ is the Doob martingale associated with $f(\widehat{A}_{N})$.
We claim that
$|M_{k}-M_{k-1}|\leq L\qquad\text{a.s. for all }k\in[m].$ (4.6)
Fix $k$ and condition on $\mathcal{F}_{k-1}$. Let $U$ denote the set of unmatched half-edges at time k1k-1, and write $a=a_{k}$. For each admissible choice $y\in U\setminus\{a\}$, define
$g(y)=\mathbb{E}\big[f(\widehat{A}_{N})\mid\mathcal{F}_{k-1},\,Y_{k}=y\big].$
We show that for any two admissible choices $y,y^{\prime}\in U\setminus\{a\}$,
$|g(y)-g(y^{\prime})|\leq L.$ (4.7)
If $y=y^{\prime}$, there is nothing to prove. Assume henceforth that $y\neq y^{\prime}$. Since we condition on $\mathcal{F}_{k-1}$, the unmatched set $U$ and the half-edge $a=a_{k}$ are deterministic. Under the conditional law given $\mathcal{F}_{k-1}$ and $Y_{k}=y$, the remaining pairs form a uniform random pairing of $U\setminus\{a,y\}$. Thus
$g(y)=\mathbb{E}_{R}\Big[f\big(A(\{\{a,y\}\}\cup R)\big)\Big],$
where $R$ is uniform over pairings of $U\setminus\{a,y\}$. Likewise,
$g(y^{\prime})=\mathbb{E}_{R^{\prime}}\Big[f\big(A(\{\{a,y^{\prime}\}\}\cup R^{\prime})\big)\Big],$
where $R^{\prime}$ is uniform over pairings of $U\setminus\{a,y^{\prime}\}$.
Now let $R$ be a uniformly random pairing of $U\setminus\{a,y\}$. Let $x$ denote the partner of $y^{\prime}$ in $R$, and define $\Phi(R)$ to be the pairing of $U\setminus\{a,y^{\prime}\}$ obtained from $R$ by replacing the pair $\{y^{\prime},x\}$ with $\{y,x\}$ and leaving all other pairs unchanged. The map $R\mapsto\Phi(R)$ is plainly a bijection from the set of pairings of $U\setminus\{a,y\}$ to the set of pairings of $U\setminus\{a,y^{\prime}\}$. Moreover, the two full pairings $\{\{a,y\}\}\cup R$ and $\{\{a,y^{\prime}\}\}\cup\Phi(R)$ differ by exactly one switching of two pairs. Therefore, by the assumed Lipschitz bound on $f$,
$\Big|f\big(A(\{\{a,y\}\}\cup R)\big)-f\big(A(\{\{a,y^{\prime}\}\}\cup\Phi(R))\big)\Big|\leq L.$
Averaging over $R$ yields (4.7).
Since $M_{k}=g(Y_{k})$ and $M_{k-1}=\mathbb{E}[g(Y_{k})\mid\mathcal{F}_{k-1}]$, the latter being an average of the values $g(y)$ over admissible $y$, (4.7) implies (4.6). Hence Lemma 4.5 gives
$\mathbb{P}\Big(\big|f(\widehat{A}_{N})-\mathbb{E}[f(\widehat{A}_{N})]\big|\geq u\Big)\leq 2\exp\Big(-\frac{u^{2}}{NdL^{2}}\Big),$ (4.8)
since $m=Nd/2$. We also record the variance bound
$\operatorname{Var}\big(f(\widehat{A}_{N})\big)=\sum_{k=1}^{m}\mathbb{E}\big[(M_{k}-M_{k-1})^{2}\big]\leq mL^{2}\leq\frac{NdL^{2}}{2}.$ (4.9)
We now pass from the pairing model to the uniform simple graph. Let
$\mu=\mathbb{E}[f(\widehat{A}_{N})],\qquad\mu_{\mathrm{s}}=\mathbb{E}\big[f(\widehat{A}_{N})\mid\mathcal{S}_{N}\big].$
Since
$\mathbb{P}(\mathcal{S}_{N})\big(\mu_{\mathrm{s}}-\mu\big)=\mathbb{E}\Big[\big(f(\widehat{A}_{N})-\mu\big)\mathbf{1}_{\mathcal{S}_{N}}\Big],$
the Cauchy–Schwarz inequality, (4.5), and (4.9) imply
$|\mu_{\mathrm{s}}-\mu|\leq\frac{\sqrt{\operatorname{Var}\big(f(\widehat{A}_{N})\big)}}{\sqrt{\mathbb{P}(\mathcal{S}_{N})}}\leq C_{1}(d)\sqrt{Nd}\,L$ (4.10)
for some constant $C_{1}(d)<\infty$.
Next, using (4.4), we have for every $u>0$,
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)=\mathbb{P}\Big(\big|f(\widehat{A}_{N})-\mu_{\mathrm{s}}\big|\geq u\,\Big|\,\mathcal{S}_{N}\Big).$
If $u>2C_{1}(d)\sqrt{Nd}\,L$, then by (4.10),
$\Big\{\big|f(\widehat{A}_{N})-\mu_{\mathrm{s}}\big|\geq u\Big\}\subset\Big\{\big|f(\widehat{A}_{N})-\mu\big|\geq\frac{u}{2}\Big\},$
and therefore
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}[f(A_{N})]\big|\geq u\Big)\leq\frac{1}{\mathbb{P}(\mathcal{S}_{N})}\mathbb{P}\Big(\big|f(\widehat{A}_{N})-\mu\big|\geq\frac{u}{2}\Big).$
Using (4.5) and (4.8), we obtain
$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}[f(A_{N})]\big|\geq u\Big)\leq 2p_{d}^{-1}\exp\Big(-\frac{u^{2}}{4NdL^{2}}\Big).$
Since $p_{d}>0$ depends only on $d$, this is bounded by
$2\exp\Big(-\frac{u^{2}}{C_{2}(d)NdL^{2}}\Big)$
for some constant $C_{2}(d)<\infty$.
Finally, if $u\leq 2C_{1}(d)\sqrt{Nd}\,L$, then the claimed bound is trivial after enlarging the constant once more, since the right-hand side can be made at least $1$ on this range of $u$. This proves the claim. ∎
4.4. Conclusion
We begin by deriving two preliminary estimates for the truncated function $\tilde{F}_{s,\delta}$.
Lemma 4.9.
Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$ (4.11)
Proof.
Conditioning on $A_{N}$ and using Lemmas 4.2 and 4.4, we obtain
$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\mid A_{N}\big]\big|\geq u/2\,\Big|\,A_{N}\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big).$
Averaging over $A_{N}$ yields
$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\mid A_{N}\big]\big|\geq u/2\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big).$
Now define
$g_{s,\delta}(A)=\mathbb{E}_{V}\big[F_{s,\delta}(z,V_{N},-tA)\big],$
where $\mathbb{E}_{V}$ denotes expectation in the Gaussian potential only. By Lemma 4.7, $g_{s,\delta}$ is $L$-Lipschitz under one simple switching with
$L=\frac{8s|t|}{N\eta^{2}}\delta^{s-1}.$
Hence Lemma 4.8 implies
$\mathbb{P}\Big(\big|\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\mid A_{N}\big]-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u/2\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$
Combining the last two estimates by the triangle inequality yields (4.11). ∎
Lemma 4.10.
Let $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
$\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\big|\geq u+2\delta^{s}\Big)\leq\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big).$
Proof.
For any $x\geq 0$, we have $(x\vee\delta)^{s}\geq x^{s}$, hence $F_{s,\delta}\geq F_{s}$. Moreover, if $x\geq\delta$ then $(x\vee\delta)^{s}-x^{s}=0$, while if $x<\delta$ then $(x\vee\delta)^{s}-x^{s}=\delta^{s}-x^{s}\leq\delta^{s}$. Therefore for all $x\geq 0$,
$0\leq(x\vee\delta)^{s}-x^{s}\leq\delta^{s}.$
Averaging over the $S_{ii}$ entries yields $0\leq\tilde{F}_{s,\delta}-\tilde{F}_{s}\leq\delta^{s}$, and taking expectations gives
$0\leq\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\leq\delta^{s}.$
Finally, by the triangle inequality,
$\big|\tilde{F}_{s}-\mathbb{E}[\tilde{F}_{s}]\big|\leq\big|\tilde{F}_{s,\delta}-\mathbb{E}[\tilde{F}_{s,\delta}]\big|+|\tilde{F}_{s,\delta}-\tilde{F}_{s}|+\big|\mathbb{E}[\tilde{F}_{s,\delta}]-\mathbb{E}[\tilde{F}_{s}]\big|\leq\big|\tilde{F}_{s,\delta}-\mathbb{E}[\tilde{F}_{s,\delta}]\big|+2\delta^{s}.$
Hence the event $\{|\tilde{F}_{s}-\mathbb{E}[\tilde{F}_{s}]|\geq u+2\delta^{s}\}$ is contained in $\{|\tilde{F}_{s,\delta}-\mathbb{E}[\tilde{F}_{s,\delta}]|\geq u\}$, proving the claim. ∎
We now prove Lemma 4.1.
Proof of Lemma 4.1.
We begin with the case $s\geq 1$. Conditioning on $A_{N}$ and using Lemmas 4.2 and 4.3, we obtain
$\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\mid A_{N}\big]\big|\geq u/2\,\Big|\,A_{N}\Big)\leq 2\exp\Big(-\frac{N\eta^{2s+2}}{8s^{2}}u^{2}\Big).$
Averaging over $A_{N}$ yields
$\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\mid A_{N}\big]\big|\geq u/2\Big)\leq 2\exp\Big(-\frac{N\eta^{2s+2}}{8s^{2}}u^{2}\Big).$
Now define
$g_{s}(A)=\mathbb{E}_{V}\big[F_{s}(z,V_{N},-tA)\big].$
By Lemma 4.7, $g_{s}$ is $L$-Lipschitz under one simple switching with
$L=\frac{8s|t|}{N\eta^{s+1}}.$
Hence Lemma 4.8 implies
$\mathbb{P}\Big(\big|\mathbb{E}\big[\tilde{F}_{s}(z)\mid A_{N}\big]-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\big|\geq u/2\Big)\leq 2\exp\Big(-\frac{N\eta^{2s+2}}{Cdt^{2}s^{2}}u^{2}\Big).$
Combining the last two bounds by the triangle inequality proves (4.1). The claim about the corresponding averages involving $\operatorname{Re}S_{ii}(z)$ is completely analogous.
Next, by Lemma 4.10 and Lemma 4.9, for any $s\in(0,1/2)$, $\delta\in(0,\eta^{-1}]$, and $u>0$,
$\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\tilde{F}_{s}(z)\big|\geq u+2\delta^{s}\Big)\leq\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\tilde{F}_{s,\delta}(z)\big|\geq u\Big)$
$\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big),$
which is (4.2). ∎
5. Bounds for the Eigenvalue Counting Function
This section is dedicated to the proof of Lemma 3.4. To obtain this result, we first show that the Stieltjes transform of $W_{N}$ converges to the Stieltjes transform of the density of the local states measure of $H_{d,t}$ at the root of the Bethe lattice. For the Anderson model on the random regular graph with a bounded potential, an analogous result for the integrated density of states was proved by Geisinger [23]. We adapt this result to $W_{N}$ by a truncation argument, then upgrade the convergence in means to convergence in probability using the $s=1$ case of Lemma 4.1. Finally, since convergence of Stieltjes transforms implies convergence of the associated spectral measures, the desired lower bound on the eigenvalue counting bound follows from a lower bound on the integrated density of states of $H_{d,t}$, which we show in Appendix B, and the upper bound follows from Wegner’s estimate.
5.1. Notation
For all $x\in[N]$, we define the spectral measure of $W_{N}$ at $x$ by
$\mu_{N,x}(B)=\langle\delta_{x},\mathbbm{1}_{B}(W_{N})\delta_{x}\rangle,$
where $B\subset\mathbb{R}$ is any Borel set, and $\mathbbm{1}_{B}(W_{N})$ is defined in the usual way (using the spectral theorem). We define
$m_{N}(z)=\frac{1}{N}\operatorname{Tr}G_{N}(z),$
which is the Stieltjes transform of the empirical eigenvalue distribution of $W_{N}$.
Next, bcause $H_{d,t}$ is self-adjoint, the spectral theorem implies the existence of a random projection-valued measure $\mu$ such that $H=\int\lambda\,d\mu(\lambda)$. The spectral measure at the root is given by
$\mu_{0}(S)=\langle\delta_{0},\mu(S)\delta_{0}\rangle$
for all Borel sets $S\subset\mathbb{R}$, and the density of states measure $\nu$ is given by
$\nu(S)=\mathbb{E}\big[\mu_{0}(S)\big].$
For all $z\in\mathbb{H}$, the Stieltjes transform of $\nu$ is
$m_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,d\nu(\lambda).$
5.2. Preliminaries
We first record an integration by parts bound for signed distribution functions.
Lemma 5.1.
Let $\mu_{1}$ and $\mu_{2}$ be finite Borel measures on $\mathbb{R}$, and set
$F(t)=\mu_{1}\big((-\infty,t]\big)-\mu_{2}\big((-\infty,t]\big).$
If $f\in C^{1}(\mathbb{R})$ satisfies $f(\lambda)\to 0$ as $|\lambda|\to\infty$ and $f^{\prime}\in L^{1}(\mathbb{R})$, then
$\Big|\int_{\mathbb{R}}f\,d\mu_{1}-\int_{\mathbb{R}}f\,d\mu_{2}\Big|\leq\|F\|_{\infty}\int_{\mathbb{R}}\big|f^{\prime}(\lambda)\big|\,d\lambda.$
Proof.
We apply integration by parts to the signed measure $\mu_{1}-\mu_{2}$ (see, e.g., [21, Theorem 3.36]), yielding
$\int_{\mathbb{R}}f\,d(\mu_{1}-\mu_{2})=-\int_{\mathbb{R}}F(\lambda)f^{\prime}(\lambda)\,d\lambda,$
then bound the right-hand side by by $\|F\|_{\infty}\int|f^{\prime}|$. ∎
Next we establish an estimate on the stability of resolvents under perturbations.
Lemma 5.2.
Let $H$ and $\widetilde{H}$ be symmetric $N\times N$ matrices, and set $\Delta=\widetilde{H}-H$. Then for any $z\in\mathbb{H}$,
$\Big|\frac{1}{N}\operatorname{Tr}(\widetilde{H}-z)^{-1}-\frac{1}{N}\operatorname{Tr}(H-z)^{-1}\Big|\leq\frac{1}{N\eta^{2}}\operatorname{Tr}|\Delta|.$ (5.1)
Proof.
By the definition of a resolvent,
$(\widetilde{H}-z)^{-1}-(H-z)^{-1}=-(H-z)^{-1}\Delta(\widetilde{H}-z)^{-1}.$
Taking traces and using the trace norm bound $|\operatorname{Tr}(ABC)|\leq\|A\|\|C\|\operatorname{Tr}|B|$ yields
$\Big|\operatorname{Tr}\big((\widetilde{H}-z)^{-1}-(H-z)^{-1}\big)\Big|\leq\big\|(H-z)^{-1}\big\|\big\|(\widetilde{H}-z)^{-1}\big\|\operatorname{Tr}|\Delta|\leq\eta^{-2}\operatorname{Tr}|\Delta|.$
Dividing by $N$ yields the claim. ∎
Given a probability density $\tilde{\rho}$, we let $\widetilde{W}_{N}=-tA_{N}+\widetilde{V}_{N}$ and $H_{d,t}=-tA^{(d)}+\widetilde{V}^{(d)}$, where $\widetilde{V}_{N}$ and $\widetilde{V}^{(d)}$ are defined analogously to $V_{N}$ and $V^{(d)}$ with the Gaussian potential variables replaced by i.i.d. variables with density $\tilde{\rho}$. We set $\tilde{\mu}_{N,x}$ and $\tilde{\mu}_{0}$ denote the corresponding spectral measures, defined analogously to $\mu_{N,x}$ and $\mu_{0}$.
The following theorem is [23, Theorem 2.3].
Theorem 5.3.
Suppose that $\|\tilde{\rho}\|$$\infty$ and there exists $\rho_{0}>0$ such that $\mathrm{supp}(\tilde{\rho})\subset(-\rho_{0},\rho_{0})$. Let $\widetilde{\mathbb{E}}$ denote expectation with respect to the potentials $\tilde{V}_{N}$ and $\tilde{V}^{(d)}$, conditional on $A_{N}$. Define the event
$\mathcal{A}_{N}=\left\{\sup_{s\in\mathbb{R}}\Bigg|\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}\big[\tilde{\mu}_{N,x}\big((-\infty,s]\big)\big]-\widetilde{\mathbb{E}}\big[\tilde{\mu}_{0}\big((-\infty,s]\big)\big]\Bigg|\leq\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.2)
which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{A}_{N})=1$.
Define the disorder-averaged measures
$\tilde{\nu}^{*}_{N}=\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}[\tilde{\mu}_{N,x}],\qquad\tilde{\nu}^{*}=\widetilde{\mathbb{E}}[\tilde{\mu}_{0}].$
The corresponding Stieltjes measures are
$m^{*}_{N}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}_{N}(d\lambda),\qquad m^{*}_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}(d\lambda).$
Using Theorem 5.3, we deduce convergence of the disorder-averaged Stieltjes transform in the bounded potential setting.
Proposition 5.4.
Retain the hypotheses and notation of Theorem 5.3. For every $z\in\mathbb{H}$, define the event
$\mathcal{B}_{N}(z)=\left\{\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|\leq\frac{\pi}{\eta}\cdot\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.3)
which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N}(z))=1$.
Proof.
Define $f_{z}(\lambda)=(\lambda-z)^{-1}$. Then $f^{\prime}_{z}(\lambda)=-(\lambda-z)^{-2}$ and
$\int_{\mathbb{R}}|f^{\prime}_{z}(\lambda)|\,d\lambda=\int_{\mathbb{R}}\frac{d\lambda}{(\lambda-E)^{2}+\eta^{2}}=\frac{\pi}{\eta}.$ (5.4)
Apply Lemma 5.1 with $\mu_{1}=\tilde{\nu}^{*}_{N}$, $\mu_{2}=\tilde{\nu}^{*}$, and $f=f_{z}$. Further, define
$F_{N}(s)=\tilde{\nu}^{*}_{N}\big((-\infty,s]\big)\qquad F_{\mathbb{T}}(s)=\tilde{\nu}^{*}\big((-\infty,s]\big).$
Then, using (5.4),
$\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|=\Big|\int f_{z}\,d\tilde{\nu}^{*}_{N}-\int f_{z}\,d\tilde{\nu}^{*}\Big|\leq\|F_{N}-F_{\mathbb{T}}\|_{\infty}\int|f^{\prime}_{z}|=\|F_{N}-F_{\mathbb{T}}\|_{\infty}\cdot\frac{\pi}{\eta}.$
Now insert the bound (5.2) to complete the proof. ∎
Next, we state an upper bound on the integrated density of states of $H_{d,t}$, which is a consequence of a well-known estimate of Wegner [8, (5.27)].
Lemma 5.5.
Fix $g,\mathfrak{L}>0$ such that (2.1) holds. For every $d\geq 3$, there exists a constant $C(d,\mathfrak{L})>0$ such that for all intervals $I\subset\mathbb{R}$, we have $\nu(I)\leq C|I|$.
Finally, we state a lower bound that complements the previous Wegner estimate. It is proved in Appendix B.
Lemma 5.6.
Fix $g,\mathfrak{L}>0$ such that (2.1) holds, and fix $D>0$. For every $d\geq 3$, there exists a constant $c(d,D,\mathfrak{L})>0$ such that for all intervals $I\subset[-D,D]$, we have $\nu(I)\geq c|I|$.
5.3. Global Law for Gaussian Disorder
To pass from Gaussian disorder to bounded support, we introduce a coupled truncation scheme. Fix $L\geq 1$. Let $X\sim N(0,1)$ and let $U^{(L)}\sim\mathrm{Unif}[-L,L]$ be independent of $V$. Define
$X^{(L)}=X\mathbbm{1}_{\{|X|\leq L\}}+U^{(L)}\mathbbm{1}_{\{|X|>L\}}.$ (5.5)
Let $(X_{x},U_{x}^{(L)})_{x\in[N]}$ be i.i.d. with law $(X,U)$, define $X_{x}^{(L)}$ by (5.5), and set
$V_{N}^{(L)}=\mathrm{diag}(X_{1}^{(L)},\dots,X_{N}^{(L)}),\qquad W_{N}^{(L)}=-tA_{N}+V_{N}^{(L)}.$
Define the truncated Stieltjes transform
$m_{N}^{(L)}(z)=\frac{1}{N}\operatorname{Tr}(W_{N}^{(L)}-z)^{-1},$
and define similarly $H_{t,d}^{(L)}=-tA^{(d)}+V^{(d),(L)}$ on the Bethe lattice and
$m_{\mathbb{T}}^{(L)}(z)=\mathbb{E}\big[\langle\delta_{0},(H_{d,t}^{(L)}-z)^{-1}\delta_{0}\rangle\big].$
Proposition 5.7.
Fix $z\in\mathbb{H}$. Then $\lim_{N\rightarrow\infty}m_{N}(z)=m_{\mathbb{T}}(z)$ in probability.
Proof.
We proceed through a sequence of reductions that isolate truncation and approximation errors. We first apply Lemma 5.2 with $H=W_{N}$ and $\widetilde{H}=W_{N}^{(L)}$. Here $\Delta$ is diagonal with entries $V_{N,i}^{(L)}-V_{N,i}$, hence
$|m_{N}(z)-m_{N}^{(L)}(z)|\leq\frac{1}{\eta^{2}}\cdot\frac{1}{N}\sum_{i=1}^{N}|V_{N,i}^{(L)}-V_{N,i}|.$
Taking expectation and using that the entries of each potential are i.i.d., we find
$\sup_{N\geq 1}\mathbb{E}\Big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\Big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[|V_{N,1}^{(L)}-V_{N,1}|\big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[\big(|g_{1}|+L\big)\mathbbm{1}_{\{|g_{1}|>L\}}\big].$ (5.6)
For each $L$ and $i\in[N]$, the law of $V^{(L)}_{N,i}$ is supported on [L,L][-L,L] and has a bounded density. Therefore Proposition 5.4 applies to $W_{N}^{(L)}$ and yields, for fixed $L$ and fixed $z$,
$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}^{(L)}(z)\big]=m_{\mathbb{T}}^{(L)}(z).$ (5.7)
We now remove the truncation on the tree. Because the tree has infinitely many sites, we do not compare $m_{\mathbb{T}}^{(L)}$ and $m_{\mathbb{T}}$ by summing over sites of $\mathbb{T}_{d}$. Instead, we use strong resolvent convergence. Let $C_{c}(\mathbb{T})$ denote finitely supported functions on $T_{d}$. It is well known that $C_{c}(\mathbb{T})$ is a core for the multiplication operator $V^{(d)}$ and hence for $H_{d,t}=-tA^{(d)}+V^{(d)}$ (since $tA^{(d)}$ is bounded on $\ell^{2}(\mathbb{T})$). Moreover, for each $\psi\in C_{c}(\mathbb{T})$, we have $H_{d,t}^{(L)}\psi\to H_{d,t}\psi$ a.s. as $L\to\infty$ because $\psi$ has finite support and $V_{x}^{(L)}\to V_{x}$ pointwise. By the core convergence criterion for strong resolvent convergence (see, e.g., [31, Theorem VIII.25]), $(H_{d,t}^{(L)}-z)^{-1}$ converges in the strong operator topology to $(H_{d,t}-z)^{-1}$. Applying this fact to $\delta_{0}$ and using the uniform bound $\|(H_{d,t}^{(L)}-z)^{-1}\|\leq\eta^{-1}$ together with the dominated convergence theorem yields
$\lim_{L\rightarrow\infty}m_{\mathbb{T}}^{(L)}(z)=m_{\mathbb{T}}(z).$ (5.8)
Finally, we combine the previous estimates to conclude convergence in mean, then upgrade to convergence in probability. Fix $\varepsilon>0$. Choose $L$ large enough so that (5.6) and (5.8) yield
$\sup_{n}\mathbb{E}\big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\big]<\varepsilon/3,\qquad|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)|<\varepsilon/3.$
Then take $N$ large enough so that (5.7) gives $|\mathbb{E}m_{N}^{(L)}(z)-m_{\mathbb{T}}^{(L)}(z)|<\varepsilon/3$. By the triangle inequality,
$\big|\mathbb{E}\big[m_{N}(z)\big]-m_{\mathbb{T}}(z)\big|\leq\big|\mathbb{E}\big[m_{N}(z)\big]-\mathbb{E}\big[m_{N}^{(L)}(z)\big]\big|+\big|\mathbb{E}\big[m_{N}^{(L)}(z)\big]-m_{\mathbb{T}}^{(L)}(z)\big|+\big|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)\big|<\varepsilon.$
Hence
$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}(z)\big]=m_{\mathbb{T}}(z)$ (5.9)
in probability. By the $s=1$ case of Lemma 3.3, there exists $C(d)>0$ such that for all $u\geq 0$,
$\mathbb{P}\big(\big|m_{N}(z)-\mathbb{E}m_{N}(z)\big|\geq u\big)\leq C\exp\big(-N\eta^{4}u^{2}/C\big).$ (5.10)
Together with (5.9), this completes the proof. ∎
5.4. Conclusion
Proof of Lemma 3.4.
Let
$\widehat{\mu}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$
denote the empirical measure of $W_{N}$. By a basic property of Stieltjes transforms, the convergence of Stieltjes transforms in probability at every $z\in\mathbb{H}$ provided by Proposition 5.7 implies that $\widehat{\mu}_{N}$ converges weakly to $\nu$ in probability (see, e.g., [14, Lemma 2.1]). Since $\nu$ is absolutely continuous, by Lemma 5.5, we have that $\widehat{\mu}_{N}(I)$ converges to $\nu(I)$ in probability for every interval $I\subset\mathbb{R}$. The conclusion now follows from Lemma 5.5 and Lemma 5.6. ∎
6. Diagonal Resolvent Convergence
This section proves Lemma 3.5, which concerns convergence of the diagonal resolvent entries of $W_{N}$ to $R_{00}$. The two main technical ingredients of the proof are (i) a Combes–Thomas-type exponential decay bound for resolvents entries at fixed $\eta>0$ and (ii) the fact that a random $d$-regular graph is locally tree-like with high probability. Using these inputs, we prove Lemma 3.5 by truncating the $d$-regular graph around a fixed vertex and then controlling the differences of the resolvent entries of the original and truncated graph by the Combes–Thomas bound.
6.1. Preliminary Bounds
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be any (finite or infinite) graph with maximum degree at most $d$. Here $\mathcal{V}$ denotes the vertex set and $\mathcal{E}$ denotes the edge set. Let $A$ denote the adjacency operator on $\ell^{2}(\mathcal{V})$, let $V$ also denote a real diagonal potential (not necessarily bounded), fix $t>0$, and define H=tA+VH=-tA+V. Since this notation conflicts with the shorthand introduced in in Section 2.1 for the Anderson model on the Bethe lattice, we will always use the more precise notation $H_{d,t}=-tA^{(d)}+V^{(d)}$ when discussing the Bethe lattice in this section.
For all $X\subset\mathcal{V}$, write $\chi_{X}$ for the orthogonal projection onto $\ell^{2}(X)$. For all $X,Y\subset\mathcal{V}$ define the set-distance
$\operatorname{dist}(X,Y)=\inf\{\operatorname{dist}(x,y):x\in X,\ y\in Y\}\in\{0,1,2,\dots\}\cup\{\infty\}.$
We record the following straightforward observation, omitting the proof.
Lemma 6.1.
If $\mathcal{G}$ is a graph with maximum degree at most $d$, then $\|A\|\leq d$ on $\ell^{2}(\mathcal{V})$.
The following Combes–Thomas bound is a key technical input for this section. Its proof is given in Appendix A.
Lemma 6.2.
Let $\mathcal{G}$ be a graph with maximum degree at most $d$. Fix $z=E+\mathrm{i}\eta$ with $\eta>0$ and set
$\mu_{\mathrm{CT}}=\ln\Big(1+\frac{\eta}{4td}\Big).$
Then for any subsets $X,Y\subset\mathcal{V}$,
$\big\|\chi_{X}(H-z)^{-1}\chi_{Y}\big\|\leq\frac{2}{\eta}\exp\big(-\mu_{\mathrm{CT}}\operatorname{dist}(X,Y)\big).$
6.2. Decoupling and Local Tree-Likeness
We next define a decoupling of the Hamiltonian and quantify its effect on the root resolvent. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be as above and fix a vertex $o\in\mathcal{V}$. For every $r\in\mathbb{N}$ define the ball
$B_{r}(o)=\{x\in\mathcal{V}:\operatorname{dist}(o,x)\leq r\}.$
Define the decoupling at radius $r$ of the Anderson model on $\mathcal{G}$ by considering the Anderson model on the graph formed by removing all edges between $B_{r}(o)$ and its complement $B_{r}(o)^{c}$, yielding an operator
$H^{(r)}=H_{B_{r}(o)}\oplus H_{B_{r}(o)^{c}}$
on $\ell^{2}(\mathcal{V})$. For this section only, we always use $G_{N}$ when referring to the resolvent of $W_{N}$; we let $G(z)=(H-z)^{-1}$ and $G^{(r)}(z)=(H^{(r)}-z)^{-1}$ and write $G_{oo}(z)=\langle\delta_{o},G(z)\delta_{o}\rangle$, and similarly for $G^{(r)}_{oo}(z)$.
Lemma 6.3.
Let $\mathcal{G}$ be a graph with maximum degree at most $d$. With $\mu_{\mathrm{CT}}$ as in Lemma 6.2, for all $r\in\mathbb{N}$,
$\big|G_{oo}(z)-G^{(r)}_{oo}(z)\big|\leq\frac{4td}{\eta^{2}}\exp(-2\mu_{\mathrm{CT}}r).$
Proof.
Let $K=H-H^{(r)}$. Then $K$ is supported on edges crossing the cut between $B_{r}(o)$ and $B_{r}(o)^{c}$, and it and equals t-t times the adjacency matrix of the graph formed by the cut edges. Hence, by Lemma 6.1, $\|K\|\leq t\|A\|\leq td$. It follows from the definitions of $G$ and $G^{(r)}$ that
$G-G^{(r)}=GKG^{(r)}.$
Let $\partial B_{r}(o)$ denote the set of vertices in $\mathcal{B}$ incident to at least one cut edge (both inside and outside $B_{r}(o)$). Then $\operatorname{dist}(\{o\},\partial B_{r}(o))=r$ and $K=\chi_{\partial B_{r}(o)}K\chi_{\partial B_{r}(o)}$. Therefore
$|G_{oo}-G^{(r)}_{oo}|=\big|\langle\delta_{o},GKG^{(r)}\delta_{o}\rangle\big|\leq\|\chi_{\{o\}}G\chi_{\partial B_{r}(o)}\|\,\|K\|\,\|\chi_{\partial B_{r}(o)}G^{(r)}\chi_{\{o\}}\|.$
By Lemma 6.2 applied to both $H$ and $H^{(r)}$, the first and last factors on the right-hand side are bounded by $\frac{2}{\eta}\exp(-\mu_{\mathrm{CT}}r)$. Hence
$|G_{oo}-G^{(r)}_{oo}|\leq\Big(\frac{2}{\eta}e^{-\mu_{\mathrm{CT}}r}\Big)^{2}\cdot td=\frac{4td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r},$
as desired. ∎
We then use local tree-likeness of random $d$-regular graphs to compare finite neighborhoods with the tree. For a rooted graph $(\mathcal{G},o)$ we write $B_{r}(\mathcal{G},o)$ for the rooted radius-$r$ ball.
Lemma 6.4.
Fix $r\in\mathbb{N}$ and let $\mathcal{G}_{N}$ be a uniformly random simple $d$-regular graph on $N$ vertices labeled $1$ to $N$. Then there exists a constant $C(d)>0$ such that for all $N\geq C$,
$\mathbb{P}\big(B_{r}(\mathcal{G}_{N},1)\text{ contains a cycle}\big)\leq\frac{C(d-1)^{2r}}{N}.$
Proof.
We briefly sketch the standard configuration-model argument.
Let $G^{\mathrm{conf}}_{N,d}$ be the random $d$-regular multigraph from the configuration model (which places $d$ half-edges on of the $N$ vertices, then forms a perfect matching of the half-edges uniformly at random), and let $\mathbb{Q}$ be the induced measure on graph events. We let $C(d)>0$ denote a $d$-dependent constant that may increase at each appearance in this proof. For fixed $d$, we have $\mathbb{Q}(G^{\mathrm{conf}}_{N,d}\text{is simple})>C^{-1}$ for all $N\geq C$ by (4.5). Therefore, by the definition of conditional probability, we have for any event $\mathcal{E}$ measurable with respect to the graph and $N\geq C$ that
$\mathbb{P}(\mathcal{E})\leq C\mathbb{Q}(\mathcal{E}).$ (6.1)
We can construct $B_{r}(1)$ by breadth-first search in the associated configuration model, pairing one half-edge at each step. Let $|B_{r}(1)|$ denote the number of vertices in $B_{r}(1)$, and let $M$ be the total number of half-edges in $B_{r}(1)$ Deterministically,
$\big|B_{r}(1)\big|\leq 1+d\sum_{k=0}^{r-1}(d-1)^{k}\leq C(d-1)^{r},$
which implies
$M\leq d\big|B_{r}(1)\big|\leq C(d-1)^{r},$
where we increase $C$ at each step if necessary. During breadth-first pairing procedure, at any step, the probability that a newly exposed half-edge pairs with one of the at most $M$ previously exposed half-edges is bounded by $CM/Nd$. A cycle in $B_{r}(1)$ can occur only if some such collision happens. By a union bound over the at most $M$ steps,
$\mathbb{Q}\big(B_{r}(1)\text{ contains a cycle in }G^{\mathrm{conf}}_{N,d}\big)\leq M\cdot\frac{CM}{Nd}\leq\frac{C(d-1)^{2r}}{N}.$
This estimate then transfers to the uniform simple $d$-regular model using (6.1). ∎
6.3. Proof of Diagonal Resolvent Convergence
We now construct the coupling between d-regular graph potentials and tree potentials used in the diagonal resolvent comparison. Sample the tree potential $(V_{u})_{u\in\mathbb{V}}$ as i.i.d. $N(0,1)$ random variables. Independently sample the graphs $(\mathcal{G}_{N})_{N\geq 1}$, each uniform over simple $d$-regular graphs on $[N]$. Let $r_{N}=\ln\ln N$, and define the event
$\mathcal{T}_{N}=\{B_{r_{N}}(\mathcal{G}_{N},1)\text{ is a tree}\}.$
Recall the shorthand $\mathbb{T}=\mathbb{T}_{d}$. On $\mathcal{T}_{N}$, choose a rooted graph isomorphism
$\varphi_{N}\colon B_{r_{N}}(\mathbb{T},0)\to B_{r_{N}}(\mathcal{G}_{N},1).$
For concreteness, one may select the isomorphism induced by the breadth-first exploration of the half-edges in the configuration model construction of $\mathcal{G}_{N}$. We define a potential $V^{(N)}$ on $\mathcal{G}_{N}$ in two parts. In the ball, we set
$V^{(N)}_{x}=V_{\varphi_{N}^{-1}(x)},\qquad x\in B_{r_{N}}(\mathcal{G}_{N},1).$
Outside the ball, meaning for indices $x\notin B_{r_{N}}$, we sample the $V^{(N)}_{x}$ variables as i.i.d. $N(0,1)$, which are additionally independent from all other variables in this construction.
Let
$H_{N}=-tA_{N}+V^{(N)}$
and note that $H_{N}$ has the same distribution as $W_{N}$. Hence, we will identify $H_{N}$ with $W_{N}$ for the remainder of this section.
Let $H_{N}^{(r)}$ be the decoupling of $H_{N}$ at radius $r$ around the root $1\in[N]$, and let $G_{N}^{(r)}(z)=(H_{N}^{(r)}-z)^{-1}$. Define $H^{(r)}$ and $R^{(r)}(z)$ similarly on $\mathbb{T}$ (decoupling at radius $r$ around the root 0). The following lemma is an immediate consequence of these definitions.
Lemma 6.5.
On the event $\mathcal{T}_{N}$ we have
$G^{(r_{N})}_{N,11}(z)=R^{(r_{N})}_{00}(z)$
under the above coupling.
Proof.
On $\mathcal{T}_{N}$, the rooted balls $B_{r_{N}}(\mathcal{G}_{N},1)$ and $B_{r_{N}}(\mathbb{T},0)$ are isomorphic via $\varphi_{N}$, and by construction the potentials match under this isomorphism. The decoupled operators $H_{N}^{(r_{N})}$ and $H^{(r_{N})}$ are direct sums of the inside-ball operator and the outside-ball operator. The diagonal resolvent entry at the root depends only on the inside-ball part. Since the inside-ball operators are identical under the isomorphism, the diagonal resolvent entries coincide. ∎
We can now prove Lemma 3.5.
Proof of Lemma 3.5.
By Lemma 6.3 applied to $H_{N}$ (root 11) and to $H$ (root 0) with $r=r_{N}$, we obtain
$\big|G_{N,11}(z)-G^{(r_{N})}_{N,11}(z)\big|\leq\frac{4td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r_{N}},\qquad\big|R_{00}(z)-R^{(r_{N})}_{00}(z)\big|\leq\frac{4td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r_{N}}.$
On the event $\mathcal{T}_{N}$, Lemma 6.5 gives $G^{(r_{N})}_{N,11}(z)=R^{(r_{N})}_{00}(z)$, hence
$\big|G_{N,11}(z)-R_{00}(z)\big|\leq\big|G_{N,11}-G^{(r_{N})}_{N,11}\big|+\big|G^{(r_{N})}_{N,11}-R^{(r_{N})}_{00}\big|+\big|R^{(r_{N})}_{00}-R_{00}\big|\leq\frac{8td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r_{N}}.$
Since $r_{N}\to\infty$, the deterministic bound on the right tends to 0. Let $\varepsilon>0$ be a parameter. Then there exists $C_{0}(\varepsilon)>0$ such that for $N>C_{0}$,
$\mathbb{P}\Big(\big|G_{N,11}(z)-R_{00}(z)\big|>\varepsilon\Big)\leq\mathbb{P}(\mathcal{T}_{N}^{c}).$
Finally, by Lemma 6.4,
$\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{T}_{N}^{c})=0.$
This proves $G_{N,11}(z)\to R_{00}(z)$ in probability under the coupling. ∎
7. Uniform Moment Bound for the Resolvent
In this section, we prove Lemma 3.6.
7.1. Preliminaries
It is well known (see, e.g., [5, (2.2)]) that for every $z\in\mathbb{H}$, the resolvent entry $R^{\prime}_{00}(z)$ for the Anderson model on a variant of $\mathbb{T}_{d}$, where the root degree is $d$ and the degree of all other vertices is $d+1$, solves the recursive distribution equation
$\Gamma_{\eta}\stackrel{{\scriptstyle d}}{{=}}\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d}\Gamma_{\eta,j}},$ (7.1)
where $V_{0}\sim N(0,1)$, the variables $\Gamma_{\eta,1},\dots,\Gamma_{\eta,K}$ and $\Gamma_{\eta}$ are i.i.d. and all variables on the right-hand side are independent.
We write
$\Gamma_{\eta}=X_{\eta}+\mathrm{i}Y_{\eta},$ (7.2)
with $X_{\eta}$ and $Y_{\eta}$ are real. We have $Y_{\eta}>0$ almost surely and
$X_{\eta}=\frac{U_{\eta}}{U_{\eta}^{2}+S_{\eta}^{2}},\qquad Y_{\eta}=\frac{S_{\eta}}{U_{\eta}^{2}+S_{\eta}^{2}},$ (7.3)
where
$U_{\eta}=V-E-t^{2}\sum_{j=1}^{d}X_{\eta,j},\qquad S_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta,j},$ (7.4)
with the $X_{\eta,j}$ and $Y_{\eta,j}$ defined as in (7.2). Moreover, the Schur complement formula (see, e.g., [8]) implies
$R_{00}(E+\mathrm{i}\eta)=\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d+1}\Gamma_{\eta}^{(j)}},$ (7.5)
where $V_{0}\sim N(0,1)$ is independent of the i.i.d. random variables $\Gamma_{\eta}^{(1)},\dots,\Gamma_{\eta}^{(d+1)}$ of $\Gamma_{\eta}$. In particular, defining
$B_{\eta}=\eta+t^{2}\sum_{j=1}^{d+1}Y_{\eta}^{(j)},$ (7.6)
where $Y_{\eta}^{(j)}$ is defined analogously to $Y_{\eta,j}$, there is a real random variable $A_{\eta}$, independent of $V_{0}$, such that
$\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=\frac{B_{\eta}}{(V_{0}-A_{\eta})^{2}+B_{\eta}^{2}}.$ (7.7)
For all lemmas in this section, we retain the notation and hypothesis of Lemma 3.6. In particular, $E\in\mathbb{R}$ is always fixed.
Lemma 7.1.
There exist constants $\eta_{0}(E)\in(0,1]$, $y_{*}(E)>0$, and $p_{*}(E)>0$, such that for all $\eta\in(0,\eta_{0}]$,
$\mathbb{P}(Y_{\eta}>y_{*})\geq p_{*}.$ (7.8)
Proof.
Choose $p_{0}>0$ and $\eta_{1}\in(0,1]$ so that, for all $0<\eta\leq\eta_{1}$,
$\mathbb{P}\bigl(\operatorname{Im}R(E+\mathrm{i}\eta)>c\bigr)\geq 2p_{0}.$ (7.9)
Fix $0<\eta\leq\eta_{1}$. Conditional on $(A_{\eta},B_{\eta})$ in (7.7), the event $\{\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\}$ is empty unless $B_{\eta}<1/c$, in which case it is equivalent to
$|V_{0}-A_{\eta}|<\sqrt{\frac{B_{\eta}}{c}-B_{\eta}^{2}}.$
Since $\|\rho\|_{\infty}=(2\pi)^{-1/2}$,
$\mathbb{P}\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\mid A_{\eta},B_{\eta}\bigr)\leq 2\|\rho\|_{\infty}c^{-1/2}B_{\eta}^{1/2}\mathbbm{1}_{\{B_{\eta}<1/c\}}.$ (7.10)
Set
$C_{\mathrm{G}}=2\|\rho\|_{\infty}c^{-1/2}.$
Taking expectations and using (7.9),
$2p_{0}\leq C_{\mathrm{G}}\mathbb{E}\bigl[B_{\eta}^{1/2}\mathbbm{1}_{\{B_{\eta}<1/c\}}\bigr].$ (7.11)
Choose
$b_{*}=\min\left\{\frac{1}{4c},\left(\frac{p_{0}}{C_{\mathrm{G}}}\right)^{2}\right\}.$
Then
$\mathbb{E}\bigl[B_{\eta}^{1/2}\mathbbm{1}_{\{B_{\eta}<1/c\}}\bigr]\leq b_{*}^{1/2}+c^{-1/2}\mathbb{P}(B_{\eta}>b_{*}),$
and therefore (7.11) implies
$2p_{0}\leq p_{0}+C_{\mathrm{G}}c^{-1/2}\mathbb{P}(B_{\eta}>b_{*}).$
Hence, for all $0<\eta\leq\eta_{1}$,
$\mathbb{P}(B_{\eta}>b_{*})\geq p_{1}>0,\qquad p_{1}=\frac{p_{0}c^{1/2}}{C_{\mathrm{G}}}.$ (7.12)
Now set
$\eta_{0}=\min\{\eta_{1},b_{*}/2\},\qquad y_{*}=\frac{b_{*}}{2(d+1)t^{2}}.$
If $0<\eta\leq\eta_{0}$ and $B_{\eta}>b_{*}$, then
$t^{2}\sum_{j=1}^{d+1}Y_{\eta}^{(j)}\geq B_{\eta}-\eta>b_{*}/2,$
so for at least one index $j$ we have $Y_{\eta}^{(j)}>y_{*}$. By the union bound and symmetry,
$p_{1}\leq\mathbb{P}\left(\bigcup_{j=1}^{d+1}\{Y_{\eta}^{(j)}>y_{*}\}\right)\leq(d+1)\mathbb{P}(Y_{\eta}>y_{*}).$
Thus (7.8) holds with
$p_{*}=\frac{p_{1}}{d+1}>0.\qed$
Lemma 7.2.
There exists a constant $C(t,E,d)>0$, such that for every $0<\eta\leq\eta_{0}$, the following bounds hold.
  1. (1)
    For all $u>0$,
    $\mathbb{P}(|X_{\eta}|>u)\leq\frac{C}{u},\qquad\mathbb{P}(Y_{\eta}>u)\leq\frac{C}{u}.$ (7.13)
  2. (2)
    For all $u\geq 1$,
    $\mathbb{P}(|U_{\eta}|>u)\leq\frac{C}{u}.$ (7.14)
  3. (3)
    For all $u\geq 2\eta_{0}$,
    $\mathbb{P}(S_{\eta}>u)\leq\frac{C}{u}.$ (7.15)
Proof.
Fix $a\in\mathbb{R}$, $s>0$, and $u>0$. The set
$\left\{x\in\mathbb{R}:\frac{|x|}{x^{2}+s^{2}}>u\right\}$
has Lebesgue measure at most $2/u$. Therefore, conditioning on
$A=E+t^{2}\sum_{j=1}^{d}X_{\eta,j}\qquad\text{and}\qquad S_{\eta}=s,$
we obtain from (7.3) that
$\mathbb{P}\bigl(|X_{\eta}|>u\mid A,S_{\eta}\bigr)\leq\frac{2\|\rho\|_{\infty}}{u}.$
Taking expectations yields the first bound of (7.13).
Likewise, the set
$\left\{x\in\mathbb{R}:\frac{s}{x^{2}+s^{2}}>u\right\}$
has Lebesgue measure at most $1/u$, hence
$\mathbb{P}\bigl(Y_{\eta}>u\mid A,S_{\eta}\bigr)\leq\frac{\|\rho\|_{\infty}}{u},$
and therefore the second bound of (7.13) follows after taking expectation.
For (7.14),we use (7.4) and the union bound to get
$\mathbb{P}\big(|U_{\eta}|>u\big)\leq\mathbb{P}\big(|V-E|>u/2\big)+\sum_{j=1}^{d}\mathbb{P}\left(|X_{\eta,j}|>\frac{u}{2dt^{2}}\right).$
The first term is bounded by $C/u$ for $u\geq 1$ and the second by (7.13), which gives (7.14).
Finally, if $u\geq 2\eta_{0}$ and $0<\eta\leq\eta_{0}$, then
$S_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta,j}>u$
implies
$\sum_{j=1}^{d}Y_{\eta,j}>\frac{u}{2t^{2}},$
so at least one index satisfies $Y_{\eta,j}>u/(2dt^{2})$. Hence
$\mathbb{P}(S_{\eta}>u)\leq\sum_{j=1}^{d}\mathbb{P}\left(Y_{\eta,j}>\frac{u}{2dt^{2}}\right),$
and (7.15) follows from (7.13). ∎
7.2. Bootstrap and Conclusion
Lemma 7.3.
For $\varepsilon>0$, define
$F_{\eta}(\varepsilon)=\mathbb{P}(Y_{\eta}\leq\varepsilon).$
There exist constants $C_{1}(t,E,d),C_{2}(t,E,d)>0$ and $\varepsilon_{1}(t,E,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$ and $0<\eta\leq\eta_{0}$,
$F_{\eta}(\varepsilon)\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d}+C_{2}\varepsilon^{1/3}.$ (7.16)
Proof.
We use that the event $\{Y_{\eta}\leq\varepsilon\}$ is contained in
$\{S_{\eta}\leq\varepsilon^{1/3}\}\cup\{\varepsilon^{1/3}<S_{\eta}\leq\varepsilon^{-1/3},Y_{\eta}\leq\varepsilon\}\cup\{S_{\eta}>\varepsilon^{-1/3}\}.$
On the first event, each summand in (7.4) satisfies
$Y_{\eta,j}\leq t^{-2}\varepsilon^{1/3},$
so by independence,
$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(t^{-2}\varepsilon^{1/3})^{d}.$ (7.17)
On the third event, (7.15) yields
$\mathbb{P}(S_{\eta}>\varepsilon^{-1/3})\leq C\varepsilon^{1/3}$ (7.18)
for all $\varepsilon\leq\varepsilon_{1}$ small enough that $\varepsilon^{-1/3}\geq 2\eta_{0}$.
Finally, on the middle event,
$\varepsilon\geq Y_{\eta}=\frac{S_{\eta}}{U_{\eta}^{2}+S_{\eta}^{2}},$
which implies
$U_{\eta}^{2}\geq\frac{S_{\eta}}{\varepsilon}-S_{\eta}^{2}.$
For fixed $\varepsilon$, the function $s\mapsto s/\varepsilon-s^{2}$ is increasing on $\varepsilon^{1/3},\varepsilon^{-1/3}]$ once $\varepsilon$ is small enough, because its derivative is $\varepsilon^{-1}-2s>0$ there. Hence, for $0<\varepsilon\leq\varepsilon_{1}$ sufficiently small,
$\frac{S_{\eta}}{\varepsilon}-S_{\eta}^{2}\geq\frac{\varepsilon^{1/3}}{\varepsilon}-\varepsilon^{2/3}=\varepsilon^{-2/3}-\varepsilon^{2/3}\geq\frac{1}{2}\varepsilon^{-2/3}.$
Therefore,
$\mathbb{P}\bigl(\varepsilon^{1/3}<S_{\eta}\leq\varepsilon^{-1/3},Y_{\eta}\leq\varepsilon\bigr)\leq\mathbb{P}\bigl(|U_{\eta}|\geq 2^{-1/2}\varepsilon^{-1/3}\bigr)\leq C\varepsilon^{1/3}$
by (7.14), for all $\varepsilon\leq\varepsilon_{1}$ small enough that $2^{-1/2}\varepsilon^{-1/3}\geq 1$. Combining this with (7.17) and (7.18) proves (7.16) with $C_{1}=t^{-2}$. ∎
Lemma 7.4.
There exist constants $C_{*}(t,E,d)>0$, and $\varepsilon_{*}(E,t,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{*}$ and $0<\eta\leq\eta_{0}$,
$\mathbb{P}(Y_{\eta}\leq\varepsilon)\leq C_{*}\varepsilon^{\beta},\qquad\beta=\frac{1}{20}.$ (7.19)
Proof.
Let $C_{1},C_{2},\varepsilon_{1}$ be as in Lemma 7.3. Choose
$\tilde{\varepsilon}_{0}\in(0,\min\{y_{*},\varepsilon_{1},1\}]$
so small that the sequence $(\tilde{\varepsilon}_{n})_{n\geq 0}$, defined by
$\tilde{\varepsilon}_{n+1}=\left(\frac{\tilde{\varepsilon}_{n}}{2C_{1}}\right)^{3},$ (7.20)
satisfies, for every $n\geq 0$,
$L_{n}=-\ln\tilde{\varepsilon}_{n}\geq\max\{12\ln C_{2},20\ln 2\},\qquad L_{n+1}\leq 4L_{n}.$ (7.21)
This is possible because $L_{n+1}=3L_{n}+3\ln(2C_{1})$.
Define
$q_{n}=\sup_{0<\eta\leq\eta_{0}}F_{\eta}(\tilde{\varepsilon}_{n}),\qquad\nu_{n}=-\ln q_{n}.$
Since $\tilde{\varepsilon}_{0}\leq y_{*}$ and Lemma 7.1 gives $\mathbb{P}(Y_{\eta}>y_{*})\geq p_{*}$, we have
$q_{0}\leq 1-p_{*}<1,\qquad\nu_{0}>0.$ (7.22)
Moreover, because $C_{1}\tilde{\varepsilon}_{n+1}^{1/3}=\tilde{\varepsilon}_{n}/2<\tilde{\varepsilon}_{n}<\varepsilon_{1}$, Lemma 7.3 implies
$q_{n+1}\leq q_{n}^{d}+C_{2}\tilde{\varepsilon}_{n+1}^{1/3}.$ (7.23)
By (7.21),
$C_{2}\tilde{\varepsilon}_{n+1}^{1/3}=\exp\left(-\frac{1}{3}L_{n+1}+\ln C_{2}\right)\leq\exp\left(-\frac{1}{4}L_{n+1}\right).$
Therefore,
$\nu_{n+1}\geq\min\left\{d\nu_{n},\frac{1}{4}L_{n+1}\right\}-\ln 2.$ (7.24)
We claim that there exists an index $n_{*}$ such that
$d\nu_{n_{*}}\geq\frac{1}{4}L_{n_{*}+1}.$ (7.25)
If not, then (7.24) would imply, for all $n$, that
$\nu_{n+1}\geq d\nu_{n}-\ln 2.$
Setting
$w_{n}=\nu_{n}+\frac{\ln 2}{d-1},$
we would have $w_{n+1}\geq dw_{n}$, hence
$w_{n}\geq d^{n}w_{0}.$
On the other hand, the negation of (7.25) gives
$\nu_{n}<\frac{L_{n+1}}{4d}\leq\frac{4^{n}L_{1}}{4d}$
by (7.21), so $w_{n}\leq C4^{n}$. Since $d\geq 20>4$, this is impossible for large $n$. Thus there exists some $n_{*}\in\mathbb{N}$ such that (7.25) holds.
Now (7.24), (7.25), and (7.21) give
$\nu_{n_{*}+1}\geq\frac{1}{4}L_{n_{*}+1}-\ln 2\geq\frac{1}{5}L_{n_{*}+1}.$
Because $L_{n_{*}+2}\leq 4L_{n_{*}+1}$, we obtain
$\nu_{n_{*}+1}\geq\frac{1}{20}L_{n_{*}+2}.$
We now prove by induction that, for all $n\geq n_{*}+1$,
$\nu_{n}\geq\beta L_{n+1},\qquad\beta=\frac{1}{20}.$ (7.26)
The base case has just been established. If (7.26) holds at index $n$, then
$d\nu_{n}\geq d\beta L_{n+1}=\frac{d}{20}L_{n+1}\geq L_{n+1}\geq\frac{1}{4}L_{n+2},$
where the last inequality again uses (7.21). Hence (7.24) and $L_{n+2}\geq 20\ln 2$ yield
$\nu_{n+1}\geq\frac{1}{4}L_{n+2}-\ln 2\geq\beta L_{n+2}.$
This proves (7.26).
Consequently, for all $n\geq n_{*}+1$,
$q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}.$
Now let $0<\varepsilon\leq\varepsilon_{*}=\tilde{\varepsilon}_{n_{*}+2}$ and choose $n\geq n_{*}+1$ so that
$\tilde{\varepsilon}_{n+1}<\varepsilon\leq\tilde{\varepsilon}_{n}.$
Since $F_{\eta}$ is nondecreasing,
$F_{\eta}(\varepsilon)\leq q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}<\varepsilon^{\beta}.$
Thus (7.19) holds with $C_{*}=1$ and $\beta=1/20$. ∎
Lemma 7.5.
There exists a constant $C(E,t,d)>0$ such that
$\sup_{0<\eta\leq\eta_{0}}\mathbb{E}[B_{\eta}^{-1}]\leq C.$ (7.27)
Proof.
Let $\beta=1/20$ and $\varepsilon_{*}>0$ be given by Lemma 7.4. If $0<\varepsilon\leq t^{2}\varepsilon_{*}$ and $B_{\eta}\leq\varepsilon$, then each summand in (7.6) with $1\leq j\leq d+1$ satisfies
$Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}},$
since all summands are nonnegative. Therefore, by independence and Lemma 7.4,
$\mathbb{P}(B_{\eta}\leq\varepsilon)\leq\prod_{j=1}^{d+1}\mathbb{P}\left(Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}}\right)\leq C\varepsilon^{\beta(d+1)}.$
Because $\beta(d+1)=(d+1)/20>1$, we conclude that
$\sup_{0<\eta\leq\eta_{0}}\int_{1}^{\infty}\mathbb{P}(B_{\eta}<s^{-1})\,ds<\infty,$
which is equivalent to (7.27). ∎
We now return to Lemma 3.6.
Proof.
Let $\eta\in(0,\eta_{0}]$. Conditioning on $(A_{\eta},B_{\eta})$ in (7.7),
$\mathbb{E}\left[\big(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\bigr)^{2}\mid A_{\eta},B_{\eta}\right]=\int_{\mathbb{R}}\rho(v)\frac{B_{\eta}^{2}}{((v-A_{\eta})^{2}+B_{\eta}^{2})^{2}}\,dv.$
Since $\rho$ is bounded,
$\mathbb{E}\left[\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\bigr)^{2}\mid A_{\eta},B_{\eta}\right]\leq\|\rho\|_{\infty}\int_{\mathbb{R}}\frac{B_{\eta}^{2}}{(x^{2}+B_{\eta}^{2})^{2}}\,dx=\frac{\pi\|\rho\|_{\infty}}{2B_{\eta}}.$
Taking expectations and applying Lemma 7.5,
$\sup_{0<\eta\leq\eta_{0}}\mathbb{E}\left[\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\bigr)^{2}\right]<\infty.$ (7.28)
For $\eta\in[\eta_{0},1]$, the trivial resolvent bound $|R_{00}(E+\mathrm{i}\eta)|\leq\eta_{0}^{-1}$ gives
$\sup_{\eta\in[\eta_{0},1]}\mathbb{E}\left[\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\bigr)^{2}\right]\leq\eta_{0}^{-2}<\infty.$
Together with (7.28), this proves (3.6). ∎
Appendix A Combes–Thomas Bound
In the proof of Combes–Thomas estimates, one typically conjugates the Hamiltonian $H$ by an exponential weight operator
$(W\psi)(x)=e^{\mu w_{X}(x)}\psi(x),\qquad w_{X}(x)=\operatorname{dist}(x,X),$
then writes $H_{\mu}=W^{-1}HW$ and manipulates resolvents. On an infinite graph, $W$ is usually an unbounded operator on $\ell^{2}(\mathcal{V})$, so one may worry about domain issues. In this appendix, we address these concerns in a standard way. We first prove the desired Combes–Thomas estimate on finite volumes, then pass to the infinite graph by strong resolvent convergence and a finite-rank sandwiching argument.
Throughout this appendix, we let $\mathcal{G}=(V,E)$ be a graph (finite or finite) with maximum degree at most $d$, and we retain the notation introduced in Section 6. Additionally, for vertices $v,w$ in a graph, we write $v\sim w$ if there is an edge between $v$ and $w$.
A.1. Finite-Volume Estimate
Fix a finite set $X\subset\mathcal{V}$. For all $r\in\mathbb{N}$ define the (finite) neighborhood
$\Lambda_{r}=\{x\in\mathcal{V}:\operatorname{dist}(x,X)\leq r\}.$
If $\mathcal{G}$ has degree bounded by $d$ and $X$ is finite, then $\Lambda_{r}$ is finite for every $r$. Define the restriction $H_{\Lambda_{r}}$ as the operator on $\ell^{2}(\Lambda_{r})$ obtained by removing all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$, and write
$H_{\Lambda_{r}}=-tA_{\Lambda_{r}}+V_{\Lambda_{r}}.$
Lemma A.1.
Fix $z=E+\mathrm{i}\eta\in\mathbb{H}$ and $t>0$. Recall $\mu_{\mathrm{CT}}$ from Lemma 6.2. Then for any finite sets $X,Y\subset\Lambda_{r}$,
$\big\|\chi_{X}\,(H_{\Lambda_{r}}-z)^{-1}\,\chi_{Y}\big\|\leq\frac{2}{\eta}\,e^{-\mu_{\mathrm{CT}}\operatorname{dist}(X,Y)}.$
Proof.
We abbreviate $\mu=\mu_{\mathrm{CT}}$. Since $\Lambda_{r}$ is finite, $H_{\Lambda_{r}}$ is a finite Hermitian matrix and all algebraic manipulations below are permissible. Fix a nonempty set $X\subset\Lambda_{r}$ and set $w_{X}(x)=\operatorname{dist}(x,X)$ for $x\in\Lambda_{r}$. Define the bounded weight operator $W_{r}$ on $\ell^{2}(\Lambda_{r})$ by
$(W_{r}\psi)(x)=e^{\mu w_{X}(x)}\psi(x),\qquad x\in\Lambda_{r},$
and set
$H_{\Lambda_{r},\mu}=W_{r}^{-1}H_{\Lambda_{r}}W_{r}=-tW_{r}^{-1}A_{\Lambda_{r}}W_{r}+V_{\Lambda_{r}}.$
In the last equality, we used that since $V_{\Lambda_{r}}$ is diagonal, it commutes with $W_{r}$.
For neighbors $x\sim y$ in $\Lambda_{r}$ we have $|w_{X}(x)-w_{X}(y)|\leq 1$, so
$\big|(W_{r}^{-1}A_{\Lambda_{r}}W_{r})_{xy}\big|=e^{\mu(w_{X}(y)-w_{X}(x))}\leq e^{\mu}.$
Hence $J_{r}=W_{r}^{-1}A_{\Lambda_{r}}W_{r}-A_{\Lambda_{r}}$ is supported on edges and satisfies $|(J_{r})_{xy}|\leq e^{\mu}-1$ for $x\sim y$. Analogously to Lemma 6.1,
$\|J_{r}\|\leq d(e^{\mu}-1).$
Therefore
$\|H_{\Lambda_{r},\mu}-H_{\Lambda_{r}}\|\leq td(e^{\mu}-1).$
By the definition of $\mu$, we have $td(e^{\mu}-1)=\eta/4$. Also, $\|(H_{\Lambda_{r}}-z)^{-1}\|\leq 1/\eta$. Thus
$\|(H_{\Lambda_{r},\mu}-H_{\Lambda_{r}})(H_{\Lambda_{r}}-z)^{-1}\|\leq\frac{1}{4},$
so $H_{\Lambda_{r},\mu}-z$ is invertible, and
$\|(H_{\Lambda_{r},\mu}-z)^{-1}\|\leq\frac{1}{1-1/4}\,\|(H_{\Lambda_{r}}-z)^{-1}\|\leq\frac{2}{\eta}.$
Now $\chi_{X}W_{r}=\chi_{X}$ and, since $w_{X}(y)\geq\operatorname{dist}(X,Y)$ for all $y\in Y$,
$\|W_{r}^{-1}\chi_{Y}\|\leq e^{-\mu\operatorname{dist}(X,Y)}.$
Hence
$\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y}\|=\|\chi_{X}W_{r}(H_{\Lambda_{r},\mu}-z)^{-1}W_{r}^{-1}\chi_{Y}\|\leq\frac{2}{\eta}e^{-\mu\operatorname{dist}(X,Y)}.$
This proves the claim. ∎
A.2. Passage to Infinite Volume
We define the decoupled operator
$\widehat{H}_{r}=H_{\Lambda_{r}}\oplus H_{\Lambda_{r}^{c}},$
on $\ell^{2}(\mathcal{V})$ by by cutting all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ in the underlying graph and removing the corresponding entries from $H$ (so that $\widehat{H}_{r}$ is self-adjoint). If $X,Y\subset\Lambda_{r}$, then
$\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}=\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y}.$ (A.1)
Lemma A.2.
For every $z\in\mathbb{H}$,
$\lim_{r\rightarrow\infty}(\widehat{H}_{r}-z)^{-1}\varphi=(H-z)^{-1}\varphi$
in norm for all $\varphi\in\ell^{2}(\mathcal{V})$.
Proof.
Let $C_{c}(\mathcal{V})$ be the finitely supported functions on $\mathcal{V}$. Since $A$ is bounded and $V$ is a self-adjoint multiplication operator, $C_{c}(\mathcal{V})$ is a core for $H$ and every $\widehat{H}_{r}$. For any $\psi\in C_{c}(V)$, we have $\operatorname{supp}\psi\subset\Lambda_{r}$ for all sufficiently large $r$, in which case the cut edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ do not affect the value of $H\psi$ because $\psi$ vanishes on $\Lambda_{r}^{c}$. Hence for all sufficiently large $r$,
$\widehat{H}_{r}\psi=H\psi,$
and therefore $\widehat{H}_{r}\psi\to H\psi$ as $r\rightarrow\infty$ for every $\psi\in C_{c}(V)$. By the core criterion for strong resolvent convergence (e.g., [31, Theorem VIII.25]), this implies $\widehat{H}_{r}\to H$ in the strong resolvent sense, which is exactly the desired conclusion. ∎
Lemma A.3.
Let $(T_{r})_{r=1}^{\infty}$ and $T$ be bounded operators on a Hilbert space such that $\sup_{r}\|T_{r}\|<\infty$, and suppose $\lim_{r\rightarrow\infty}T_{r}\to T$ in the strong operator topology. If $P$ and $Q$ are finite-rank orthogonal projections, then
$\lim_{r\rightarrow\infty}\|PT_{r}Q-PTQ\|=0.$
Proof.
Let $F=\mathrm{Ran}(Q)$; then $F$ is finite-dimensional. Define $S_{r}=P(T_{r}-T)Q$, which maps $F$ into the finite-dimensional space $\mathrm{Ran}(P)$. Strong convergence implies that for each $v\in F$, $\|S_{r}v\|\to 0$. Since $F\cap\{v:\|v\|=1\}$ is compact and $\sup_{r}\|S_{r}\|<\infty$, pointwise convergence on the unit sphere implies uniform convergence,
$\sup_{v\in F,\ \|v\|=1}\|S_{r}v\|\to 0,$
which is exactly $\|S_{r}\|\to 0$. ∎
We can now prove Lemma 6.2.
Proof of Lemma 6.2.
We first prove the claim when both $X$ and $Y$ are finite. Fix finite $X,Y\subset V$ and choose $r$ large enough so that $X\cup Y\subset\Lambda_{r}$. By Lemma A.1 and the identity (A.1),
$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}\big\|=\big\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y}\big\|\leq\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$ (A.2)
By Lemma A.2, $(\widehat{H}_{r}-z)^{-1}\to(H-z)^{-1}$ in the strong operator topology. Further, we have the elementary bound $\sup_{r}\|(\widehat{H}_{r}-z)^{-1}\|\leq 1/\eta$. This allows us to apply Lemma A.3 with $T_{r}=(\widehat{H}_{r}-z)^{-1}$, $T=(H-z)^{-1}$, $P=\chi_{X}$, and $Q=\chi_{Y}$ to obtain
$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}-\chi_{X}(H-z)^{-1}\chi_{Y}\big\|\to 0.$
Together with (A.2), this proves the claim when $X$ and $Y$ are finite.
For the general case, pick increasing sequences of finite sets $(X_{n})_{n=1}^{\infty}$ and $(Y_{n})_{n=1}^{\infty}$ such that $\lim_{n\rightarrow\infty}X_{n}=X$ and $\lim_{n\rightarrow\infty}Y_{n}=Y$. Since $X_{n}\subset X$ and $Y_{n}\subset Y$, we have
$\operatorname{dist}(X_{n},Y_{n})\geq\operatorname{dist}(X,Y).$
Therefore, by the finite-set Combes–Thomas estimate,
$\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\|\leq C\,e^{-\mu\,\operatorname{dist}(X_{n},Y_{n})}\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\qquad\text{for all }n.$ (A.3)
Next, for any $\psi\in\ell^{2}(\mathcal{V})$,
$\bigl(\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}-\chi_{X}(H-z)^{-1}\chi_{Y}\bigr)\psi=$ $(\chi_{X_{n}}-\chi_{X})(H-z)^{-1}\chi_{Y_{n}}\psi$
$+\chi_{X}(H-z)^{-1}(\chi_{Y_{n}}-\chi_{Y})\psi.$
Since $(H-z)^{-1}$ is bounded and $\chi_{X_{n}}\to\chi_{X}$, $\chi_{Y_{n}}\to\chi_{Y}$ strongly, both terms on the right-hand side converge to zero in norm. In combination with , this shows that for any $\psi\in\ell^{2}(\mathcal{V})$,
$\|\chi_{X}R(z)\chi_{Y}\psi\|=\lim_{n\to\infty}\|\chi_{X_{n}}R(z)\chi_{Y_{n}}\psi\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\|\psi\|.$
Taking the supremum over all $\psi$ with $\|\psi\|=1$ yields
$\|\chi_{X}R(z)\chi_{Y}\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}=\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$
This proves the claim. ∎
Appendix B Lower Bound on the Density of States
This appendix is devoted to the proof of Lemma 5.6. We begin with an elementary fact about self-adjoint operators.
Lemma B.1.
Let $H$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\varphi\in\mathcal{H}$ such that $\varphi\neq 0$, $E\in\mathbb{R}$, and $\varepsilon>0$. For all Borel sets $J\subset\mathbb{R}$, define.
$\mu_{\varphi}(J)=\langle\varphi,\mathbbm{1}_{J}(H)\varphi\rangle.$
Then
$\mu_{\varphi}([E-\varepsilon,E+\varepsilon])\geq\|\varphi\|^{2}-\frac{\|(H-E)\varphi\|^{2}}{\varepsilon^{2}}.$
Proof.
The spectral theorem implies
$\|(H-E)\varphi\|^{2}=\int_{\mathbb{R}}(\lambda-E)^{2}\,\mu_{\varphi}(d\lambda)\geq\varepsilon^{2}\,\mu_{\varphi}\big(\mathbb{R}\setminus[E-\varepsilon,E+\varepsilon]\big),$
which yields the conclusion. ∎
For all $M>0$, define the $M$-strength barrier event at distance $1$ from the root of $\mathbb{T}$ by
$\mathcal{B}_{M}=\Big\{\min_{y\sim 0}|V_{y}-V_{0}|\geq M\Big\}.$
On $\mathcal{B}_{M}$, define
$\psi=\delta_{0}+\sum_{y\sim 0}a_{y}\delta_{y},\qquad a_{y}=\frac{t}{V_{y}-V_{0}}.$
Lemma B.2.
On $\mathcal{B}_{M}$,
$\|\psi\|^{2}\leq 1+\frac{dt^{2}}{M^{2}},\qquad\|(H-V_{0})\psi\|^{2}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
Proof.
By definition, $|a_{y}|\leq t/M$, so
$\|\psi\|^{2}=1+\sum_{y\sim 0}|a_{y}|^{2}\leq 1+\frac{dt^{2}}{M^{2}}.$
Also,
$\big((H-V_{0})\psi\big)(0)=-t\sum_{y\sim 0}a_{y},$
hence $|((H-V_{0})\psi)(0)|\leq t^{2}d/M$.
At any $y\sim 0$, by definition of $a_{y}$ we have
$\big((H-V_{0})\psi\big)(y)=(V_{y}-V_{0})a_{y}-t=0.$
At vertices $u$ with $\operatorname{dist}(u,0)=2$, writing $y$ for the parent of $u$,
$\big((H-V_{0})\psi\big)(u)=-ta_{y},\qquad\big|\big((H-V_{0})\psi\big)(u)\big|\leq\frac{t^{2}}{M}.$
There are exactly d(d1)d(d-1) such vertices, therefore
$\|(H-V_{0})\psi\|^{2}\leq\frac{t^{4}d^{2}}{M^{2}}+\frac{t^{4}d(d-1)}{M^{2}}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
Lemma B.3.
Let $I\subset\mathbb{R}$ be Borel and write $P_{I}=\mathbbm{1}_{I}(H)$. If $\psi=\delta_{0}+\xi$, then
$\mu_{0}(I)\geq\big(\sqrt{\mu_{\psi}(I)}-\|\xi\|\big)^{2}.$
Proof.
Since $\delta_{0}=\psi-\xi$,
$\sqrt{\mu_{0}(I)}=\|P_{I}\delta_{0}\|=\|P_{I}(\psi-\xi)\|\geq\|P_{I}\psi\|-\|P_{I}\xi\|\geq\sqrt{\mu_{\psi}(I)}-\|\xi\|.$
Squaring gives the claim. ∎
Proposition B.4.
Assume $t>0$, fix $M>0$, and set $\varepsilon=4t^{2}d/M$. Then on $\mathcal{B}_{M}$,
$\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq\left(\sqrt{\frac{7}{8}}-\frac{\sqrt{d}\,t}{M}\right)^{2}.$
Proof.
By Lemma B.1 and Lemma B.2,
$\mu_{\psi}([V_{0}-\varepsilon,V_{0}+\varepsilon])\geq\|\psi\|^{2}-\frac{\|(H-V_{0})\psi\|^{2}}{\varepsilon^{2}}\geq 1-\frac{2t^{4}d^{2}/M^{2}}{16t^{4}d^{2}/M^{2}}=\frac{7}{8}.$
Also,
$\xi=\sum_{y\sim 0}a_{y}\delta_{y},\qquad\|\xi\|^{2}=\sum_{y\sim 0}|a_{y}|^{2}\leq\frac{dt^{2}}{M^{2}},$
so $\|\xi\|\leq\sqrt{d}\,t/M$. Apply Lemma B.3. ∎
We are now ready for the proof of Lemma 5.6.
Proof of Lemma 5.6.
Fix $E\in[-D,D]$. Set $M=4\sqrt{d}t$ and $\varepsilon=4t^{2}d/M$, and define
$\mathcal{A}_{E,\varepsilon,M}=\big\{|V_{0}-E|\leq\varepsilon/2\big\}\cap\mathcal{B}_{M}.$
On $\mathcal{A}_{E,\varepsilon,M}$ we have $|V_{0}-E|\leq\varepsilon/2$, hence
$V_{0}-\varepsilon,V_{0}+\varepsilon]\subset(E-2\varepsilon,E+2\varepsilon).$
By Proposition B.4, there exists $c_{0}>0$ such that on $\mathcal{B}_{M}$ (hence on $\mathcal{A}_{E,\varepsilon,M}$),
$\mu_{0}\big((E-2\varepsilon,E+2\varepsilon)\big)\geq\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq c_{0}.$
Taking expectation gives
$\nu\big((E-2\varepsilon,E+2\varepsilon)\big)\geq c_{0}\,\mathbb{P}(\mathcal{A}_{E,\varepsilon,M}).$
Observe that
$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})=\int_{E-\varepsilon/2}^{E+\varepsilon/2}\rho(v)\,\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)\,dv.$
Given $V_{0}=v$, the $d$ neighbor potentials are i.i.d. with law $\rho$, so
$\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)=\mathbb{P}(|V-v|\geq M)^{d}.$
Lower bounding both factors on $|v-E|\leq\varepsilon/2$ yields
$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})\geq\varepsilon\Big(\inf_{|v-E|\leq\varepsilon/2}\rho(v)\Big)\Big(\inf_{|v-E|\leq\varepsilon/2}\mathbb{P}(|V-v|\geq M)\Big)^{d}.$
Both factors on the right-hand side can be bounded below by a constant depending only on $d$, $D$, and $\mathfrak{L}$. This completes the proof. ∎

### 19 Mar 2026 v2

@@
 Abstract.
-We determine the phase diagram of the Anderson tight-binding model on regular random graphs with Gaussian disorder distribution and sufficiently large degree.
+We determine the phase diagram of the Anderson tight-binding model on random regular graphs with Gaussian disorder and sufficiently large degree.
 In particular, we prove that if the degree is fixed and the number of vertices goes to infinity, the spectrum asymptotically consists of a finite delocalized interval surrounded by two unbounded localized components.
 Our argument uses a recent description of the spectrum of the tight-binding model on the Bethe lattice (Aggarwal–Lopatto, 2025).
-By viewing the Bethe lattice as the local limit of a random regular graph, and proving several preliminary estimates, we transfer this characterization of the spectrum of the limiting model to its finite approximants.
+By viewing the Bethe lattice as the local limit of a random regular graph, and establishing suitable concentration, eigenvalue-counting, and resolvent estimates, we transfer this characterization of the spectrum of the limiting model to the finite-volume setting.
 1. Introduction
@@
 A standard setting for studying wave transport in disordered media is the tight-binding model, which served as the basis for Anderson’s original investigations [12].
-Let $\mathcal{G}$ be a graph (finite or infinite), let $A$ denote the adjacency operator of $\mathcal{G}$, and let $V$ denote a diagonal operator of independent, identically distributed (i.i.d) random variables.
-A particle whose time evolution is governed by $A$ moves freely by nearest-neighbor hopping, displaying diffusive behavior.
+Let $\mathcal{G}$ be a graph (finite or infinite), let $A$ denote the adjacency operator of $\mathcal{G}$, and let $V$ denote a diagonal operator of independent, identically distributed (i.i.d.) random variables.
+A particle whose time evolution is governed by $A$ moves freely by nearest-neighbor hopping.
 In contrast, the time evolution induced by $V$ traps a particle at its initial site.
 The tight-binding Hamiltonian is defined as $H_{t}=-tA+V$; it represents a contest between the opposite influences of $A$ and $V$ mediated by $t$, the inverse disorder strength.
-When $t$ is sufficiently small, $V$ dominates, and particle remains frozen in position space.
-However, when $t$ is sufficiently large, it is expected that the particle may diffuse, at least for certain energies and graph geometries, leading to a region (or regions) of delocalization in the spectrum separated from regions of localization by sharp transitions called mobility edges [9].
+When $t$ is sufficiently small, $V$ dominates, and the particle remains localized near its initial position.
+However, when $t$ is sufficiently large, one expects that extended states may emerge, at least for certain energies and graph geometries, leading to a region (or regions) of delocalization in the spectrum separated from regions of localization by sharp transitions called mobility edges [9].
 An important special case is when $\mathcal{G}$ is the infinite $d$-regular tree, also known as the Bethe lattice.
@@
 For instance, the self-consistent theory of Anderson localization was developed using equations which are approximate on a true lattice, but become exact on the Bethe lattice [2, 1].
-Much is now known rigorously about this model, including delocalization at small energies [27, 26, 7, 22], localization at large energies [10, 6], and a nearly complete description of the phase diagram at low disorder (omitting only small intervals around the expected locations of the mobility edges) [20].
+Much is now known rigorously about this model, including delocalization at small energies [26, 28, 27, 7, 22], localization at large energies [10, 6], and a nearly complete description of the phase diagram at low disorder (omitting only small intervals around the expected locations of the mobility edges) [21].
 Further, [8] established a criterion for absolutely continuous spectrum based on the fractional moments of resolvent entries.
 This was used in [14], along with the Simon–Wolff criterion for localization [33, 6], to provide quantitative bounds on the endpoints of the regions of localization and delocalization, mapping a large part of the spectrum.
-As the degree of the Bethe lattice tends to infinity, these bounds become increasingly better and characterize the entire the spectrum in the limit (but not for any finite degree).
+As the degree of the Bethe lattice tends to infinity, these bounds become increasingly better and characterize the entire spectrum in the limit (but not for any finite degree).
 Recently, building on the work of [14, 33, 6, 8, 4], mobility edges separating regions of delocalization and localization were shown to exist in [5] when the degree $d$ is sufficiently large and the potential has unbounded support.
@@
 (Simply truncating the Bethe lattice to a finite tree leads to a graph where most vertices are leaves, which produce undesirable boundary effects.) Since random regular graphs converge locally to the Bethe lattice, their spectral and eigenfunction observables measure how Bethe-lattice physics is modified by long loops and finite-size effects, providing a controlled high-dimensional reference for comparison with finite-dimensional Anderson localization [17, 18, 34].
-Further, there has been a resurgence of interest in the tight-binding model on the random regular graph due to its connection with many-body localization, a phenomenon that has recently attracted significant interested in the physics community [30].
+Further, there has been a resurgence of interest in the tight-binding model on the random regular graph due to its connection with many-body localization, a phenomenon that has recently attracted significant interest in the physics community [31].
 Specifically, in suitable many-body problems, certain hierarchical models in Fock space can be approximately mapped to tight-binding models on locally tree-like graphs, for which a random regular graph is a useful proxy [35].
@@
 In [23], it was shown that for compactly supported and bounded disorder densities, the integrated density of states converges to the integrated density of states of the Bethe lattice, as $d$ is fixed and the number of vertices $N$ tends to infinity, with a rate of convergence of order $1/\ln N$.
-This work also showed for that if the eigenvectors in a finite energy window in the absolutely continuous spectrum for the Bethe lattice are averaged, weighted by their squared amplitudes at a uniformly random vertex $x_{0}$, the resulting weighted mass does not concentrate in small balls around $x_{0}$; this is a form of eigenvector delocalization.
+This work also showed that if the eigenvectors in a finite energy window in the absolutely continuous spectrum for the Bethe lattice are averaged, weighted by their squared amplitudes at a uniformly random vertex $x_{0}$, the resulting weighted mass does not concentrate in small balls around $x_{0}$; this is a form of eigenvector delocalization.
 Further, [24] showed a criterion for the presence of Poisson spectral statistics based on the fractional moment decay of off-diagonal resolvent entries and applied it to the edges of the spectrum (for compactly supported disorder distributions).
@@
 Fix an integer $d\geq 3$, and let $\mathbb{N}$ denote the positive integers.
-For each $N\in\mathbb{N}$, let $G_{N}$ be a (uniformly) random $d$-regular graph on the vertex set $N]=\{1,\dots,N\}$, and let $A_{N}$ denote its adjacency matrix.
+When discussing $d$-regular random graphs with $N$ vertices, we always restrict to the subset of integers $N$ such that $N\geq d+1$ and $Nd$ is even (so that the configuration model construction of $d$-regular graphs is valid).
+Statements quantifying over $N$ implicitly quantify over only this subset, and all limits $N\rightarrow\infty$ are implicitly along the subsequence of these values.
+Let $\mathcal{G}_{N}$ be a (uniformly) random $d$-regular graph on the vertex set $N]=\{1,\dots,N\}$, and let $A_{N}$ denote its adjacency matrix.
 Let $(g_{i})_{i=1}^{\infty}$ be i.i.d.
-Gaussian random variables with mean zero and variance one.
+Gaussian random variables with mean zero and variance one (independent of $\mathcal{G}_{N}$).
 Define the diagonal matrix
@@
 where $\delta_{k\ell}$ is the Kronecker delta function.
-The eigenvectors $u_{k}$ are determined only up to a sign, which we may select arbitrarily.
-We let $\Lambda$ denote the multiset of eigenvalues of $W_{N}$ (which may have duplicated elements due to repeated eigenvalues), and given an interval $I\subset\mathbb{R}$, we let $\Lambda_{I}$ be the multiset of eigenvalues contained in $\mathbb{R}$.
+The eigenvectors $u_{k}$ are generically determined only up to a sign, which we may select arbitrarily.
+We let $\Lambda$ denote the multiset of eigenvalues of $W_{N}$ (which may have duplicated elements due to repeated eigenvalues), and given an interval $I\subset\mathbb{R}$, we let $\Lambda_{I}$ be the multiset of eigenvalues contained in $I$.
 While these objects depend on $d$, $t$, and $N$, we suppress this dependence in the notation.
@@
 $P_{I}(j)=\frac{N}{|\Lambda_{I}|}\sum_{\lambda_{i}\in\Lambda_{I}}u_{i}(j)^{2},\qquad Q_{I}=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{2}.$
+When $|\Lambda_{I}|=0$, we set $P_{I}(j)=0$.
 To understand the meaning of these observables, we consider two extreme cases.
-First, suppose that all eigenvectors with eigenvalues in $I$ are as delocalized as possible, so that $u_{i}(j)=N^{-1/2}$ for all $j\in[N]$ and $i$ such that $\lambda_{i}\in I$.
+First, suppose that all eigenvectors with eigenvalues in $I$ are as delocalized as possible, so that $u_{i}(j)^{2}=N^{-1}$ for all $j\in[N]$ and $i$ such that $\lambda_{i}\in I$.
 Then $P_{I}(j)=1$ for all $j$, implying $Q_{I}=1$.
@@
 Fix an energy $E\in\mathbb{R}$. We say that delocalization holds at $E$ for $W_{N}$ if there exist constants $C(E)>1$ and $c(E)>0$ such that for all $\varepsilon\in(0,c)$,
-$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\leq C\Big)=1.$
+$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\leq C\Big)=1,$
 where $I(\varepsilon)=[E-\varepsilon,\,E+\varepsilon]$. We say that localization holds at $E$ if for every $D>0$, there exists $c(D)>0$ such that for all $\varepsilon\in(0,c)$,
@@
 We now state our main theorem.
-Let $\rho$ denote the density function of a mean zero, variance one Gaussian random variable.
-The scaling of the strength disorder as $(d\ln d)^{-1}$ is conventional; it ensures the mobility edge locations tend to a constant-order limit as $d$ grows large.
+The scaling of the inverse disorder strength as $(d\ln d)^{-1}$ is conventional; it ensures the mobility edge locations tend to constant-order limits as $d$ grows large.
+The constants arise from the fact that the supremum of the density of a mean zero, variance one Gaussian random variable is $(2\pi)^{-1/2}$.
 Theorem 1.3.
 For every real number $\mathfrak{L}>0$, there exists a constant $d_{0}(\mathfrak{L})>0$ such that the following holds for all $d\geq d_{0}$. Fix $g\in\mathbb{R}$ such that
-$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$
-and let $t=g(d\ln d)^{-1}$.
-Then there exists $\mathfrak{E}>0$ such that the set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$.
-Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$, localization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|>\mathfrak{M}\}$, and delocalization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|<\mathfrak{M}\}$.
+$\sqrt{\frac{\pi}{8}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$
+and let
+$t=g(d\ln d)^{-1},\qquad\mathfrak{E}=\sqrt{2\ln\!\left(\frac{4g}{\sqrt{2\pi}}\right)}.$
+Then there exists $\mathfrak{M}(d,g)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathfrak{L}^{-1}$, localization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|>\mathfrak{M}\}$, and delocalization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|<\mathfrak{M}\}$.
 1.3. Proof Ideas
-As mentioned previously, the spectral decomposition of the tight-binding model of the Bethe lattice was characterized in [5] (for the disorder regime considered in our main result).
-Since the Bethe lattice is the local limit of random regular graphs, it natural to attempt to use the Bethe lattice result to prove Theorem 1.3, and this is indeed our strategy.
+As mentioned previously, the spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] (for the disorder regime considered in our main result).
+Since the Bethe lattice is the local limit of random regular graphs, it is natural to attempt to use the Bethe lattice result to prove Theorem 1.3, and this is indeed our strategy.
 However, this route requires suitable estimates for the tight-binding model on the regular random graph, which we enumerate in Section 3.
-Specifically, we require concentration for fractional moments of the off-diagonal resolvent entries (Lemma 3.3), upper and lower bounds on the eigenvalue density (Lemma 3.4), convergence of the diagonal resolvent entries of the random regular graph to the diagonal resolvent entry at the root of the Bethe lattice (Lemma 3.5), and, in the delocalized regime, a bound on the second moment of the imaginary of part of the resolvent at the root of the Bethe lattice that is uniform in the imaginary part of the spectral parameter (Lemma 3.6).
+Specifically, we require concentration for fractional moments of the diagonal resolvent entries (Lemma 3.3), upper and lower bounds on the eigenvalue density (Lemma 3.5), convergence of the diagonal resolvent entries of the random regular graph to the diagonal resolvent entry at the root of the Bethe lattice (Lemma 3.4), and, in the delocalized regime, a bound on the second moment of the imaginary part of the resolvent at the root of the Bethe lattice that is uniform in the imaginary part of the spectral parameter (Lemma 3.6).
 Of these, the last is perhaps the most technically challenging.
@@
 First, we have chosen to concentrate on the case of Gaussian disorder to streamline the presentation.
-The spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] for a wide class of unbounded potentials; in the general case, it is slightly more complicated to describe, with interleaved regions of localization and delocalized determined by the disorder density.
-The techniques we present suffice to establish the analogous result for random regular graphs with corresponding potentials under the additional assumption that density decays sufficiently rapidly (for example, exponentially), so that the concentration argument for Lemma 3.3 below, based on Gaussian concentration for Lipschitz functions, can be replaced by one based on truncation and Talagrand’s concentration inequality for bounded distributions.
+The spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] for a wide class of unbounded potentials; in the general case, it is slightly more complicated to describe, with interleaved regions of localization and delocalization determined by the disorder density.
+The techniques we present suffice to establish the analogous result for random regular graphs with corresponding potentials under the additional assumption that the density decays sufficiently rapidly (for example, exponentially), so that the concentration argument for Lemma 3.3 below, based on Gaussian concentration for Lipschitz functions, can be replaced by one based on truncation and Talagrand’s concentration inequality for bounded distributions.
 Second, the requirement that $d$ be very large in Theorem 1.3 comes primarily from the analogous Bethe lattice result in [5], not our arguments that show its conclusions transfer to random regular graphs.
-The only input we require from the limiting model is an understanding of how the imaginary part of the resolvent at the root behaves in the upper half plane as the imaginary part of its spectral parameter goes to zero (see Lemma 2.3), along with a mild lower bound on $d$ (which could be improved).
+The main input we require from the limiting model is an understanding of how the imaginary part of the resolvent at the root behaves in the upper half plane as the imaginary part of its spectral parameter goes to zero (see Lemma 2.3).
 In particular, it seems likely that our methods suffice to transfer the nearly complete description of the phase diagram for the Bethe lattice at low disorder recently established in [21] to random regular graphs.
@@
 It extends by linearity to a bounded operator $\ell^{2}(\mathbb{V})$.
-Further, we let $(g_{v})_{v\in\mathbb{V}}$ be collection of i.i.d.
+Further, we let $(g_{v})_{v\in\mathbb{V}}$ be a collection of i.i.d.
 Gaussian random variables with mean zero and variance one.
@@
 $H_{d,t}=-tA^{(d)}+V^{(d)}.$
-It is well-known that $H_{d,t}$ is unbounded and self-adjoint on the domain of $V^{(d)}$; see, e.g., [25].
+It is well-known that $H_{d,t}$ is unbounded and self-adjoint on the domain of $V^{(d)}$; see, e.g., [25]. For all $z\in\mathbb{H}$, we let $R(z)$ denote the resolvent of $H_{d,t}$.
 2.2. Mobility Edge on the Bethe Lattice
@@
 For every real number $\mathfrak{L}>0$, there exists a constant $d_{0}(\mathfrak{L})>0$ such that the following holds for all $d\geq d_{0}$. Fix $g\in\mathbb{R}$ such that
-$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$ (2.1)
-and let $t=g(d\ln d)^{-1}$.
-Then there exists $\mathfrak{E}>0$ such that the solution set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$.
-Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$ and the following claims hold.
+$\sqrt{\frac{\pi}{8}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$ (2.1)
+and let
+$t=g(d\ln d)^{-1},\qquad\mathfrak{E}=\sqrt{2\ln\!\left(\frac{4g}{\sqrt{2\pi}}\right)}.$
+Then there exists $\mathfrak{M}(d,g)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathfrak{L}^{-1}$ and the following claims hold.
   1. (1)
@@
     For all $E\in\mathbb{R}$ such that $|E|<\mathfrak{M}$, there exists $c(E)>0$ such that
-    $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
+    $\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c.$
 Remark 2.2.
 In [5], the conclusion of the theorem is stated in a slightly different way, involving absolutely continuous and pure point spectrum.
-However, it is readily verified by inspecting the proof that the claims in Theorem 2.1 are proved as intermediate steps (see, e.g., [5, Lemma 2.7].) We also note that the proof in [5] is for a tree whose root degree is one less than each other vertex.
-However, as mentioned in that reference, is it straightforward to see that the same argument gives the conclusion for $\mathbb{T}_{d}$.
+However, it is readily verified by inspecting the proof that the claims in Theorem 2.1 are proved as intermediate steps (see, e.g., [5, Lemma 2.7]).
+We also note that the proof in [5] is for a tree whose root degree is one less than each other vertex.
+However, as mentioned in that reference, it is straightforward to see that the same argument gives the conclusion for $\mathbb{T}_{d}$.
 2.3. Resolvent Criteria for Localization and Delocalization
 The following lemma is the primary novelty of this paper.
-It shows that the conclusions of Theorem 2.1 regarding the resolvent of the Anderson model at the root of Bethe lattice imply the corresponding (de)localization behavior for the matrices $W_{N}$.
+It shows that the conclusions of Theorem 2.1 regarding the resolvent of the Anderson model at the root of the Bethe lattice imply the corresponding (de)localization behavior for the matrices $W_{N}$.
 The proof is given in Section 3.
@@
   2. (2)
-    If there exists $c>0$ such that $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c$, then delocalization holds at $E$.
+    If there exists $c>0$ such that $\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c$, then delocalization holds at $E$ for $W_{N}$.
 2.4. Proof of Main Result
@@
     For every $i\in[N]$, we have
-    $\sum_{j=1}^{N}\big|R_{ij}(z)\big|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (3.2)
-We also require the next lemma, on the uniform integrability of $R_{00}$.
+    $\sum_{j=1}^{N}\big|G_{ij}(z)\big|^{2}=\frac{\operatorname{Im}G_{ii}}{\eta}.$ (3.2)
+We also require the next lemma, on the uniform integrability of $\operatorname{Im}R_{00}$.
 Lemma 3.2.
-Fix $E\in\mathbb{R}$. The set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\}_{\eta\in(0,1)}$ is uniformly integrable.
+Fix $E\in\mathbb{R}$, $t>0$, and $r\in(0,1)$. The set of random variables $\{(\operatorname{Im}R_{00}(E+\mathrm{i}\eta))^{r}\}_{\eta\in(0,1)}$ is uniformly integrable.
 Proof.
-This is immediate from [5, Lemma 3.3] applied to the moments of order $(s+1)/2$. ∎
+Set $\alpha=(1+r)/2$. By [5, Lemma 3.3], applied with $L=0$ and exponent $\alpha$,
+$\sup_{\eta\in(0,1)}\mathbb{E}\!\left[\left(\big(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\big)^{r}\right)^{\alpha/r}\right]\leq\sup_{\eta\in(0,1)}\mathbb{E}\bigl[|R_{00}(E+\mathrm{i}\eta)|^{\alpha}\bigr]<\infty.$
+Because $\alpha/r>1$, the claim follows from de la Vallée–Poussin’s criterion. ∎
 3.2. Preliminary Lemmas
@@
 Lemma 3.3.
-Fix $s\in(0,1/2)\cup[1,\infty)$. There exists $C(s)>1$ such that for all $E\in\mathbb{R}$ and $\eta\geq(\ln N)^{-1}$,
+Fix $s\in(0,\infty)$, $t>0$, and $d\geq 3$. There exists $C(d,s,t)>1$ such that for all $E\in\mathbb{R}$ and $\eta\geq(\ln N)^{-1}$,
 $\mathbb{P}\!\left(\left|\frac{1}{N}\sum_{j=1}^{N}\big(\operatorname{Im}G_{jj}(z)\big)^{s}-\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s}\Big]\right|\leq\eta\right)\geq 1-CN^{-100}.$ (3.3)
-The following lemma provides upper and lower bounds on the eigenvalue density of $W_{N}$. It is proved in Section 5.
-Lemma 3.4.
-For all $M>0$, there exists $c(M)>0$ such that the following holds. Fix $E\in[-M,M]$ and $\eta\in(0,1)$, and set $I=[E-\eta,E+\eta]$. Then
-$\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\geq cN\eta\right)=1,\qquad\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\leq c^{-1}N\eta\right)=1.$ (3.4)
 The next lemma shows that the diagonal resolvent entries of the Anderson model on the $d$-regular graph converge to the resolvent entry at the root of the $d$-regular tree.
@@
 Lemma 3.4.
-Fix $d\geq 3$, $t>0$, and $z=E+\mathrm{i}\eta\in\mathbb{C}$ with $\eta>0$. There exists a coupling of $W_{N,t}$ and $H_{d,t}$ such that for all $k\in\mathbb{N}$,
-$\lim_{N\rightarrow\infty}G_{kk}(z)=R_{00}(z)$ (3.5)
+Fix $d\geq 3$, $t>0$, and $z\in\mathbb{H}$. There exists a coupling of $W_{N,t}$ and $H_{d,t}$ such that
+$\lim_{N\rightarrow\infty}G_{11}(z)=R_{00}(z)$ (3.4)
 in probability.
-Finally, we state a bound on the second moment of $\operatorname{Im}R_{00}(z)$ that is uniform in $\eta$. The proof is given in Section 7.
+The following lemma provides upper and lower bounds on the eigenvalue density of $W_{N}$. It is proved in Section 6.
 Lemma 3.5.
-Fix $E\in\mathbb{R}$ and $d\geq 20$. Suppose that there exists $c>0$ such that
-$\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
+For all $t,M>0$ and $d\geq 10$, there exists $c(d,t,M)>0$ such that the following holds. Fix $E\in[-M,M]$ and $\eta\in(0,1)$, and set $I=[E-\eta,E+\eta]$. Then
+$\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\geq cN\eta\right)=1,\qquad\lim_{N\to\infty}\mathbb{P}\!\left(|\Lambda_{I}|\leq c^{-1}N\eta\right)=1.$ (3.5)
+Finally, we state a bound on the second moment of $\operatorname{Im}R_{00}(z)$ that is uniform in $\eta$.
+The proof is given in Section 7.
+The condition $d\geq 10$ could likely be improved, but we refrain from pursuing this to avoid unnecessary complications.
+Lemma 3.6.
+Fix $E\in\mathbb{R}$, $d\geq 10$, and $t>0$. Suppose that there exists $c>0$ such that
+$\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c.$
 Then
@@
 Fix $s,\delta\in(0,1)$ and $E\in\mathbb{R}$, and suppose that
-$\lim_{\eta\rightarrow 0}R_{00}(E+\mathrm{i}\eta)=0$
+$\lim_{\eta\rightarrow 0}\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=0$
 in probability. Let $I(\eta)=[E-\eta,E+\eta]$. Then there exists $\eta_{0}(s,\delta,E)>0$ such that for all $\eta\in(0,\eta_{0})$,
@@
 We set $z=E+\mathrm{i}\eta$, where $\eta$ is a parameter that may vary throughout the proof (but never depends on $N$). By the spectral theorem, we have for all $\eta>0$ that
-$G_{jl}(z)=\sum_{k=1}^{N}\frac{u_{k}(j)u_{k}(l)}{\lambda_{k}-z},\qquad\operatorname{Im}G_{jj}(z)=\sum_{k=1}^{N}\frac{\eta\,|u_{k}(j)|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}.$ (3.8)
+$\operatorname{Im}G_{jj}(z)=\sum_{k=1}^{N}\frac{\eta\,|u_{k}(j)|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}.$ (3.8)
 For $\lambda_{k}\in I(\eta)$ we have $(\lambda_{k}-E)^{2}+\eta^{2}\leq 2\eta^{2}$, hence
-$\sum_{k\in\Lambda_{I}}\big|u_{k}(j)\big|^{2}\leq\sum_{k=1}^{N}\frac{2\eta^{2}\big|u_{k}(j)\big|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}=2\eta\operatorname{Im}G_{jj}(z).$
-Multiplying by $N/|\Lambda_{I}|$ gives
+$\sum_{\lambda_{k}\in\Lambda_{I}}\big|u_{k}(j)\big|^{2}\leq\sum_{k=1}^{N}\frac{2\eta^{2}\big|u_{k}(j)\big|^{2}}{(\lambda_{k}-E)^{2}+\eta^{2}}=2\eta\operatorname{Im}G_{jj}(z).$
+By the first bound in Lemma 3.5, we may restrict to the set where $|\Lambda_{I}|\neq 0$. Multiplying by $N/|\Lambda_{I}|$ gives
 $P_{I}(j)\leq\frac{2N\eta}{|\Lambda_{I}|}\,\operatorname{Im}G_{jj}(z),$ (3.9)
@@
 $\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\!\left(\frac{s}{2}\right)\leq c_{1}^{-s/2}\left(\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]+\eta\right)\right)=1.$ (3.11)
-Next, recalling (3.5) and using $|G_{11}(z)|\leq 1/\eta$ (from (3.1)), the dominated convergence theorem yields
+Next, recalling (3.4) and using $|G_{11}(z)|\leq 1/\eta$ (from (3.1)), uniform integrability yields
 $\lim_{N\to\infty}\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]=\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(z)\big)^{s/2}\Big].$ (3.12)
@@
 $\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{s/2}\Big]\leq\frac{c_{1}^{s/2}\delta}{4}.$
-Combining this with (3.11) and taking $\eta_{0}<c_{1}^{s/2}\delta/2$ completes the proof. ∎
+Combining this with (3.11) and taking $\eta_{0}<\min\{c_{2},\,c_{1}^{s/2}\delta/2\}$ completes the proof. ∎
 Proposition 3.8.
@@
 Proof.
-Let $\varepsilon\in(0,1)$ and $p,q>1$ be parameters such that $1/p+1/q=1$. By Hölder’s inequality,
-$N=\sum_{j=1}^{N}P_{I}(j)=\frac{1}{N}\sum_{j=1}^{N}(NP_{I}(j))^{\varepsilon}(NP_{I}(j))^{1-\varepsilon}\leq\Bigl(N^{\varepsilon p}Q_{I}(\varepsilon p)\Bigr)^{1/p}\,\Bigl(N^{(1-\varepsilon)q}Q_{I}((1-\varepsilon)q)\Bigr)^{1/q}.$
+By the first bound in Lemma 3.5, we may restrict to the set where $|\Lambda_{I}|\neq 0$. Let $\varepsilon\in(0,1)$ and $p,q>1$ be parameters such that $1/p+1/q=1$. By Hölder’s inequality,
+$N=\sum_{j=1}^{N}P_{I}(j)=\frac{1}{N}\sum_{j=1}^{N}(NP_{I}(j))^{\varepsilon}(NP_{I}(j))^{1-\varepsilon}\leq\Bigl(N^{\varepsilon p}Q_{I}(\varepsilon p)\Bigr)^{1/p}\,\Bigl(N^{(1-\varepsilon)q}Q_{I}((1-\varepsilon)q)\Bigr)^{1/q}.$ (3.13)
 With $s\in(0,1)$, we fix
-$\varepsilon=\frac{4}{4-s},\qquad p=2-\frac{s}{2}.$
-Then $\varepsilon p=s/2$, $(1-\varepsilon)q=2$, and (p/q)=1s/2(p/q)=1-s/2, yielding
+$\varepsilon=\frac{s}{4-s},\qquad p=2-\frac{s}{2}.$
+Then $\varepsilon p=s/2$, $(1-\varepsilon)q=2$, and (p/q)=1s/2(p/q)=1-s/2. Raising both sides of (3.13) to the $p$th power yields
 $1\leq Q_{I}(s/2)\cdot Q_{I}(2)^{\,1-s/2}=Q_{I}(s/2)\cdot Q_{I}^{\,1-s/2}.$
@@
 Proof.
-Using (3.9), we have
+Let $c_{1}(E)>0$ be such that (3.5) yields $\lim_{N\to\infty}\mathbb{P}(|\Lambda_{I}|\geq c_{1}N\eta)=1$. We restrict to the event where $|\Lambda_{I}|\geq c_{1}N\eta$. Using (3.9), we have
 $Q_{I}=\frac{1}{N}\sum_{j=1}^{N}P_{I}(j)^{2}\leq\frac{4N^{2}\eta^{2}}{|\Lambda_{I}|^{2}}\cdot\frac{1}{N}\sum_{j=1}^{N}(\operatorname{Im}G_{jj}(z))^{2}.$ (3.14)
-Using the concentration bound (3.3) for $r=2$, we obtain
-$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\leq 4c_{1}^{-2}\cdot\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]+\eta\right)=1.$ (3.14)
-From (3.5) and together with the deterministic bound $|G_{11}|\leq\eta^{-1}$ (from (3.1)), the dominated convergence theorem gives
+Using the concentration bound (3.3) for $s=2$, we obtain
+$\lim_{N\to\infty}\mathbb{P}\left(Q_{I}\leq 4c_{1}^{-2}\left(\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]+\eta\right)\right)=1.$ (3.15)
+From (3.4), together with the deterministic bound $|G_{11}|\leq\eta^{-1}$ (from (3.1)), uniform integrability gives
 $\lim_{N\to\infty}\mathbb{E}\Big[\big(\operatorname{Im}G_{11}(z)\big)^{2}\Big]=\mathbb{E}\Big[\big(\operatorname{Im}R_{00}(z)\big)^{2}\Big].$ (3.16)
@@
 4.1. Preliminary Reduction
-For all symmetric $N\times N$ matrices $M$, $v\in\mathbb{R}^{N}$, $z\in\mathbb{H}$ and $s>0$, define
+For all symmetric $N\times N$ matrices $M$, vectors $v\in\mathbb{R}^{N}$, $z\in\mathbb{H}$ and $s>0$, define
 $F_{s}(z,v,M)=\frac{1}{N}\sum_{i=1}^{N}\big(\operatorname{Im}S_{ii}(z)\big)^{s},\qquad S=\big(M+\mathrm{diag}(v_{1},\dots,v_{N})-z\big)^{-1}.$
-We frequently abbreviate $F_{s}(z)=F_{s}(z,v,M)$, and we set $\tilde{F}_{s}=F_{s}(z,V_{N},-tA_{N})$, so that with this choice of arguments, $S(z)=G(z)$.
+We frequently abbreviate $F_{s}(z)=F_{s}(z,v,M)$, and we write $v_{N}=(g_{1},\dots,g_{N})$ and set $\tilde{F}_{s}=F_{s}(z,v_{N},-tA_{N})$, so that with this choice of arguments, $S(z)=G(z)$.
 We establish the following preliminary concentration estimate in Section 4.4 below, after some preliminary estimates on $F_{s}(z)$ (and a truncated counterpart).
-The first exponential term comes from Gaussian concentration in $V_{N}$, while the second comes from a coarse Azuma–Freedman bound for the graph randomness in $A_{N}$.
+The first exponential term comes from Gaussian concentration in $V_{N}$, while the second comes from a coarse Azuma–Hoeffding bound for the graph randomness in $A_{N}$.
 Lemma 4.1.
@@
     $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u\right)\leq 2\exp\left(-\frac{N\eta^{2s+2}}{8s^{2}}u^{2}\right)+2\exp\left(-\frac{N\eta^{2s+2}}{Ct^{2}s^{2}}u^{2}\right).$ (4.1)
-    Further, the same bound holds for the corresponding averages with $\operatorname{Im}S_{ii}(z)$ replaced by $\operatorname{Re}S_{ii}(z)$.
+    Further, when $s=1$, the same bound holds for the corresponding averages with $\operatorname{Im}S_{ii}(z)$ replaced by $\operatorname{Re}S_{ii}(z)$.
   2. (2)
-    Fix $s\in(0,1/2)$. Then for every $u>0$ and every $\delta\in(0,\eta^{-1}]$,
-    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u+2\delta^{s}\right)\leq 2\exp\left(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\right)+2\exp\left(-\frac{N\eta^{4}}{Ct^{2}s^{2}}u^{2}\right).$ (4.2)
+    Fix $s\in(0,1)$. Then for every $u>0$ and every $\delta\in(0,\eta^{-1}]$,
+    $\mathbb{P}\left(\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\geq u+2\delta^{s}\right)\leq 2\exp\left(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\right)+2\exp\left(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\right).$ (4.2)
 Lemma 3.3 is an immediate consequence of Lemma 4.1.
 Proof of Lemma 3.3.
-We begin with the case $s\geq 1$.
-This follows from setting $u=\eta$ in (4.1).
-Next, for $s\in(0,1/2)$, we may suppose that $\eta\leq 1$ (by decreasing $\eta$ if necessary).
-Then the estimate follows from setting $u=\eta/2$ and $\delta=(\eta/4)^{1/s}$ in (4.2).
-∎
+We begin with the case $s\geq 1$. This follows from setting $u=\eta$ in (4.1). Next, for $s\in(0,1)$, if $\eta\geq 1$, then $0\leq\tilde{F}_{s}(z)\leq\eta^{-s}\leq 1$, and therefore
+$\Big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\Big|\leq 1\leq\eta$
+deterministically, so the claim is trivial. Thus we may assume $\eta\leq 1$. Then the estimate follows from setting $u=\eta/2$ and $\delta=(\eta/4)^{1/s}$ in (4.2). ∎
 4.2. Gaussian Concentration
-The following Gaussian concentration inequality is standard (see, e.g., [19, Theorem 5.6]).
+The following Gaussian concentration inequality is standard (see, e.g., [20, Theorem 5.6]).
 We say a function $f\colon\mathbb{R}^{N}\rightarrow\mathbb{R}$ is $L$-Lipschitz if $|f(x)-f(y)|\leq L\|x-y\|_{2}$ for all $x,y\in\mathbb{R}^{N}$.
@@
 For all $t>0$,
-$\mathbb{P}\left(f(X)-\mathbb{E}\big[f(X)\big]\geq t\right)\leq\exp\left(-\frac{t^{2}}{2L^{2}}\right).$
+$\mathbb{P}\left(\Big|f(X)-\mathbb{E}\big[f(X)\big]\Big|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2L^{2}}\right).$
 We use the notation $\partial_{k}F_{s}(z)$ to denote the partial derivative in $v_{k}$, and we write $\nabla F_{s}(z)$ for the gradient with respect to $v\in\mathbb{R}^{N}$.
@@
 $\partial_{k}S_{ii}(z)=-S_{ik}(z)S_{ki}(z).$ (4.3)
-We now derive some gradient bounds that are necessary to apply Lemma 4.2. These bounds are uniform in the matrix $M$, and hence may be applied conditionally on $A_{N}$.
+We now derive some Lipschitz bounds that are necessary to apply Lemma 4.2. These bounds are uniform in the matrix $M$, and hence may be applied conditionally on $A_{N}$.
 Lemma 4.3.
@@
 $\partial_{k}F_{s}(z)=\frac{1}{N}\sum_{i=1}^{N}s\big(\operatorname{Im}S_{ii}(z)\big)^{s-1}\partial_{k}\big(\operatorname{Im}S_{ii}(z)\big).$
-Using (4.3) and the bound $\operatorname{Im}S_{ii}(z)\leq\eta^{-1}$ from (3.1),
+Using (4.3) and the bound $\operatorname{Im}S_{ii}(z)\leq\eta^{-1}$ (analogous to (3.1)),
 $\big|\partial_{k}F_{s}(z)\big|\leq\frac{s}{N}\sum_{i=1}^{N}(\operatorname{Im}S_{ii})^{s-1}|S_{ik}(z)|^{2}\leq\frac{s}{N}\cdot\eta^{-(s-1)}\sum_{i=1}^{N}|S_{ik}(z)|^{2}.$
-By the Ward identity (3.2),
-$\sum_{i=1}^{N}|G_{ik}(z)|^{2}=\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
+By the Ward identity (3.2) for $S$,
+$\sum_{i=1}^{N}|S_{ik}(z)|^{2}=\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
 Therefore,
@@
 which implies the claim. ∎
-For every $s\in(0,1/2)$, the map $x\mapsto x^{s}$ is not Lipschitz near $x=0$. We therefore introduce a truncation at level $\delta>0$ and define, for any $\delta\in(0,\eta^{-1}]$,
+For every $s\in(0,1)$, the map $x\mapsto x^{s}$ is not Lipschitz near $x=0$. We therefore introduce a truncation at level $\delta>0$ and define, for any $\delta\in(0,\eta^{-1}]$,
 $F_{s,\delta}(z)=\frac{1}{N}\sum_{i=1}^{N}\big(\operatorname{Im}S_{ii}(z)\vee\delta\big)^{s},$
@@
 Lemma 4.4.
-Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then
-$\|\nabla F_{s,\delta}(z)\|_{2}\leq\frac{s}{\eta^{2}\sqrt{N}}\ \delta^{s-1}.$
+Fix $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. Then, viewed as a function of $v\in\mathbb{R}^{N}$, the map $F_{s,\delta}(z)$ is Lipschitz and
+$\operatorname{Lip}(F_{s,\delta}(z))\leq\frac{s}{\eta^{2}\sqrt{N}}\ \delta^{s-1}.$
 Proof.
-Let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$. By the chain rule and (4.3),
-$\partial_{k}F_{s,\delta}(z)=\frac{1}{N}\sum_{i=1}^{N}\varphi_{\delta}^{\prime}\big(\operatorname{Im}S_{ii}(z)\big)\partial_{k}\big(\operatorname{Im}S_{ii}(z)\big),\qquad\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{1}{N}\sum_{i=1}^{N}\|\varphi_{\delta}^{\prime}\|_{\infty}\big|\partial_{k}\operatorname{Im}S_{ii}(z)\big|.$
+Let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\varphi_{\delta}$ is Lipschitz on $0,\infty)$, with
+$\operatorname{Lip}(\varphi_{\delta})\leq s\delta^{s-1}.$
+Fix $k\in\{1,\dots,N\}$ and $h\in\mathbb{R}$. Then
+$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|$ $\leq\frac{1}{N}\sum_{i=1}^{N}\big|\varphi_{\delta}(\operatorname{Im}S_{ii}(v+he_{k};z))-\varphi_{\delta}(\operatorname{Im}S_{ii}(v;z))\big|$
+$\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S_{ii}(v+he_{k};z)-\operatorname{Im}S_{ii}(v;z)\big|.$
+Since $v\mapsto\operatorname{Im}S_{ii}(v;z)$ is $C^{1}$, the fundamental theorem of calculus yields
+$\big|\operatorname{Im}S_{ii}(v+he_{k};z)-\operatorname{Im}S_{ii}(v;z)\big|\leq|h|\,\sup_{t\in[0,1]}\big|\partial_{k}(\operatorname{Im}S_{ii})(v+the_{k};z)\big|.$
+Hence
+$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|\leq\frac{s\delta^{s-1}|h|}{N}\sup_{t\in[0,1]}\sum_{i=1}^{N}\big|\partial_{k}(\operatorname{Im}S_{ii})(v+the_{k};z)\big|.$
 Note that (4.3) implies $|\partial_{k}(\operatorname{Im}S_{ii})|\leq|S_{ik}(z)|^{2}$. Then combining this estimate with (3.1) and (3.2), we find
-$\big|\partial_{k}F_{s,\delta}(z)\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}|S_{ik}(z)|^{2}\leq\frac{s\delta^{s-1}}{N}\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{s\delta^{s-1}}{N\eta^{2}}.$
-Summing over $k$ gives
-$\|\nabla F_{s,\delta}(z)\|_{2}^{2}\leq N\left(\frac{s}{N\eta^{2}}\delta^{s-1}\right)^{2}=\frac{s^{2}}{N\eta^{4}}\delta^{2s-2}.$
-Taking square roots yields the claim. ∎
+$\sum_{i=1}^{N}\big|\partial_{k}(\operatorname{Im}S_{ii})(z)\big|\leq\sum_{i=1}^{N}|S_{ik}(z)|^{2}\leq\frac{\operatorname{Im}S_{kk}(z)}{\eta}\leq\frac{1}{\eta^{2}}.$
+Therefore
+$\big|F_{s,\delta}(v+he_{k};z)-F_{s,\delta}(v;z)\big|\leq\frac{s\delta^{s-1}}{N\eta^{2}}|h|.$
+So the Lipschitz constant in each coordinate direction is at most $\frac{s\delta^{s-1}}{N\eta^{2}}$.
+For arbitrary $v,w\in\mathbb{R}^{N}$, varying the coordinates one at a time gives
+$|F_{s,\delta}(v;z)-F_{s,\delta}(w;z)|\leq\frac{s\delta^{s-1}}{N\eta^{2}}\,\|v-w\|_{1}\leq\frac{s\delta^{s-1}}{\eta^{2}\sqrt{N}}\,\|v-w\|_{2}.$
+This proves the Lipschitz bound. ∎
 4.3. Dependence on the Graph
@@
 Distinct pairs of $\omega$ may project to the same unordered pair $\{i,j\}$, in which case this produces multiple edges.
-We write $A(\omega)$ for the adjacency matrix of $A(\omega)$; if a loop occurs, it contributes $2$ to the corresponding diagonal entry, so that every row sum is equal to $d$.
+We write $A(\omega)$ for the adjacency matrix of $\mathcal{G}(\omega)$; if a loop occurs, it contributes $2$ to the corresponding diagonal entry, so that every row sum is equal to $d$.
 Let $\widehat{A}_{N}=A(\omega)$ for $\omega$ chosen uniformly from the set of all pairings of $W$, and let
@@
 $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{S}_{N})=\exp\left(\frac{1-d^{2}}{4}\right)$
-(see, e.g., [35, equation (5)]), and therefore there exists a constant $p_{d}>0$ such that
+(see, e.g., [36, equation (5)]), and therefore there exists a constant $p_{d}>0$ such that
 $\mathbb{P}(\mathcal{S}_{N})\geq p_{d}\qquad\text{for all }N.$ (4.5)
-We also record the Azuma–Hoeffding martingale concentration estimate that we use. See, e.g., [19, Chapter 2].
+We also record the Azuma–Hoeffding martingale concentration estimate that we use. See, e.g., [20, Chapter 2].
 Lemma 4.5.
@@
     For every $s\geq 1$,
-    $\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{8s|t|}{N\eta^{s+1}}.$
+    $\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{32s|t|}{N\eta^{s+1}}.$
   2. (2)
-    For every $s\in(0,1/2)$ and every $\delta\in(0,\eta^{-1}]$,
-    $\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
+    For every $s\in(0,1)$ and every $\delta\in(0,\eta^{-1}]$,
+    $\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{32s|t|}{N\eta^{2}}\,\delta^{s-1}.$
 Proof.
 Set $\Delta=-t(A^{\prime}-A)$.
-Since $A$ and $A^{\prime}$ differ by one switching of two pairs, the matrix $A^{\prime}-A$ has at most eight nonzero entries, each of magnitude at most $1$.
+Since $A$ and $A^{\prime}$ differ by one switching of two pairs, the matrix $A^{\prime}-A$ is supported on the at most four vertices incident to the switched half-edges.
 In particular,
-$\mathrm{rank}(\Delta)\leq 4,\qquad\|\Delta\|\leq 2|t|.$
+$\mathrm{rank}(\Delta)\leq 4.$
+Moreover, $A^{\prime}-A$ has at most eight nonzero entries, each of magnitude at most $2$ (since loops contribute $2$ to the diagonal).
+Hence every row and every column of $A^{\prime}-A$ has $\ell^{1}$-norm at most $8$, so by the Schur test,
+$\|\Delta\|\leq 8|t|.$
 By the resolvent identity,
-$S^{\prime}-S=S\Delta S^{\prime}.$
+$S^{\prime}-S=-S\Delta S^{\prime}.$
 Hence
-$\mathrm{rank}(S^{\prime}-S)\leq 4,\qquad\|S^{\prime}-S\|\leq\|S\|\,\|\Delta\|\,\|S^{\prime}\|\leq\frac{2|t|}{\eta^{2}}.$
+$\mathrm{rank}(S^{\prime}-S)\leq 4,\qquad\|S^{\prime}-S\|\leq\|S\|\,\|\Delta\|\,\|S^{\prime}\|\leq\frac{8|t|}{\eta^{2}}.$
 Therefore
-$\frac{1}{N}\sum_{i=1}^{N}\big|S^{\prime}_{ii}-S_{ii}\big|\leq\frac{1}{N}\|S^{\prime}-S\|_{*}\leq\frac{4}{N}\|S^{\prime}-S\|\leq\frac{8|t|}{N\eta^{2}},$
+$\frac{1}{N}\sum_{i=1}^{N}\big|S^{\prime}_{ii}-S_{ii}\big|\leq\frac{1}{N}\|S^{\prime}-S\|_{*}\leq\frac{4}{N}\|S^{\prime}-S\|\leq\frac{32|t|}{N\eta^{2}},$
 where $\|\cdot\|_{*}$ denotes the trace norm.
 For $s\geq 1$, the map $x\mapsto x^{s}$ is $s\eta^{-(s-1)}$-Lipschitz on $0,\eta^{-1}]$, so
-$\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{s}{N}\eta^{-(s-1)}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{s+1}}.$
+$\big|F_{s}(z,v,-tA)-F_{s}(z,v,-tA^{\prime})\big|\leq\frac{s}{N}\eta^{-(s-1)}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{32s|t|}{N\eta^{s+1}}.$
 This proves the first claim.
-Next, let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Then $\|\varphi_{\delta}^{\prime}\|_{\infty}\leq s\delta^{s-1}$, and therefore
-$\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|\leq\frac{8s|t|}{N\eta^{2}}\,\delta^{s-1}.$
+Next, let $\varphi_{\delta}(x)=(x\vee\delta)^{s}$. Since $s\in(0,1)$, the map $x\mapsto x^{s}$ is concave on $(0,\infty)$, and therefore $\varphi_{\delta}$ is Lipschitz on $0,\infty)$ with
+$\operatorname{Lip}(\varphi_{\delta})\leq s\delta^{s-1}.$
+Since $\operatorname{Im}S_{ii},\operatorname{Im}S^{\prime}_{ii}\geq 0$, it follows that
+$\big|F_{s,\delta}(z,v,-tA)-F_{s,\delta}(z,v,-tA^{\prime})\big|$ $\leq\frac{1}{N}\sum_{i=1}^{N}\big|\varphi_{\delta}(\operatorname{Im}S^{\prime}_{ii})-\varphi_{\delta}(\operatorname{Im}S_{ii})\big|$
+$\leq\frac{s\delta^{s-1}}{N}\sum_{i=1}^{N}\big|\operatorname{Im}S^{\prime}_{ii}-\operatorname{Im}S_{ii}\big|$
+$\leq\frac{32s|t|}{N\eta^{2}}\delta^{s-1}.$
 This proves the second claim. ∎
@@
 whenever $A$ and $A^{\prime}$ arise from two pairings that differ by one switching of two pairs. Then for every $u>0$,
-$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)\leq 2\exp\left(-\frac{u^{2}}{CNdL^{2}}\right).$
+$\mathbb{P}\Big(\big|f(A_{N})-\mathbb{E}\big[f(A_{N})\big]\big|\geq u\Big)\leq 2\exp\left(-\frac{u^{2}}{CNL^{2}}\right).$
 Proof.
@@
 Lemma 4.9.
-Fix $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
-$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$ (4.11)
+Fix $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. There exists $C(d)>0$ such that for every $u>0$,
+$\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big)\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big).$ (4.11)
 Proof.
@@
 where $\mathbb{E}_{V}$ denotes expectation in the Gaussian potential only. By Lemma 4.7, $g_{s,\delta}$ is $L$-Lipschitz under one simple switching with
-$L=\frac{8s|t|}{N\eta^{2}}\delta^{s-1}.$
+$L=\frac{32s|t|}{N\eta^{2}}\delta^{s-1}.$
 Hence Lemma 4.8 implies
@@
 Lemma 4.10.
-Let $s\in(0,1/2)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
+Let $s\in(0,1)$ and $\delta\in(0,\eta^{-1}]$. Then for every $u>0$,
 $\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\big[\tilde{F}_{s}(z)\big]\big|\geq u+2\delta^{s}\Big)\leq\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\big[\tilde{F}_{s,\delta}(z)\big]\big|\geq u\Big).$
@@
 By Lemma 4.7, $g_{s}$ is $L$-Lipschitz under one simple switching with
-$L=\frac{8s|t|}{N\eta^{s+1}}.$
+$L=\frac{32s|t|}{N\eta^{s+1}}.$
 Hence Lemma 4.8 implies
@@
 Combining the last two bounds by the triangle inequality proves (4.1). The claim about the corresponding averages involving $\operatorname{Re}S_{ii}(z)$ is completely analogous.
-Next, by Lemma 4.10 and Lemma 4.9, for any $s\in(0,1/2)$, $\delta\in(0,\eta^{-1}]$, and $u>0$,
+Next, by Lemma 4.10 and Lemma 4.9, for any $s\in(0,1)$, $\delta\in(0,\eta^{-1}]$, and $u>0$,
 $\mathbb{P}\Big(\big|\tilde{F}_{s}(z)-\mathbb{E}\tilde{F}_{s}(z)\big|\geq u+2\delta^{s}\Big)\leq\mathbb{P}\Big(\big|\tilde{F}_{s,\delta}(z)-\mathbb{E}\tilde{F}_{s,\delta}(z)\big|\geq u\Big)$
-$\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Cdt^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big),$
+$\leq 2\exp\Big(-\frac{N\eta^{4}}{8s^{2}}\delta^{2-2s}\,u^{2}\Big)+2\exp\Big(-\frac{N\eta^{4}}{Ct^{2}s^{2}}\delta^{2-2s}\,u^{2}\Big),$
 which is (4.2). ∎
-5. Bounds for the Eigenvalue Counting Function
-This section is dedicated to the proof of Lemma 3.4.
-To obtain this result, we first show that the Stieltjes transform of $W_{N}$ converges to the Stieltjes transform of the density of the local states measure of $H_{d,t}$ at the root of the Bethe lattice.
-For the Anderson model on the random regular graph with a bounded potential, an analogous result for the integrated density of states was proved by Geisinger [23].
-We adapt this result to $W_{N}$ by a truncation argument, then upgrade the convergence in means to convergence in probability using the $s=1$ case of Lemma 4.1.
-Finally, since convergence of Stieltjes transforms implies convergence of the associated spectral measures, the desired lower bound on the eigenvalue counting bound follows from a lower bound on the integrated density of states of $H_{d,t}$, which we show in Appendix B, and the upper bound follows from Wegner’s estimate.
-5.1. Notation
-For all $x\in[N]$, we define the spectral measure of $W_{N}$ at $x$ by
-$\mu_{N,x}(B)=\langle\delta_{x},\mathbbm{1}_{B}(W_{N})\delta_{x}\rangle,$
-where $B\subset\mathbb{R}$ is any Borel set, and $\mathbbm{1}_{B}(W_{N})$ is defined in the usual way (using the spectral theorem). We define
-$m_{N}(z)=\frac{1}{N}\operatorname{Tr}G_{N}(z),$
-which is the Stieltjes transform of the empirical eigenvalue distribution of $W_{N}$.
-Next, bcause $H_{d,t}$ is self-adjoint, the spectral theorem implies the existence of a random projection-valued measure $\mu$ such that $H=\int\lambda\,d\mu(\lambda)$.
-The spectral measure at the root is given by
-$\mu_{0}(S)=\langle\delta_{0},\mu(S)\delta_{0}\rangle$
-for all Borel sets $S\subset\mathbb{R}$, and the density of states measure $\nu$ is given by
-$\nu(S)=\mathbb{E}\big[\mu_{0}(S)\big].$
-For all $z\in\mathbb{H}$, the Stieltjes transform of $\nu$ is
-$m_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,d\nu(\lambda).$
-5.2. Preliminaries
-We first record an integration by parts bound for signed distribution functions.
-Lemma 5.1.
-Let $\mu_{1}$ and $\mu_{2}$ be finite Borel measures on $\mathbb{R}$, and set
-$F(t)=\mu_{1}\big((-\infty,t]\big)-\mu_{2}\big((-\infty,t]\big).$
-If $f\in C^{1}(\mathbb{R})$ satisfies $f(\lambda)\to 0$ as $|\lambda|\to\infty$ and $f^{\prime}\in L^{1}(\mathbb{R})$, then
-$\Big|\int_{\mathbb{R}}f\,d\mu_{1}-\int_{\mathbb{R}}f\,d\mu_{2}\Big|\leq\|F\|_{\infty}\int_{\mathbb{R}}\big|f^{\prime}(\lambda)\big|\,d\lambda.$
-Proof.
-We apply integration by parts to the signed measure $\mu_{1}-\mu_{2}$ (see, e.g., [21, Theorem 3.36]), yielding
-$\int_{\mathbb{R}}f\,d(\mu_{1}-\mu_{2})=-\int_{\mathbb{R}}F(\lambda)f^{\prime}(\lambda)\,d\lambda,$
-then bound the right-hand side by by $\|F\|_{\infty}\int|f^{\prime}|$. ∎
-Next we establish an estimate on the stability of resolvents under perturbations.
-Lemma 5.2.
-Let $H$ and $\widetilde{H}$ be symmetric $N\times N$ matrices, and set $\Delta=\widetilde{H}-H$. Then for any $z\in\mathbb{H}$,
-$\Big|\frac{1}{N}\operatorname{Tr}(\widetilde{H}-z)^{-1}-\frac{1}{N}\operatorname{Tr}(H-z)^{-1}\Big|\leq\frac{1}{N\eta^{2}}\operatorname{Tr}|\Delta|.$ (5.1)
-Proof.
-By the definition of a resolvent,
-$(\widetilde{H}-z)^{-1}-(H-z)^{-1}=-(H-z)^{-1}\Delta(\widetilde{H}-z)^{-1}.$
-Taking traces and using the trace norm bound $|\operatorname{Tr}(ABC)|\leq\|A\|\|C\|\operatorname{Tr}|B|$ yields
-$\Big|\operatorname{Tr}\big((\widetilde{H}-z)^{-1}-(H-z)^{-1}\big)\Big|\leq\big\|(H-z)^{-1}\big\|\big\|(\widetilde{H}-z)^{-1}\big\|\operatorname{Tr}|\Delta|\leq\eta^{-2}\operatorname{Tr}|\Delta|.$
-Dividing by $N$ yields the claim. ∎
-Given a probability density $\tilde{\rho}$, we let $\widetilde{W}_{N}=-tA_{N}+\widetilde{V}_{N}$ and $H_{d,t}=-tA^{(d)}+\widetilde{V}^{(d)}$, where $\widetilde{V}_{N}$ and $\widetilde{V}^{(d)}$ are defined analogously to $V_{N}$ and $V^{(d)}$ with the Gaussian potential variables replaced by i.i.d.
-variables with density $\tilde{\rho}$.
-We set $\tilde{\mu}_{N,x}$ and $\tilde{\mu}_{0}$ denote the corresponding spectral measures, defined analogously to $\mu_{N,x}$ and $\mu_{0}$.
-The following theorem is [23, Theorem 2.3].
-Theorem 5.3.
-Suppose that $\|\tilde{\rho}\|$$\infty$ and there exists $\rho_{0}>0$ such that $\mathrm{supp}(\tilde{\rho})\subset(-\rho_{0},\rho_{0})$.
-Let $\widetilde{\mathbb{E}}$ denote expectation with respect to the potentials $\tilde{V}_{N}$ and $\tilde{V}^{(d)}$, conditional on $A_{N}$.
-Define the event
-$\mathcal{A}_{N}=\left\{\sup_{s\in\mathbb{R}}\Bigg|\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}\big[\tilde{\mu}_{N,x}\big((-\infty,s]\big)\big]-\widetilde{\mathbb{E}}\big[\tilde{\mu}_{0}\big((-\infty,s]\big)\big]\Bigg|\leq\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.2)
-which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{A}_{N})=1$.
-Define the disorder-averaged measures
-$\tilde{\nu}^{*}_{N}=\frac{1}{N}\sum_{x\in[N]}\widetilde{\mathbb{E}}[\tilde{\mu}_{N,x}],\qquad\tilde{\nu}^{*}=\widetilde{\mathbb{E}}[\tilde{\mu}_{0}].$
-The corresponding Stieltjes measures are
-$m^{*}_{N}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}_{N}(d\lambda),\qquad m^{*}_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,\tilde{\nu}^{*}(d\lambda).$
-Using Theorem 5.3, we deduce convergence of the disorder-averaged Stieltjes transform in the bounded potential setting.
-Proposition 5.4.
-Retain the hypotheses and notation of Theorem 5.3. For every $z\in\mathbb{H}$, define the event
-$\mathcal{B}_{N}(z)=\left\{\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|\leq\frac{\pi}{\eta}\cdot\frac{5\pi\|\tilde{\rho}\|_{\infty}(2\sqrt{d-1}+\rho_{0})}{\operatorname{log}_{d-1}N}\right\},$ (5.3)
-which depends only on the randomness in $A_{N}$. Then $\lim_{N\rightarrow\infty}\mathbb{P}(\mathcal{B}_{N}(z))=1$.
-Proof.
-Define $f_{z}(\lambda)=(\lambda-z)^{-1}$. Then $f^{\prime}_{z}(\lambda)=-(\lambda-z)^{-2}$ and
-$\int_{\mathbb{R}}|f^{\prime}_{z}(\lambda)|\,d\lambda=\int_{\mathbb{R}}\frac{d\lambda}{(\lambda-E)^{2}+\eta^{2}}=\frac{\pi}{\eta}.$ (5.4)
-Apply Lemma 5.1 with $\mu_{1}=\tilde{\nu}^{*}_{N}$, $\mu_{2}=\tilde{\nu}^{*}$, and $f=f_{z}$. Further, define
-$F_{N}(s)=\tilde{\nu}^{*}_{N}\big((-\infty,s]\big)\qquad F_{\mathbb{T}}(s)=\tilde{\nu}^{*}\big((-\infty,s]\big).$
-Then, using (5.4),
-$\big|m^{*}_{N}(z)-m^{*}_{\mathbb{T}}(z)\big|=\Big|\int f_{z}\,d\tilde{\nu}^{*}_{N}-\int f_{z}\,d\tilde{\nu}^{*}\Big|\leq\|F_{N}-F_{\mathbb{T}}\|_{\infty}\int|f^{\prime}_{z}|=\|F_{N}-F_{\mathbb{T}}\|_{\infty}\cdot\frac{\pi}{\eta}.$
-Now insert the bound (5.2) to complete the proof. ∎
-Next, we state an upper bound on the integrated density of states of $H_{d,t}$, which is a consequence of a well-known estimate of Wegner [8, (5.27)].
-Lemma 5.5.
-Fix $g,\mathfrak{L}>0$ such that (2.1) holds. For every $d\geq 3$, there exists a constant $C(d,\mathfrak{L})>0$ such that for all intervals $I\subset\mathbb{R}$, we have $\nu(I)\leq C|I|$.
-Finally, we state a lower bound that complements the previous Wegner estimate. It is proved in Appendix B.
-Lemma 5.6.
-Fix $g,\mathfrak{L}>0$ such that (2.1) holds, and fix $D>0$.
-For every $d\geq 3$, there exists a constant $c(d,D,\mathfrak{L})>0$ such that for all intervals $I\subset[-D,D]$, we have $\nu(I)\geq c|I|$.
-5.3. Global Law for Gaussian Disorder
-To pass from Gaussian disorder to bounded support, we introduce a coupled truncation scheme. Fix $L\geq 1$. Let $X\sim N(0,1)$ and let $U^{(L)}\sim\mathrm{Unif}[-L,L]$ be independent of $V$. Define
-$X^{(L)}=X\mathbbm{1}_{\{|X|\leq L\}}+U^{(L)}\mathbbm{1}_{\{|X|>L\}}.$ (5.5)
-Let $(X_{x},U_{x}^{(L)})_{x\in[N]}$ be i.i.d. with law $(X,U)$, define $X_{x}^{(L)}$ by (5.5), and set
-$V_{N}^{(L)}=\mathrm{diag}(X_{1}^{(L)},\dots,X_{N}^{(L)}),\qquad W_{N}^{(L)}=-tA_{N}+V_{N}^{(L)}.$
-Define the truncated Stieltjes transform
-$m_{N}^{(L)}(z)=\frac{1}{N}\operatorname{Tr}(W_{N}^{(L)}-z)^{-1},$
-and define similarly $H_{t,d}^{(L)}=-tA^{(d)}+V^{(d),(L)}$ on the Bethe lattice and
-$m_{\mathbb{T}}^{(L)}(z)=\mathbb{E}\big[\langle\delta_{0},(H_{d,t}^{(L)}-z)^{-1}\delta_{0}\rangle\big].$
-Proposition 5.7.
-Fix $z\in\mathbb{H}$. Then $\lim_{N\rightarrow\infty}m_{N}(z)=m_{\mathbb{T}}(z)$ in probability.
-Proof.
-We proceed through a sequence of reductions that isolate truncation and approximation errors.
-We first apply Lemma 5.2 with $H=W_{N}$ and $\widetilde{H}=W_{N}^{(L)}$.
-Here $\Delta$ is diagonal with entries $V_{N,i}^{(L)}-V_{N,i}$, hence
-$|m_{N}(z)-m_{N}^{(L)}(z)|\leq\frac{1}{\eta^{2}}\cdot\frac{1}{N}\sum_{i=1}^{N}|V_{N,i}^{(L)}-V_{N,i}|.$
-Taking expectation and using that the entries of each potential are i.i.d., we find
-$\sup_{N\geq 1}\mathbb{E}\Big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\Big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[|V_{N,1}^{(L)}-V_{N,1}|\big]\leq\frac{1}{\eta^{2}}\mathbb{E}\big[\big(|g_{1}|+L\big)\mathbbm{1}_{\{|g_{1}|>L\}}\big].$ (5.6)
-For each $L$ and $i\in[N]$, the law of $V^{(L)}_{N,i}$ is supported on [L,L][-L,L] and has a bounded density.
-Therefore Proposition 5.4 applies to $W_{N}^{(L)}$ and yields, for fixed $L$ and fixed $z$,
-$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}^{(L)}(z)\big]=m_{\mathbb{T}}^{(L)}(z).$ (5.7)
-We now remove the truncation on the tree.
-Because the tree has infinitely many sites, we do not compare $m_{\mathbb{T}}^{(L)}$ and $m_{\mathbb{T}}$ by summing over sites of $\mathbb{T}_{d}$.
-Instead, we use strong resolvent convergence.
-Let $C_{c}(\mathbb{T})$ denote finitely supported functions on $T_{d}$.
-It is well known that $C_{c}(\mathbb{T})$ is a core for the multiplication operator $V^{(d)}$ and hence for $H_{d,t}=-tA^{(d)}+V^{(d)}$ (since $tA^{(d)}$ is bounded on $\ell^{2}(\mathbb{T})$).
-Moreover, for each $\psi\in C_{c}(\mathbb{T})$, we have $H_{d,t}^{(L)}\psi\to H_{d,t}\psi$ a.s.
-as $L\to\infty$ because $\psi$ has finite support and $V_{x}^{(L)}\to V_{x}$ pointwise.
-By the core convergence criterion for strong resolvent convergence (see, e.g., [31, Theorem VIII.25]), $(H_{d,t}^{(L)}-z)^{-1}$ converges in the strong operator topology to $(H_{d,t}-z)^{-1}$.
-Applying this fact to $\delta_{0}$ and using the uniform bound $\|(H_{d,t}^{(L)}-z)^{-1}\|\leq\eta^{-1}$ together with the dominated convergence theorem yields
-$\lim_{L\rightarrow\infty}m_{\mathbb{T}}^{(L)}(z)=m_{\mathbb{T}}(z).$ (5.8)
-Finally, we combine the previous estimates to conclude convergence in mean, then upgrade to convergence in probability. Fix $\varepsilon>0$. Choose $L$ large enough so that (5.6) and (5.8) yield
-$\sup_{n}\mathbb{E}\big[\big|m_{N}(z)-m_{N}^{(L)}(z)\big|\big]<\varepsilon/3,\qquad|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)|<\varepsilon/3.$
-Then take $N$ large enough so that (5.7) gives $|\mathbb{E}m_{N}^{(L)}(z)-m_{\mathbb{T}}^{(L)}(z)|<\varepsilon/3$. By the triangle inequality,
-$\big|\mathbb{E}\big[m_{N}(z)\big]-m_{\mathbb{T}}(z)\big|\leq\big|\mathbb{E}\big[m_{N}(z)\big]-\mathbb{E}\big[m_{N}^{(L)}(z)\big]\big|+\big|\mathbb{E}\big[m_{N}^{(L)}(z)\big]-m_{\mathbb{T}}^{(L)}(z)\big|+\big|m_{\mathbb{T}}^{(L)}(z)-m_{\mathbb{T}}(z)\big|<\varepsilon.$
-Hence
-$\lim_{N\rightarrow\infty}\mathbb{E}\big[m_{N}(z)\big]=m_{\mathbb{T}}(z)$ (5.9)
-in probability. By the $s=1$ case of Lemma 3.3, there exists $C(d)>0$ such that for all $u\geq 0$,
-$\mathbb{P}\big(\big|m_{N}(z)-\mathbb{E}m_{N}(z)\big|\geq u\big)\leq C\exp\big(-N\eta^{4}u^{2}/C\big).$ (5.10)
-Together with (5.9), this completes the proof. ∎
-5.4. Conclusion
-Proof of Lemma 3.4.
-Let
-$\widehat{\mu}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$
-denote the empirical measure of $W_{N}$.
-By a basic property of Stieltjes transforms, the convergence of Stieltjes transforms in probability at every $z\in\mathbb{H}$ provided by Proposition 5.7 implies that $\widehat{\mu}_{N}$ converges weakly to $\nu$ in probability (see, e.g., [14, Lemma 2.1]).
-Since $\nu$ is absolutely continuous, by Lemma 5.5, we have that $\widehat{\mu}_{N}(I)$ converges to $\nu(I)$ in probability for every interval $I\subset\mathbb{R}$.
-The conclusion now follows from Lemma 5.5 and Lemma 5.6.
-∎
 5. Diagonal Resolvent Convergence
 This section proves Lemma 3.4, which concerns convergence of the diagonal resolvent entries of $W_{N}$ to $R_{00}$.
-The two main technical ingredients of the proof are (i) a Combes–Thomas-type exponential decay bound for resolvents entries at fixed $\eta>0$ and (ii) the fact that a random $d$-regular graph is locally tree-like with high probability.
+The two main technical ingredients of the proof are (i) a Combes–Thomas-type exponential decay bound for resolvent entries at fixed $\eta>0$ and (ii) the fact that a random $d$-regular graph is locally tree-like with high probability.
 Using these inputs, we prove Lemma 3.4 by truncating the $d$-regular graph around a fixed vertex and then controlling the differences of the resolvent entries of the original and truncated graph by the Combes–Thomas bound.
@@
 Let $A$ denote the adjacency operator on $\ell^{2}(\mathcal{V})$, let $V$ also denote a real diagonal potential (not necessarily bounded), fix $t>0$, and define H=tA+VH=-tA+V.
-Since this notation conflicts with the shorthand introduced in in Section 2.1 for the Anderson model on the Bethe lattice, we will always use the more precise notation $H_{d,t}=-tA^{(d)}+V^{(d)}$ when discussing the Bethe lattice in this section.
+Since this notation conflicts with the shorthand introduced in Section 2.1 for the Anderson model on the Bethe lattice, we will always use the more precise notation $H_{d,t}=-tA^{(d)}+V^{(d)}$ when discussing the Bethe lattice in this section.
 For all $X\subset\mathcal{V}$, write $\chi_{X}$ for the orthogonal projection onto $\ell^{2}(X)$. For all $X,Y\subset\mathcal{V}$ define the set-distance
@@
 Lemma 5.2.
-Let $\mathcal{G}$ be a graph with maximum degree at most $d$. Fix $z=E+\mathrm{i}\eta$ with $\eta>0$ and set
+Let $\mathcal{G}$ be a graph with maximum degree at most $d$. Fix $z=E+\mathrm{i}\eta\in\mathbb{H}$ and set
 $\mu_{\mathrm{CT}}=\ln\Big(1+\frac{\eta}{4td}\Big).$
@@
 Let $K=H-H^{(r)}$.
-Then $K$ is supported on edges crossing the cut between $B_{r}(o)$ and $B_{r}(o)^{c}$, and it and equals t-t times the adjacency matrix of the graph formed by the cut edges.
+Then $K$ is supported on edges crossing the cut between $B_{r}(o)$ and $B_{r}(o)^{c}$, and it equals t-t times the adjacency matrix of the graph formed by the cut edges.
 Hence, by Lemma 5.1, $\|K\|\leq t\|A\|\leq td$.
 It follows from the definitions of $G$ and $G^{(r)}$ that
-$G-G^{(r)}=GKG^{(r)}.$
-Let $\partial B_{r}(o)$ denote the set of vertices in $\mathcal{B}$ incident to at least one cut edge (both inside and outside $B_{r}(o)$).
+$G-G^{(r)}=-GKG^{(r)}.$
+Let $\partial B_{r}(o)$ denote the set of vertices in $\mathcal{V}$ incident to at least one cut edge (both inside and outside $B_{r}(o)$).
 Then $\operatorname{dist}(\{o\},\partial B_{r}(o))=r$ and $K=\chi_{\partial B_{r}(o)}K\chi_{\partial B_{r}(o)}$.
@@
 as desired. ∎
-We then use local tree-likeness of random $d$-regular graphs to compare finite neighborhoods with the tree.
+We now use local tree-likeness of random $d$-regular graphs to compare finite neighborhoods with the tree.
 For a rooted graph $(\mathcal{G},o)$ we write $B_{r}(\mathcal{G},o)$ for the rooted radius-$r$ ball.
@@
 We briefly sketch the standard configuration-model argument.
-Let $G^{\mathrm{conf}}_{N,d}$ be the random $d$-regular multigraph from the configuration model (which places $d$ half-edges on of the $N$ vertices, then forms a perfect matching of the half-edges uniformly at random), and let $\mathbb{Q}$ be the induced measure on graph events.
+Let $G^{\mathrm{conf}}_{N,d}$ be the random $d$-regular multigraph from the configuration model (which places $d$ half-edges on each of the $N$ vertices, then forms a perfect matching of the half-edges uniformly at random), and let $\mathbb{Q}$ be the induced measure on graph events.
 We let $C(d)>0$ denote a $d$-dependent constant that may increase at each appearance in this proof.
@@
 We can construct $B_{r}(1)$ by breadth-first search in the associated configuration model, pairing one half-edge at each step.
-Let $|B_{r}(1)|$ denote the number of vertices in $B_{r}(1)$, and let $M$ be the total number of half-edges in $B_{r}(1)$ Deterministically,
+Let $|B_{r}(1)|$ denote the number of vertices in $B_{r}(1)$, and let $M$ be the total number of half-edges in $B_{r}(1)$.
+Deterministically,
 $\big|B_{r}(1)\big|\leq 1+d\sum_{k=0}^{r-1}(d-1)^{k}\leq C(d-1)^{r},$
@@
 where we increase $C$ at each step if necessary.
-During breadth-first pairing procedure, at any step, the probability that a newly exposed half-edge pairs with one of the at most $M$ previously exposed half-edges is bounded by $CM/Nd$.
+During the breadth-first pairing procedure, at any step, the probability that a newly exposed half-edge pairs with one of the at most $M$ half-edges attached to already discovered vertices is bounded by $CM/Nd$.
 A cycle in $B_{r}(1)$ can occur only if some such collision happens.
@@
 Independently sample the graphs $(\mathcal{G}_{N})_{N\geq 1}$, each uniform over simple $d$-regular graphs on $[N]$.
-Let $r_{N}=\ln\ln N$, and define the event
+Let $r_{N}=\lfloor\ln\ln N\rfloor$, and define the event
 $\mathcal{T}_{N}=\{B_{r_{N}}(\mathcal{G}_{N},1)\text{ is a tree}\}.$
@@
 $V^{(N)}_{x}=V_{\varphi_{N}^{-1}(x)},\qquad x\in B_{r_{N}}(\mathcal{G}_{N},1).$
-Outside the ball, meaning for indices $x\notin B_{r_{N}}$, we sample the $V^{(N)}_{x}$ variables as i.i.d. $N(0,1)$, which are additionally independent from all other variables in this construction.
+Outside the ball, meaning for indices $x\notin B_{r_{N}}(\mathcal{G}_{N},1)$, we sample the $V^{(N)}_{x}$ variables as i.i.d.
+$N(0,1)$, which are additionally independent of all other variables in this construction.
+On the complement of $\mathcal{T}_{N}$, we sample all $V^{(N)}_{x}$ variables as i.i.d.
+$N(0,1)$, independently from the tree potential.
 Let
@@
 Proof of Lemma 3.4.
-By Lemma 6.3 applied to $H_{N}$ (root 11) and to $H$ (root 0) with $r=r_{N}$, we obtain
+By Lemma 5.3 applied to $H_{N}$ (root 11) and to $H_{d,t}$ (root 0) with $r=r_{N}$, we obtain
 $\big|G_{N,11}(z)-G^{(r_{N})}_{N,11}(z)\big|\leq\frac{4td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r_{N}},\qquad\big|R_{00}(z)-R^{(r_{N})}_{00}(z)\big|\leq\frac{4td}{\eta^{2}}e^{-2\mu_{\mathrm{CT}}r_{N}}.$
@@
 This proves $G_{N,11}(z)\to R_{00}(z)$ in probability under the coupling. ∎
+6. Bounds for the Eigenvalue Counting Function
+This section is dedicated to the proof of Lemma 3.5.
+The main input is the convergence of the Stieltjes transform of $W_{N}$ to the Stieltjes transform of the density of states measure of $H_{d,t}$ on the Bethe lattice.
+In the Gaussian setting, this follows directly from the diagonal resolvent convergence proved in Lemma 3.4, together with the $s=1$ case of Lemma 4.1.
+Finally, since convergence of Stieltjes transforms implies convergence of the associated spectral measures, the desired lower bound on the eigenvalue counting function follows from a lower bound on the integrated density of states of $H_{d,t}$, which we show in Appendix B, and the upper bound follows from Wegner’s estimate.
+6.1. Notation
+We define
+$m_{N}(z)=\frac{1}{N}\operatorname{Tr}G_{N}(z),$
+which is the Stieltjes transform of the empirical eigenvalue distribution of $W_{N}$.
+Next, because $H_{d,t}$ is self-adjoint, the spectral theorem implies the existence of a random projection-valued measure $\mu$ such that $H_{d,t}=\int\lambda\,d\mu(\lambda)$.
+The spectral measure at the root is given by
+$\mu_{0}(S)=\langle\delta_{0},\mu(S)\delta_{0}\rangle$
+for all Borel sets $S\subset\mathbb{R}$, and the density of states measure $\nu$ is given by
+$\nu(S)=\mathbb{E}\big[\mu_{0}(S)\big].$
+For all $z\in\mathbb{H}$, the Stieltjes transform of $\nu$ is
+$m_{\mathbb{T}}(z)=\int_{\mathbb{R}}\frac{1}{\lambda-z}\,d\nu(\lambda).$
+6.2. Preliminaries
+We first state an upper bound on the integrated density of states of $H_{d,t}$, which is a consequence of a well-known estimate of Wegner [8, (5.27)].
+Lemma 6.1.
+For every $d\geq 3$ and $t>0$, there exists a constant $C(d,t)>0$ such that for all intervals $I\subset\mathbb{R}$, we have $\nu(I)\leq C|I|$.
+We also require a lower bound that complements the previous Wegner estimate. It is proved in Appendix B.
+Lemma 6.2.
+For every $d\geq 3$, $t>0$, and $D>0$, there exists a constant $c(d,t,D)>0$ such that for all intervals $I\subset[-D,D]$, we have $\nu(I)\geq c|I|$.
+We next show convergence of Stieltjes transforms.
+Proposition 6.3.
+Fix $z\in\mathbb{H}$. Then $\lim_{N\rightarrow\infty}m_{N}(z)=m_{\mathbb{T}}(z)$ in probability.
+Proof.
+Write $z=E+\mathrm{i}\eta$, where $\eta>0$ is fixed. By exchangeability of the vertices,
+$\mathbb{E}\big[m_{N}(z)\big]=\frac{1}{N}\sum_{j=1}^{N}\mathbb{E}\big[G_{jj}(z)\big]=\mathbb{E}\big[G_{11}(z)\big].$ (6.1)
+By Lemma 3.4, there exists a coupling such that $G_{11}(z)\to R_{00}(z)$ in probability.
+Since $|G_{11}(z)|\leq\eta^{-1}$ and $|R_{00}(z)|\leq\eta^{-1}$ deterministically, the bounded convergence theorem yields, after using (6.1),
+$\lim_{N\to\infty}\mathbb{E}\big[m_{N}(z)\big]=\lim_{N\to\infty}\mathbb{E}\big[G_{11}(z)\big]=\mathbb{E}\big[R_{00}(z)\big]=m_{\mathbb{T}}(z).$ (6.2)
+It remains to control the fluctuations of $m_{N}(z)$.
+By the $s=1$ case of Lemma 4.1 applied to the imaginary parts, and to the corresponding averages of the real parts, there exists $C(d)>0$ such that for all $u>0$ and all sufficiently large $N$,
+$\mathbb{P}\Big(\big|m_{N}(z)-\mathbb{E}\big[m_{N}(z)\big]\big|\geq u\Big)\leq 2C\exp\big(-N\eta^{4}u^{2}/(4C)\big),$
+which implies $m_{N}(z)-\mathbb{E}[m_{N}(z)]\to 0$ in probability. Combining this with (6.2) completes the proof. ∎
+6.3. Conclusion
+Proof of Lemma 3.5.
+Let
+$\widehat{\mu}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$
+denote the empirical measure of $W_{N}$.
+By a basic property of Stieltjes transforms, the convergence of Stieltjes transforms in probability at every $z\in\mathbb{H}$ provided by Proposition 6.3 implies that $\widehat{\mu}_{N}$ converges weakly to $\nu$ in probability (see, e.g., [15, Lemma 2.1]).
+Since $\nu$ is absolutely continuous, by Lemma 6.1, we have that $\widehat{\mu}_{N}(I)$ converges to $\nu(I)$ in probability for every interval $I\subset\mathbb{R}$.
+The conclusion now follows from Lemma 6.1 and Lemma 6.2.
+∎
 7. Uniform Moment Bound for the Resolvent
@@
 7.1. Preliminaries
-It is well known (see, e.g., [5, (2.2)]) that for every $z\in\mathbb{H}$, the resolvent entry $R^{\prime}_{00}(z)$ for the Anderson model on a variant of $\mathbb{T}_{d}$, where the root degree is $d$ and the degree of all other vertices is $d+1$, solves the recursive distribution equation
-$\Gamma_{\eta}\stackrel{{\scriptstyle d}}{{=}}\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d}\Gamma_{\eta,j}},$ (7.1)
-where $V_{0}\sim N(0,1)$, the variables $\Gamma_{\eta,1},\dots,\Gamma_{\eta,K}$ and $\Gamma_{\eta}$ are i.i.d. and all variables on the right-hand side are independent.
+It is well known (see, e.g., [5, (2.2)]) that for every $z\in\mathbb{H}$, the resolvent entry $R^{\prime}_{00}(z)$ for the Anderson model on a variant of $\mathbb{T}_{d}$, where the root degree is d1d-1 and the degree of all other vertices is $d$, solves the recursive distribution equation
+$\Gamma_{\eta}\stackrel{{\scriptstyle d}}{{=}}\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d-1}\Gamma_{\eta,j}},$ (7.1)
+where $V_{0}\sim N(0,1)$, the variables $\Gamma_{\eta,1},\dots,\Gamma_{\eta,d-1}$ and $\Gamma_{\eta}$ are i.i.d. and all variables on the right-hand side are independent.
 We write
 $\Gamma_{\eta}=X_{\eta}+\mathrm{i}Y_{\eta},$ (7.2)
-with $X_{\eta}$ and $Y_{\eta}$ are real. We have $Y_{\eta}>0$ almost surely and
+where $X_{\eta}$ and $Y_{\eta}$ are real. We have $Y_{\eta}>0$ almost surely and
 $X_{\eta}=\frac{U_{\eta}}{U_{\eta}^{2}+S_{\eta}^{2}},\qquad Y_{\eta}=\frac{S_{\eta}}{U_{\eta}^{2}+S_{\eta}^{2}},$ (7.3)
 where
-$U_{\eta}=V-E-t^{2}\sum_{j=1}^{d}X_{\eta,j},\qquad S_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta,j},$ (7.4)
+$U_{\eta}=V_{0}-E-t^{2}\sum_{j=1}^{d-1}X_{\eta,j},\qquad S_{\eta}=\eta+t^{2}\sum_{j=1}^{d-1}Y_{\eta,j},$ (7.4)
 with the $X_{\eta,j}$ and $Y_{\eta,j}$ defined as in (7.2). Moreover, the Schur complement formula (see, e.g., [8]) implies
-$R_{00}(E+\mathrm{i}\eta)=\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d+1}\Gamma_{\eta}^{(j)}},$ (7.5)
-where $V_{0}\sim N(0,1)$ is independent of the i.i.d. random variables $\Gamma_{\eta}^{(1)},\dots,\Gamma_{\eta}^{(d+1)}$ of $\Gamma_{\eta}$. In particular, defining
-$B_{\eta}=\eta+t^{2}\sum_{j=1}^{d+1}Y_{\eta}^{(j)},$ (7.6)
+$R_{00}(E+\mathrm{i}\eta)=\frac{1}{V_{0}-E-\mathrm{i}\eta-t^{2}\sum_{j=1}^{d}\Gamma_{\eta}^{(j)}},$ (7.5)
+where $V_{0}\sim N(0,1)$ is independent of the i.i.d. random variables $\Gamma_{\eta}^{(1)},\dots,\Gamma_{\eta}^{(d)}$, each distributed as $\Gamma_{\eta}$. In particular, defining
+$B_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta}^{(j)},$ (7.6)
 where $Y_{\eta}^{(j)}$ is defined analogously to $Y_{\eta,j}$, there is a real random variable $A_{\eta}$, independent of $V_{0}$, such that
 $\operatorname{Im}R_{00}(E+\mathrm{i}\eta)=\frac{B_{\eta}}{(V_{0}-A_{\eta})^{2}+B_{\eta}^{2}}.$ (7.7)
-For all lemmas in this section, we retain the notation and hypothesis of Lemma 3.6. In particular, $E\in\mathbb{R}$ is always fixed.
+For all lemmas in this section, we retain the notation and hypothesis of Lemma 3.6.
+In particular, $E\in\mathbb{R}$, $d\geq 10$, and $t>0$ are always fixed.
+We also let $\rho$ denote the density of a mean zero, variance one Gaussian random variable.
+The constants in this section may depend on the constant in the hypothesis of Lemma 3.6, but we always omit this dependence in the notation.
 Lemma 7.1.
-There exist constants $\eta_{0}(E)\in(0,1]$, $y_{*}(E)>0$, and $p_{*}(E)>0$, such that for all $\eta\in(0,\eta_{0}]$,
+There exist constants $\eta_{0}(d,t,E)\in(0,1]$, $y_{*}(d,t,E)>0$, and $p_{*}(d,t,E)>0$, such that for all $\eta\in(0,\eta_{0}]$,
 $\mathbb{P}(Y_{\eta}>y_{*})\geq p_{*}.$ (7.8)
@@
 Choose $p_{0}>0$ and $\eta_{1}\in(0,1]$ so that, for all $0<\eta\leq\eta_{1}$,
-$\mathbb{P}\bigl(\operatorname{Im}R(E+\mathrm{i}\eta)>c\bigr)\geq 2p_{0}.$ (7.9)
+$\mathbb{P}\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\bigr)\geq 2p_{0}.$ (7.9)
 Fix $0<\eta\leq\eta_{1}$. Conditional on $(A_{\eta},B_{\eta})$ in (7.7), the event $\{\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c\}$ is empty unless $B_{\eta}<1/c$, in which case it is equivalent to
@@
 Now set
-$\eta_{0}=\min\{\eta_{1},b_{*}/2\},\qquad y_{*}=\frac{b_{*}}{2(d+1)t^{2}}.$
+$\eta_{0}=\min\{\eta_{1},b_{*}/2\},\qquad y_{*}=\frac{b_{*}}{2dt^{2}}.$
 If $0<\eta\leq\eta_{0}$ and $B_{\eta}>b_{*}$, then
-$t^{2}\sum_{j=1}^{d+1}Y_{\eta}^{(j)}\geq B_{\eta}-\eta>b_{*}/2,$
+$t^{2}\sum_{j=1}^{d}Y_{\eta}^{(j)}\geq B_{\eta}-\eta>b_{*}/2,$
 so for at least one index $j$ we have $Y_{\eta}^{(j)}>y_{*}$. By the union bound and symmetry,
-$p_{1}\leq\mathbb{P}\left(\bigcup_{j=1}^{d+1}\{Y_{\eta}^{(j)}>y_{*}\}\right)\leq(d+1)\mathbb{P}(Y_{\eta}>y_{*}).$
+$p_{1}\leq\mathbb{P}\left(\bigcup_{j=1}^{d}\{Y_{\eta}^{(j)}>y_{*}\}\right)\leq d\mathbb{P}(Y_{\eta}>y_{*}).$
 Thus (7.8) holds with
-$p_{*}=\frac{p_{1}}{d+1}>0.\qed$
+$p_{*}=\frac{p_{1}}{d}>0.\qed$
 Lemma 7.2.
-There exists a constant $C(t,E,d)>0$, such that for every $0<\eta\leq\eta_{0}$, the following bounds hold.
+There exists a constant $C(d,t,E)>0$, such that for every $0<\eta\leq\eta_{0}$, the following bounds hold.
   1. (1)
@@
 Proof.
-Fix $a\in\mathbb{R}$, $s>0$, and $u>0$. The set
+For all $s>0$ and $u>0$, the set
 $\left\{x\in\mathbb{R}:\frac{|x|}{x^{2}+s^{2}}>u\right\}$
-has Lebesgue measure at most $2/u$. Therefore, conditioning on
-$A=E+t^{2}\sum_{j=1}^{d}X_{\eta,j}\qquad\text{and}\qquad S_{\eta}=s,$
+has Lebesgue measure at most $2/u$. Therefore, conditioning on $S_{\eta}$ and
+$A=E+t^{2}\sum_{j=1}^{d-1}X_{\eta,j},$
 we obtain from (7.3) that
@@
 For (7.14), we use (7.4) and the union bound to get
-$\mathbb{P}\big(|U_{\eta}|>u\big)\leq\mathbb{P}\big(|V-E|>u/2\big)+\sum_{j=1}^{d}\mathbb{P}\left(|X_{\eta,j}|>\frac{u}{2dt^{2}}\right).$
+$\mathbb{P}\big(|U_{\eta}|>u\big)\leq\mathbb{P}\big(|V_{0}-E|>u/2\big)+\sum_{j=1}^{d-1}\mathbb{P}\left(|X_{\eta,j}|>\frac{u}{2(d-1)t^{2}}\right).$
 The first term is bounded by $C/u$ for $u\geq 1$ and the second by (7.13), which gives (7.14).
 Finally, if $u\geq 2\eta_{0}$ and $0<\eta\leq\eta_{0}$, then
-$S_{\eta}=\eta+t^{2}\sum_{j=1}^{d}Y_{\eta,j}>u$
+$S_{\eta}=\eta+t^{2}\sum_{j=1}^{d-1}Y_{\eta,j}>u$
 implies
-$\sum_{j=1}^{d}Y_{\eta,j}>\frac{u}{2t^{2}},$
-so at least one index satisfies $Y_{\eta,j}>u/(2dt^{2})$. Hence
-$\mathbb{P}(S_{\eta}>u)\leq\sum_{j=1}^{d}\mathbb{P}\left(Y_{\eta,j}>\frac{u}{2dt^{2}}\right),$
+$\sum_{j=1}^{d-1}Y_{\eta,j}>\frac{u}{2t^{2}},$
+so at least one index satisfies $Y_{\eta,j}>u/(2(d-1)t^{2})$. Hence
+$\mathbb{P}(S_{\eta}>u)\leq\sum_{j=1}^{d-1}\mathbb{P}\left(Y_{\eta,j}>\frac{u}{2(d-1)t^{2}}\right),$
 and (7.15) follows from (7.13). ∎
@@
 $F_{\eta}(\varepsilon)=\mathbb{P}(Y_{\eta}\leq\varepsilon).$
-There exist constants $C_{1}(t,E,d),C_{2}(t,E,d)>0$ and $\varepsilon_{1}(t,E,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$ and $0<\eta\leq\eta_{0}$,
-$F_{\eta}(\varepsilon)\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d}+C_{2}\varepsilon^{1/3}.$ (7.16)
+There exist constants $C_{1}(d,t,E),C_{2}(d,t,E)>0$ and $\varepsilon_{1}(d,t,E)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{1}$ and $0<\eta\leq\eta_{0}$,
+$F_{\eta}(\varepsilon)\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d-1}+C_{2}\varepsilon^{1/3}.$ (7.16)
 Proof.
@@
 so by independence,
-$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(t^{-2}\varepsilon^{1/3})^{d}.$ (7.17)
+$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(t^{-2}\varepsilon^{1/3})^{d-1}.$ (7.17)
 On the third event, (7.15) yields
@@
 $\mathbb{P}\bigl(\varepsilon^{1/3}<S_{\eta}\leq\varepsilon^{-1/3},Y_{\eta}\leq\varepsilon\bigr)\leq\mathbb{P}\bigl(|U_{\eta}|\geq 2^{-1/2}\varepsilon^{-1/3}\bigr)\leq C\varepsilon^{1/3}$
-by (7.14), for all $\varepsilon\leq\varepsilon_{1}$ small enough that $2^{-1/2}\varepsilon^{-1/3}\geq 1$. Combining this with (7.17) and (7.18) proves (7.16) with $C_{1}=t^{-2}$. ∎
+by (7.14), for all $\varepsilon\leq\varepsilon_{1}$ small enough that $2^{-1/2}\varepsilon^{-1/3}\geq 1$.
+Set $C_{1}=\max\{1,t^{-2}\}$.
+Since $F_{\eta}$ is nondecreasing and $t^{-2}\leq C_{1}$, (7.17) implies
+$\mathbb{P}(S_{\eta}\leq\varepsilon^{1/3})\leq F_{\eta}(C_{1}\varepsilon^{1/3})^{d-1}.$
+Combining this with (7.18) and the bound on the middle event proves (7.16). ∎
 Lemma 7.4.
-There exist constants $C_{*}(t,E,d)>0$, and $\varepsilon_{*}(E,t,d)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{*}$ and $0<\eta\leq\eta_{0}$,
-$\mathbb{P}(Y_{\eta}\leq\varepsilon)\leq C_{*}\varepsilon^{\beta},\qquad\beta=\frac{1}{20}.$ (7.19)
+There exist constants $C_{*}(d,t,E)>0$ and $\varepsilon_{*}(d,t,E)\in(0,1]$ such that, for all $0<\varepsilon\leq\varepsilon_{*}$ and $0<\eta\leq\eta_{0}$,
+$\mathbb{P}(Y_{\eta}\leq\varepsilon)\leq C_{*}\varepsilon^{\beta},\qquad\beta=\frac{31}{301}.$ (7.19)
 Proof.
-Let $C_{1},C_{2},\varepsilon_{1}$ be as in Lemma 7.3. Choose
+Let $C_{1},C_{2},\varepsilon_{1}$ be as in Lemma 7.3, and set
+$\bar{q}=1-p_{*}\in(0,1).$
+Define a deterministic sequence for $k\geq 0$ by
+$r_{0}=\bar{q},\qquad r_{k+1}=\frac{1}{2}(r_{k}^{\,d-1}+r_{k}).$
+Since $0<r_{k}<1$, we have $r_{k+1}<r_{k}$ for all $k$, and hence $r_{k}\rightarrow 0$ (since any limit point must satisfy the previous recursion). Choose $m\in\mathbb{N}$ such that
+$r_{m}\leq\frac{1}{2}.$
+Choose
 $\tilde{\varepsilon}_{0}\in(0,\min\{y_{*},\varepsilon_{1},1\}]$
-so small that the sequence $(\tilde{\varepsilon}_{n})_{n\geq 0}$, defined by
-$\tilde{\varepsilon}_{n+1}=\left(\frac{\tilde{\varepsilon}_{n}}{2C_{1}}\right)^{3},$ (7.20)
-satisfies, for every $n\geq 0$,
-$L_{n}=-\ln\tilde{\varepsilon}_{n}\geq\max\{12\ln C_{2},20\ln 2\},\qquad L_{n+1}\leq 4L_{n}.$ (7.21)
-This is possible because $L_{n+1}=3L_{n}+3\ln(2C_{1})$.
+so small that, for the sequence defined by
+$\tilde{\varepsilon}_{n+1}=\Bigl(\frac{\tilde{\varepsilon}_{n}}{C_{1}}\Bigr)^{3},\qquad L_{n}=-\ln\tilde{\varepsilon}_{n},$
+the following hold:
+$L_{0}\geq 300\ln C_{1},\qquad L_{1}\geq\max\{50\ln(2C_{2}),\,2\},$ (7.20)
+and
+$C_{2}\tilde{\varepsilon}_{k+1}^{1/3}\leq\frac{1}{2}\bigl(r_{k}-r_{k}^{\,d-1}\bigr)\qquad\text{for all }0\leq k<m.$ (7.21)
+Since $L_{n+1}=3L_{n}+3\ln C_{1}$, the first condition implies
+$L_{n+1}\leq\frac{301}{100}L_{n}\qquad\text{for all }n\geq 0.$ (7.22)
 Define
 $q_{n}=\sup_{0<\eta\leq\eta_{0}}F_{\eta}(\tilde{\varepsilon}_{n}),\qquad\nu_{n}=-\ln q_{n}.$
-Since $\tilde{\varepsilon}_{0}\leq y_{*}$ and Lemma 7.1 gives $\mathbb{P}(Y_{\eta}>y_{*})\geq p_{*}$, we have
-$q_{0}\leq 1-p_{*}<1,\qquad\nu_{0}>0.$ (7.22)
-Moreover, because $C_{1}\tilde{\varepsilon}_{n+1}^{1/3}=\tilde{\varepsilon}_{n}/2<\tilde{\varepsilon}_{n}<\varepsilon_{1}$, Lemma 7.3 implies
-$q_{n+1}\leq q_{n}^{d}+C_{2}\tilde{\varepsilon}_{n+1}^{1/3}.$ (7.23)
-By (7.21),
-$C_{2}\tilde{\varepsilon}_{n+1}^{1/3}=\exp\left(-\frac{1}{3}L_{n+1}+\ln C_{2}\right)\leq\exp\left(-\frac{1}{4}L_{n+1}\right).$
-Therefore,
-$\nu_{n+1}\geq\min\left\{d\nu_{n},\frac{1}{4}L_{n+1}\right\}-\ln 2.$ (7.24)
-We claim that there exists an index $n_{*}$ such that
-$d\nu_{n_{*}}\geq\frac{1}{4}L_{n_{*}+1}.$ (7.25)
-If not, then (7.24) would imply, for all $n$, that
-$\nu_{n+1}\geq d\nu_{n}-\ln 2.$
-Setting
-$w_{n}=\nu_{n}+\frac{\ln 2}{d-1},$
-we would have $w_{n+1}\geq dw_{n}$, hence
-$w_{n}\geq d^{n}w_{0}.$
-On the other hand, the negation of (7.25) gives
-$\nu_{n}<\frac{L_{n+1}}{4d}\leq\frac{4^{n}L_{1}}{4d}$
-by (7.21), so $w_{n}\leq C4^{n}$. Since $d\geq 20>4$, this is impossible for large $n$. Thus there exists some $n_{*}\in\mathbb{N}$ such that (7.25) holds.
-Now (7.24), (7.25), and (7.21) give
-$\nu_{n_{*}+1}\geq\frac{1}{4}L_{n_{*}+1}-\ln 2\geq\frac{1}{5}L_{n_{*}+1}.$
-Because $L_{n_{*}+2}\leq 4L_{n_{*}+1}$, we obtain
-$\nu_{n_{*}+1}\geq\frac{1}{20}L_{n_{*}+2}.$
-We now prove by induction that, for all $n\geq n_{*}+1$,
-$\nu_{n}\geq\beta L_{n+1},\qquad\beta=\frac{1}{20}.$ (7.26)
-The base case has just been established. If (7.26) holds at index $n$, then
-$d\nu_{n}\geq d\beta L_{n+1}=\frac{d}{20}L_{n+1}\geq L_{n+1}\geq\frac{1}{4}L_{n+2},$
-where the last inequality again uses (7.21). Hence (7.24) and $L_{n+2}\geq 20\ln 2$ yield
-$\nu_{n+1}\geq\frac{1}{4}L_{n+2}-\ln 2\geq\beta L_{n+2}.$
-This proves (7.26).
-Consequently, for all $n\geq n_{*}+1$,
-$q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}.$
-Now let $0<\varepsilon\leq\varepsilon_{*}=\tilde{\varepsilon}_{n_{*}+2}$ and choose $n\geq n_{*}+1$ so that
+Because $\tilde{\varepsilon}_{0}\leq y_{*}$, Lemma 7.1 gives
+$q_{0}\leq\bar{q}=r_{0}.$
+Also, since $C_{1}\tilde{\varepsilon}_{n+1}^{1/3}=\tilde{\varepsilon}_{n}$, Lemma 7.3 yields
+$q_{n+1}\leq q_{n}^{\,d-1}+C_{2}\tilde{\varepsilon}_{n+1}^{1/3}.$ (7.23)
+We claim that
+$q_{k}\leq r_{k}\qquad\text{for all }0\leq k\leq m.$
+This is true for $k=0$. If $k<m$ and $q_{k}\leq r_{k}$, then by (7.21),
+$q_{k+1}\leq q_{k}^{\,d-1}+C_{2}\tilde{\varepsilon}_{k+1}^{1/3}\leq r_{k}^{\,d-1}+\frac{1}{2}\bigl(r_{k}-r_{k}^{\,d-1}\bigr)=\frac{1}{2}\bigl(r_{k}^{\,d-1}+r_{k}\bigr)=r_{k+1}.$
+Thus the claim follows by induction. In particular,
+$q_{m}\leq r_{m}\leq\frac{1}{2},$
+and hence
+$\nu_{m}\geq\ln 2.$ (7.24)
+From (7.23) and $a+b\leq 2\max\{a,b\}$, and the definitions of $L_{n}$ and $\nu_{n}$, we get
+$\nu_{n+1}\geq\min\Bigl\{(d-1)\nu_{n},\frac{1}{3}L_{n+1}-\ln C_{2}\Bigr\}-\ln 2.$
+Since $L_{n+1}\geq L_{1}\geq 50\ln(2C_{2})$, we obtain
+$\frac{1}{3}L_{n+1}-\ln C_{2}-\ln 2=\frac{1}{3}L_{n+1}-\ln(2C_{2})\geq\frac{31}{100}L_{n+1}.$
+Therefore
+$\nu_{n+1}\geq\min\Bigl\{(d-1)\nu_{n}-\ln 2,\frac{31}{100}L_{n+1}\Bigr\}\qquad\text{for all }n\geq 0.$ (7.25)
+We claim that there exists $n_{*}\geq m$ such that
+$\nu_{n_{*}}\geq\frac{31}{301}L_{n_{*}+1}.$ (7.26)
+Suppose not. Then
+$\nu_{n}<\frac{31}{301}L_{n+1}\qquad\text{for every }n\geq m.$
+If for some $n\geq m$ we had
+$(d-1)\nu_{n}-\ln 2\geq\frac{31}{100}L_{n+1},$
+then (7.25) would imply
+$\nu_{n+1}\geq\frac{31}{100}L_{n+1}\geq\frac{31}{301}L_{n+2},$
+by (7.22), contradicting the assumption. Hence necessarily
+$(d-1)\nu_{n}-\ln 2<\frac{31}{100}L_{n+1}\qquad\text{for all }n\geq m,$
+and then (7.25) gives
+$\nu_{n+1}\geq(d-1)\nu_{n}-\ln 2\qquad\text{for all }n\geq m.$
+Iterating,
+$\nu_{n}\geq(d-1)^{n-m}\Bigl(\nu_{m}-\frac{\ln 2}{d-2}\Bigr)+\frac{\ln 2}{d-2}\qquad\text{for all }n\geq m.$
+By (7.24) and the assumption $d\geq 10$, the coefficient
+$\nu_{m}-\frac{\ln 2}{d-2}$
+is strictly positive. On the other hand, (7.22) implies
+$L_{n+1}\leq\Bigl(\frac{301}{100}\Bigr)^{n-m}L_{m+1}.$
+Since $d-1\geq 9>301/100$, the inequality
+$(d-1)^{n-m}\Bigl(\nu_{m}-\frac{\ln 2}{d-2}\Bigr)+\frac{\ln 2}{d-2}<\frac{31}{301}\Bigl(\frac{301}{100}\Bigr)^{n-m}L_{m+1}$
+cannot hold for all large $n$, contradiction. This proves (7.26).
+Now (7.26) propagates. Indeed, if for some $n\geq n_{*}$,
+$\nu_{n}\geq\frac{31}{301}L_{n+1},$
+then, because $d-1\geq 9$ and $L_{n+1}\geq L_{1}\geq 2$,
+$(d-1)\nu_{n}-\ln 2\geq\frac{279}{301}L_{n+1}-\ln 2\geq\frac{31}{100}L_{n+1}.$
+Thus (7.25) implies
+$\nu_{n+1}\geq\frac{31}{100}L_{n+1}\geq\frac{31}{301}L_{n+2},$
+again by (7.22). By induction,
+$\nu_{n}\geq\frac{31}{301}L_{n+1}\qquad\text{for all }n\geq n_{*}.$
+Equivalently,
+$q_{n}\leq\tilde{\varepsilon}_{n+1}^{31/301}\qquad\text{for all }n\geq n_{*}.$
+Finally, let $\varepsilon_{*}=\tilde{\varepsilon}_{n_{*}}$, and fix $0<\varepsilon\leq\varepsilon_{*}$. Choose $n\geq n_{*}$ such that
 $\tilde{\varepsilon}_{n+1}<\varepsilon\leq\tilde{\varepsilon}_{n}.$
 Since $F_{\eta}$ is nondecreasing,
-$F_{\eta}(\varepsilon)\leq q_{n}\leq\tilde{\varepsilon}_{n+1}^{\beta}<\varepsilon^{\beta}.$
-Thus (7.19) holds with $C_{*}=1$ and $\beta=1/20$. ∎
+$F_{\eta}(\varepsilon)\leq q_{n}\leq\tilde{\varepsilon}_{n+1}^{31/301}<\varepsilon^{31/301}.$
+This proves (7.19) with $C_{*}=1$. ∎
 Lemma 7.5.
-There exists a constant $C(E,t,d)>0$ such that
+There exists a constant $C(t,d,E)>0$ such that
 $\sup_{0<\eta\leq\eta_{0}}\mathbb{E}[B_{\eta}^{-1}]\leq C.$ (7.27)
 Proof.
-Let $\beta=1/20$ and $\varepsilon_{*}>0$ be given by Lemma 7.4. If $0<\varepsilon\leq t^{2}\varepsilon_{*}$ and $B_{\eta}\leq\varepsilon$, then each summand in (7.6) with $1\leq j\leq d+1$ satisfies
+Let $\beta=31/301$ and $\varepsilon_{*}>0$ be given by Lemma 7.4. If $0<\varepsilon\leq t^{2}\varepsilon_{*}$ and $B_{\eta}\leq\varepsilon$, then each summand in (7.6) with $1\leq j\leq d$ satisfies
 $Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}},$
 since all summands are nonnegative. Therefore, by independence and Lemma 7.4,
-$\mathbb{P}(B_{\eta}\leq\varepsilon)\leq\prod_{j=1}^{d+1}\mathbb{P}\left(Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}}\right)\leq C\varepsilon^{\beta(d+1)}.$
-Because $\beta(d+1)=(d+1)/20>1$, we conclude that
-$\sup_{0<\eta\leq\eta_{0}}\int_{1}^{\infty}\mathbb{P}(B_{\eta}<s^{-1})\,ds<\infty,$
-which is equivalent to (7.27). ∎
+$\mathbb{P}(B_{\eta}\leq\varepsilon)\leq\prod_{j=1}^{d}\mathbb{P}\left(Y_{\eta}^{(j)}\leq\frac{\varepsilon}{t^{2}}\right)\leq C\varepsilon^{\beta d}.$
+Set $s_{0}=\max\{1,(t^{2}\varepsilon_{*})^{-1}\}$. For all $s\geq s_{0}$, we have $s^{-1}\leq t^{2}\varepsilon_{*}$, hence
+$\mathbb{P}(B_{\eta}<s^{-1})\leq Cs^{-\beta d}.$
+Therefore,
+$\mathbb{E}[B_{\eta}^{-1}]\leq 1+\int_{1}^{\infty}\mathbb{P}(B_{\eta}<s^{-1})\,ds\leq 1+\int_{1}^{s_{0}}1\,ds+C\int_{s_{0}}^{\infty}s^{-\beta d}\,ds.$
+Since $\beta d>1$, the right-hand side is finite uniformly in $0<\eta\leq\eta_{0}$. This proves (7.27). ∎
 We now return to Lemma 3.6.
-Proof.
+Proof of Lemma 3.6.
 Let $\eta\in(0,\eta_{0}]$. Conditioning on $(A_{\eta},B_{\eta})$ in (7.7),
@@
 $\sup_{\eta\in[\eta_{0},1]}\mathbb{E}\left[\bigl(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)\bigr)^{2}\right]\leq\eta_{0}^{-2}<\infty.$
-Together with (7.28), this proves (3.6). ∎
+Together with (7.28), this proves Lemma 3.6. ∎
 Appendix A Combes–Thomas Bound
@@
 We first prove the desired Combes–Thomas estimate on finite volumes, then pass to the infinite graph by strong resolvent convergence and a finite-rank sandwiching argument.
-Throughout this appendix, we let $\mathcal{G}=(V,E)$ be a graph (finite or finite) with maximum degree at most $d$, and we retain the notation introduced in Section 6.
+Throughout this appendix, we let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph (finite or infinite) with maximum degree at most $d$, and we retain the notation introduced in Section 5.
 Additionally, for vertices $v,w$ in a graph, we write $v\sim w$ if there is an edge between $v$ and $w$.
 A.1. Finite-Volume Estimate
-Fix a finite set $X\subset\mathcal{V}$. For all $r\in\mathbb{N}$ define the (finite) neighborhood
+Fix any finite set $X\subset\mathcal{V}$ and $r\in\mathbb{N}$, and define the (finite) neighborhood
 $\Lambda_{r}=\{x\in\mathcal{V}:\operatorname{dist}(x,X)\leq r\}.$
@@
 $\widehat{H}_{r}=H_{\Lambda_{r}}\oplus H_{\Lambda_{r}^{c}},$
-on $\ell^{2}(\mathcal{V})$ by by cutting all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ in the underlying graph and removing the corresponding entries from $H$ (so that $\widehat{H}_{r}$ is self-adjoint).
+on $\ell^{2}(\mathcal{V})$ by cutting all edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ in the underlying graph and removing the corresponding entries from $H$ (so that $\widehat{H}_{r}$ is self-adjoint).
 If $X,Y\subset\Lambda_{r}$, then
@@
 Since $A$ is bounded and $V$ is a self-adjoint multiplication operator, $C_{c}(\mathcal{V})$ is a core for $H$ and every $\widehat{H}_{r}$.
-For any $\psi\in C_{c}(V)$, we have $\operatorname{supp}\psi\subset\Lambda_{r}$ for all sufficiently large $r$, in which case the cut edges between $\Lambda_{r}$ and $\Lambda_{r}^{c}$ do not affect the value of $H\psi$ because $\psi$ vanishes on $\Lambda_{r}^{c}$.
-Hence for all sufficiently large $r$,
+For any $\psi\in C_{c}(\mathcal{V})$, let $\mathcal{C}_{X}\subset\mathcal{V}$ be the union of the connected components of $\mathcal{G}$ that intersect $X$.
+Since $\operatorname{supp}\psi\cap\mathcal{C}_{X}$ is finite and every vertex in $\mathcal{C}_{X}$ has finite distance to $X$, there exists $R$ such that
+$\operatorname{supp}\psi\cap\mathcal{C}_{X}\subset\Lambda_{R-1}.$
+Hence for all $r\geq R$,
+$\operatorname{supp}\psi\subset\Lambda_{r-1}\cup\mathcal{C}_{X}^{c}.$
+Now every edge removed in passing from $H$ to $\widehat{H}_{r}$ has one endpoint in $\Lambda_{r}$ and the other in $\Lambda_{r}^{c}$, so both endpoints lie in $\mathcal{C}_{X}$; moreover, the endpoint in $\Lambda_{r}$ actually lies in $\Lambda_{r}\setminus\Lambda_{r-1}$.
+Therefore no removed edge is incident to a vertex where $\psi$ is nonzero: on $\mathcal{C}_{X}$ the support of $\psi$ lies inside $\Lambda_{r-1}$, while on $\mathcal{C}_{X}^{c}$ the operators $\widehat{H}_{r}$ and $H$ already agree.
+Hence for all $r\geq R$,
 $\widehat{H}_{r}\psi=H\psi,$
-and therefore $\widehat{H}_{r}\psi\to H\psi$ as $r\rightarrow\infty$ for every $\psi\in C_{c}(V)$.
-By the core criterion for strong resolvent convergence (e.g., [31, Theorem VIII.25]), this implies $\widehat{H}_{r}\to H$ in the strong resolvent sense, which is exactly the desired conclusion.
+and therefore $\widehat{H}_{r}\psi\to H\psi$ as $r\rightarrow\infty$ for every $\psi\in C_{c}(\mathcal{V})$.
+By the core criterion for strong resolvent convergence (e.g., [32, Theorem VIII.25]), this implies $\widehat{H}_{r}\to H$ in the strong resolvent sense, which is exactly the desired conclusion.
 ∎
 Lemma A.3.
-Let $(T_{r})_{r=1}^{\infty}$ and $T$ be bounded operators on a Hilbert space such that $\sup_{r}\|T_{r}\|<\infty$, and suppose $\lim_{r\rightarrow\infty}T_{r}\to T$ in the strong operator topology.
+Let $(T_{r})_{r=1}^{\infty}$ and $T$ be bounded operators on a Hilbert space such that $\sup_{r}\|T_{r}\|<\infty$, and suppose $T_{r}\to T$ in the strong operator topology as $r\rightarrow\infty$.
 If $P$ and $Q$ are finite-rank orthogonal projections, then
@@
 Proof of Lemma 5.2.
-We first prove the claim when both $X$ and $Y$ are finite. Fix finite $X,Y\subset V$ and choose $r$ large enough so that $X\cup Y\subset\Lambda_{r}$. By Lemma A.1 and the identity (A.1),
-$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}\big\|=\big\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y}\big\|\leq\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$ (A.2)
+We abbreviate $\mu=\mu_{\mathrm{CT}}$.
+First, we prove the claim when both $X$ and $Y$ are finite.
+Fix finite $X,Y\subset\mathcal{V}$ and choose $r$ large enough so that $X\cup Y_{\mathrm{conn}}\subset\Lambda_{r}$, where $Y_{\mathrm{conn}}$ is the intersection of $Y$ with the union of the connected components of $\mathcal{G}$ meeting $X$.
+By Lemma A.1 and the identity (A.1),
+$\big\|\chi_{X}(\widehat{H}_{r}-z)^{-1}\chi_{Y}\big\|=\big\|\chi_{X}(H_{\Lambda_{r}}-z)^{-1}\chi_{Y_{\mathrm{conn}}}\big\|\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X,Y)}.$ (A.2)
 By Lemma A.2, $(\widehat{H}_{r}-z)^{-1}\to(H-z)^{-1}$ in the strong operator topology.
@@
 Therefore, by the finite-set Combes–Thomas estimate,
-$\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\|\leq C\,e^{-\mu\,\operatorname{dist}(X_{n},Y_{n})}\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\qquad\text{for all }n.$ (A.3)
+$\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\|\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X_{n},Y_{n})}\leq\frac{2}{\eta}\,e^{-\mu\operatorname{dist}(X,Y)}\qquad\text{for all }n.$ (A.3)
 Next, for any $\psi\in\ell^{2}(\mathcal{V})$,
@@
 Since $(H-z)^{-1}$ is bounded and $\chi_{X_{n}}\to\chi_{X}$, $\chi_{Y_{n}}\to\chi_{Y}$ strongly, both terms on the right-hand side converge to zero in norm.
-In combination with , this shows that for any $\psi\in\ell^{2}(\mathcal{V})$,
-$\|\chi_{X}R(z)\chi_{Y}\psi\|=\lim_{n\to\infty}\|\chi_{X_{n}}R(z)\chi_{Y_{n}}\psi\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}\|\psi\|.$
+In combination with (A.3), this shows that for any $\psi\in\ell^{2}(\mathcal{V})$,
+$\|\chi_{X}(H-z)^{-1}\chi_{Y}\psi\|=\lim_{n\to\infty}\|\chi_{X_{n}}(H-z)^{-1}\chi_{Y_{n}}\psi\|\leq\frac{2}{\eta}e^{-\mu\operatorname{dist}(X,Y)}\|\psi\|.$
 Taking the supremum over all $\psi$ with $\|\psi\|=1$ yields
-$\|\chi_{X}R(z)\chi_{Y}\|\leq C\,e^{-\mu\,\operatorname{dist}(X,Y)}=\frac{2}{\eta}\,e^{-\mu\,\operatorname{dist}(X,Y)}.$
+$\|\chi_{X}(H-z)^{-1}\chi_{Y}\|\leq\frac{2}{\eta}e^{-\mu\operatorname{dist}(X,Y)}.$
 This proves the claim. ∎
 Appendix B Lower Bound on the Density of States
-This appendix is devoted to the proof of Lemma 5.6. We begin with an elementary fact about self-adjoint operators.
+This appendix is devoted to the proof of Lemma 6.2.
+For each neighbor $y\sim 0$, let $\mathbb{T}_{y}$ denote the connected component of $\mathbb{T}\setminus\{0\}$ containing $y$, and let $H_{\mathbb{T}_{y}}$ be the restriction of $H_{d,t}$ to $\ell^{2}(\mathbb{T}_{y})$.
+For $z=E+\mathrm{i}\eta$ with $\eta>0$, define
+$\Gamma_{y}(z)=\langle\delta_{y},(H_{\mathbb{T}_{y}}-z)^{-1}\delta_{y}\rangle.$
+By the Schur complement formula,
+$R_{00}(z)=\langle\delta_{0},(H_{d,t}-z)^{-1}\delta_{0}\rangle=\frac{1}{V_{0}-z-t^{2}\sum_{y\sim 0}\Gamma_{y}(z)}.$
 Lemma B.1.
-Let $H$ be a self-adjoint operator on a Hilbert space $\mathcal{H}$, $\varphi\in\mathcal{H}$ such that $\varphi\neq 0$, $E\in\mathbb{R}$, and $\varepsilon>0$.
-For all Borel sets $J\subset\mathbb{R}$, define.
-$\mu_{\varphi}(J)=\langle\varphi,\mathbbm{1}_{J}(H)\varphi\rangle.$
-Then
-$\mu_{\varphi}([E-\varepsilon,E+\varepsilon])\geq\|\varphi\|^{2}-\frac{\|(H-E)\varphi\|^{2}}{\varepsilon^{2}}.$
+There exists a constant $p_{*}>0$ such that for all $E\in\mathbb{R}$, all $\eta>0$, and all neighbors $y\sim 0$,
+$\mathbb{P}\big(|\Gamma_{y}(E+\mathrm{i}\eta)|\leq 1\big)\geq p_{*}.$
 Proof.
-The spectral theorem implies
-$\|(H-E)\varphi\|^{2}=\int_{\mathbb{R}}(\lambda-E)^{2}\,\mu_{\varphi}(d\lambda)\geq\varepsilon^{2}\,\mu_{\varphi}\big(\mathbb{R}\setminus[E-\varepsilon,E+\varepsilon]\big),$
-which yields the conclusion. ∎
-For all $M>0$, define the $M$-strength barrier event at distance $1$ from the root of $\mathbb{T}$ by
-$\mathcal{B}_{M}=\Big\{\min_{y\sim 0}|V_{y}-V_{0}|\geq M\Big\}.$
-On $\mathcal{B}_{M}$, define
-$\psi=\delta_{0}+\sum_{y\sim 0}a_{y}\delta_{y},\qquad a_{y}=\frac{t}{V_{y}-V_{0}}.$
-Lemma B.2.
-On $\mathcal{B}_{M}$,
-$\|\psi\|^{2}\leq 1+\frac{dt^{2}}{M^{2}},\qquad\|(H-V_{0})\psi\|^{2}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
+Fix $z=E+\mathrm{i}\eta$ and $y\sim 0$. Conditioning on the potentials in $\mathbb{T}_{y}\setminus\{y\}$, the Schur complement formula gives
+$\Gamma_{y}(z)=\frac{1}{V_{y}-a-\mathrm{i}b}$
+for some random variables $a\in\mathbb{R}$ and $b>0$ that are independent of $V_{y}$. Hence
+$\mathbb{P}\big(|\Gamma_{y}(z)|\leq 1\,\big|\,a,b\big)=\mathbb{P}\big((V_{y}-a)^{2}+b^{2}\geq 1\,\big|\,a,b\big)\geq\mathbb{P}\big(|V_{y}-a|\geq 1\big).$
+Since the density of $V_{0}$ is even and strictly decreasing on $0,\infty)$, among all intervals of length $2$ the interval [1,1][-1,1] has the largest Gaussian mass. Therefore
+$\mathbb{P}(|V_{y}-a|\geq 1)\geq\mathbb{P}(|V_{y}|\geq 1)>0$
+for all $a\in\mathbb{R}$. Setting $p_{*}=\mathbb{P}(|V_{y}|\geq 1)$ and averaging over $a,b$ proves the claim. ∎
+Proposition B.2.
+Fix $d\geq 3$, $t>0$, and $D>0$. Then there exists a constant $c_{0}(d,t,D)>0$ such that for all $E\in[-D,D]$ and all $\eta\in(0,1]$,
+$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(E+\mathrm{i}\eta)\geq c_{0}.$
 Proof.
-By definition, $|a_{y}|\leq t/M$, so
-$\|\psi\|^{2}=1+\sum_{y\sim 0}|a_{y}|^{2}\leq 1+\frac{dt^{2}}{M^{2}}.$
-Also,
-$\big((H-V_{0})\psi\big)(0)=-t\sum_{y\sim 0}a_{y},$
-hence $|((H-V_{0})\psi)(0)|\leq t^{2}d/M$.
-At any $y\sim 0$, by definition of $a_{y}$ we have
-$\big((H-V_{0})\psi\big)(y)=(V_{y}-V_{0})a_{y}-t=0.$
-At vertices $u$ with $\operatorname{dist}(u,0)=2$, writing $y$ for the parent of $u$,
-$\big((H-V_{0})\psi\big)(u)=-ta_{y},\qquad\big|\big((H-V_{0})\psi\big)(u)\big|\leq\frac{t^{2}}{M}.$
-There are exactly d(d1)d(d-1) such vertices, therefore
-$\|(H-V_{0})\psi\|^{2}\leq\frac{t^{4}d^{2}}{M^{2}}+\frac{t^{4}d(d-1)}{M^{2}}\leq\frac{2t^{4}d^{2}}{M^{2}}.$
-Lemma B.3.
-Let $I\subset\mathbb{R}$ be Borel and write $P_{I}=\mathbbm{1}_{I}(H)$. If $\psi=\delta_{0}+\xi$, then
-$\mu_{0}(I)\geq\big(\sqrt{\mu_{\psi}(I)}-\|\xi\|\big)^{2}.$
-Proof.
-Since $\delta_{0}=\psi-\xi$,
-$\sqrt{\mu_{0}(I)}=\|P_{I}\delta_{0}\|=\|P_{I}(\psi-\xi)\|\geq\|P_{I}\psi\|-\|P_{I}\xi\|\geq\sqrt{\mu_{\psi}(I)}-\|\xi\|.$
-Squaring gives the claim. ∎
-Proposition B.4.
-Assume $t>0$, fix $M>0$, and set $\varepsilon=4t^{2}d/M$. Then on $\mathcal{B}_{M}$,
-$\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq\left(\sqrt{\frac{7}{8}}-\frac{\sqrt{d}\,t}{M}\right)^{2}.$
-Proof.
-By Lemma B.1 and Lemma B.2,
-$\mu_{\psi}([V_{0}-\varepsilon,V_{0}+\varepsilon])\geq\|\psi\|^{2}-\frac{\|(H-V_{0})\psi\|^{2}}{\varepsilon^{2}}\geq 1-\frac{2t^{4}d^{2}/M^{2}}{16t^{4}d^{2}/M^{2}}=\frac{7}{8}.$
-Also,
-$\xi=\sum_{y\sim 0}a_{y}\delta_{y},\qquad\|\xi\|^{2}=\sum_{y\sim 0}|a_{y}|^{2}\leq\frac{dt^{2}}{M^{2}},$
-so $\|\xi\|\leq\sqrt{d}\,t/M$. Apply Lemma B.3. ∎
+Fix $E\in[-D,D]$ and $\eta\in(0,1]$, and set $z=E+\mathrm{i}\eta$. Define the event
+$\mathcal{G}_{z}=\bigcap_{y\sim 0}\big\{|\Gamma_{y}(z)|\leq 1\big\}.$
+Since the forward subtrees $(\mathbb{T}_{y})_{y\sim 0}$ are disjoint and the potentials are i.i.d., the random variables $(\Gamma_{y}(z))_{y\sim 0}$ are independent. Hence, by Lemma B.1,
+$\mathbb{P}(\mathcal{G}_{z})\geq p_{*}^{d}.$
+Conditioning on the sigma-field generated by $(V_{x})_{x\neq 0}$, define
+$A_{z}=E+t^{2}\sum_{y\sim 0}\operatorname{Re}\Gamma_{y}(z),\qquad B_{z}=\eta+t^{2}\sum_{y\sim 0}\operatorname{Im}\Gamma_{y}(z).$
+Then $B_{z}>0$ and
+$R_{00}(z)=\frac{1}{V_{0}-A_{z}-\mathrm{i}B_{z}}.$
+Therefore,
+$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(z)=\frac{1}{\pi}\mathbb{E}\big[\operatorname{Im}R_{00}(z)\big]=\mathbb{E}\!\left[\int_{\mathbb{R}}\rho(v)\,\frac{1}{\pi}\frac{B_{z}}{(v-A_{z})^{2}+B_{z}^{2}}\,dv\right],$
+where $\rho$ is the density of $V_{0}$. On the event $\mathcal{G}_{z}$ we have
+$|A_{z}|\leq D+t^{2}d,\qquad 0<B_{z}\leq 1+t^{2}d.$
+Set
+$K_{1}=D+t^{2}d,\qquad K_{2}=1+t^{2}d.$
+Define, for $a\in\mathbb{R}$ and $b>0$,
+$F(a,b)=\int_{\mathbb{R}}\rho(v)\,\frac{1}{\pi}\frac{b}{(v-a)^{2}+b^{2}}\,dv.$
+Because $\rho$ is continuous and strictly positive, $F$ extends continuously to $\mathbb{R}\times[0,\infty)$ by setting $F(a,0)=\rho(a)$ (see, e.g., [13, Theorem 7.4]). Hence
+$\inf_{|a|\leq K_{1},\;0\leq b\leq K_{2}}F(a,b)>0.$
+Denote this infimum by $c_{1}$. On $\mathcal{G}_{z}$ we have $|A_{z}|\leq K_{1}$ and $0<B_{z}\leq K_{2}$, so
+$F(A_{z},B_{z})\geq c_{1}.$
+Consequently,
+$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(z)=\mathbb{E}\big[F(A_{z},B_{z})\big]\geq c_{1}\,\mathbb{P}(\mathcal{G}_{z})\geq c_{1}p_{*}^{d}.$
+Thus the claim holds with $c_{0}=c_{1}p_{*}^{d}$. ∎
 We are now ready for the proof of Lemma 6.2.
 Proof of Lemma 6.2.
-Fix $E\in[-D,D]$. Set $M=4\sqrt{d}t$ and $\varepsilon=4t^{2}d/M$, and define
-$\mathcal{A}_{E,\varepsilon,M}=\big\{|V_{0}-E|\leq\varepsilon/2\big\}\cap\mathcal{B}_{M}.$
-On $\mathcal{A}_{E,\varepsilon,M}$ we have $|V_{0}-E|\leq\varepsilon/2$, hence
-$V_{0}-\varepsilon,V_{0}+\varepsilon]\subset(E-2\varepsilon,E+2\varepsilon).$
-By Proposition B.4, there exists $c_{0}>0$ such that on $\mathcal{B}_{M}$ (hence on $\mathcal{A}_{E,\varepsilon,M}$),
-$\mu_{0}\big((E-2\varepsilon,E+2\varepsilon)\big)\geq\mu_{0}\big([V_{0}-\varepsilon,V_{0}+\varepsilon]\big)\geq c_{0}.$
-Taking expectation gives
-$\nu\big((E-2\varepsilon,E+2\varepsilon)\big)\geq c_{0}\,\mathbb{P}(\mathcal{A}_{E,\varepsilon,M}).$
-Observe that
-$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})=\int_{E-\varepsilon/2}^{E+\varepsilon/2}\rho(v)\,\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)\,dv.$
-Given $V_{0}=v$, the $d$ neighbor potentials are i.i.d. with law $\rho$, so
-$\mathbb{P}(\mathcal{B}_{M}\mid V_{0}=v)=\mathbb{P}(|V-v|\geq M)^{d}.$
-Lower bounding both factors on $|v-E|\leq\varepsilon/2$ yields
-$\mathbb{P}(\mathcal{A}_{E,\varepsilon,M})\geq\varepsilon\Big(\inf_{|v-E|\leq\varepsilon/2}\rho(v)\Big)\Big(\inf_{|v-E|\leq\varepsilon/2}\mathbb{P}(|V-v|\geq M)\Big)^{d}.$
-Both factors on the right-hand side can be bounded below by a constant depending only on $d$, $D$, and $\mathfrak{L}$. This completes the proof. ∎
+By Lemma 6.1, the measure $\nu$ is absolutely continuous with respect to Lebesgue measure, so there exists $n\in L^{\infty}(\mathbb{R})$ such that
+$\nu(d\lambda)=n(\lambda)\,d\lambda.$
+For $\eta>0$, let
+$P_{\eta}(x)=\frac{1}{\pi}\frac{\eta}{x^{2}+\eta^{2}}$
+denote the Poisson kernel. Since $m_{\mathbb{T}}$ is the Stieltjes transform of $\nu$, we have
+$\frac{1}{\pi}\operatorname{Im}m_{\mathbb{T}}(E+\mathrm{i}\eta)=\int_{\mathbb{R}}P_{\eta}(E-\lambda)\,n(\lambda)\,d\lambda=(P_{\eta}*n)(E).$
+By Proposition B.2, for every $E\in[-D,D]$ and every $\eta\in(0,1]$,
+$(P_{\eta}*n)(E)\geq c_{0}.$
+Now fix an interval $I\subset[-D,D]$. Integrating over $E\in I$ gives
+$\int_{I}(P_{\eta}*n)(E)\,dE\geq c_{0}|I|\qquad\text{for all }\eta\in(0,1].$
+Since $(P_{\eta})_{\eta>0}$ is an approximate identity and $n\in L^{\infty}(\mathbb{R})$, we have
+$\lim_{\eta\downarrow 0}\int_{I}(P_{\eta}*n)(E)\,dE=\int_{I}n(E)\,dE=\nu(I).$
+Therefore, $\nu(I)\geq c_{0}|I|$, as desired. ∎